### v1

# Climate in a Bottle: Towards a Generative Foundation Model for the Kilometer-Scale Global Atmosphere

Abstract
AI emulators offer a path to compressing, boosting limited ensembles, and improving the latency of interacting with petabyte-scale climate prediction data. However, prevailing auto-regressive paradigms offer limited flexibility, and are challenging to train on climate time horizons due to drifts, instabilities and component-coupling challenges. Conditionally generative models offer an appealing alternative. In this context we demonstrate a generative diffusion-based framework—Climate in a Bottle (cBottle)—for emulating global km-scale climate simulations and reanalysis on the equal-area HEALPix grid. cBottle consists of two model stages: a globally-trained coarse-resolution image generator that generates 100km (50k-pixel) fields given monthly average sea surface temperatures and solar conditioning, followed by a locally-trained 16x super-resolution stage that generates 5km (12.5M-pixel) fields; global super-resolution is made affordable using an overlapping patch-based multi-diffusion. Overall, cBottle shows promise as an emulator across a battery of climate model diagnostics, including diurnal-to-seasonal scale variability, large-scale modes of variability, tropical cyclone statistics, and trends of climate change and weather extremes. Moreover, cBottle is a step towards a foundation model, by bridging multiple data modalities (reanalysis and simulation) with corresponding utility beyond emulation to tasks such as zero-shot bias correction, climate downscaling, and channel in-filling. The code is available at https://github.com/NVlabs/cBottle.
1 Introduction
Knowing more, and planning better for the consequences of climate change are driving tremendous advances in the realms of climate simulation and Earth observation. With exascale computing and energy-efficient computational systems, petabytes of kilometer-scale climate data are being generated. Digital-twin efforts like the Destination Earth initiative (Bauer et al., 2021) and the Earth Virtualization Engines (EVE) project (Stevens et al., 2024) envision a world where high-fidelity global climate projections with local granularity are available to all with a degree of access and interactivity that provides a scientifically grounded basis for action. Further, digital twins aim to efficiently expose the data to new methods of analysis and proactive information production that will enable stakeholders to construct and interact with their own climate scenarios.
Easy access and low-latency interactivity, however, are nearly impossible with the current paradigm of petabyte-scale datasets generated and hosted by a handful of climate centers. Even if the access problem is unthrottled by moving data around, interacting with petabytes of data is virtually impossible. Generative artificial intelligence (AI) is poised to disrupt this paradigm by fully leveraging state-of-the-art architectures and algorithms optimized for performance on accelerated computing hardware.
For our purposes, we define a generative foundation model as a self-supervised climate simulator capable of producing realistic-looking samples that are climate-faithful measured by a set of diagnostics, and can be used to examine what-if scenarios not explicitly built into the model. Of course, climate science has long had such foundation models in the form of physics-based general circulation models. However, advances in AI are already fundamentally changing how weather and climate are simulated due to their ease of use, enhanced computational performance, and improved accuracy. Specifically, emulators of weather have proven to be skillful and orders-of-magnitude faster than traditional numerical simulation-based approaches to simulating the weather across spatial and temporal scales (Pathak et al., 2022; Lam et al., 2023; Bi et al., 2023; Price et al., 2024; Bodnar et al., 2024). Similarly, AI is proving to be efficient and faithful at emulating climate models and climate data, evaluated across a range of metrics on a wide variety of tasks (Watt-Meyer et al., 2024; Chapman et al., 2025; Cresswell-Clay et al., 2024; Watson-Parris et al., 2022). A majority of these developments have resorted to the auto-regressive modeling paradigm, much like numerical weather and climate models that step forward in time evolving the state of the system modeled by governing systems of equations. Masked training and representation learning have also been proposed (Lessig et al., 2023).
Diffusion modeling (Ho et al., 2020; Song et al., 2020a; Karras et al., 2022) offers an alternative to the auto-regressive time-stepping approach by learning—and sampling—from a distribution of states. This is especially attractive for rapid analyses and informatics that do not require the full temporal evolution of the system. By the same token, they are not susceptible to instabilities or drifts (i.e., hallucinations) inherent to auto-regressive models, though admittedly many auto-regressive emulators perform surprisingly well for large numbers of steps (Watt-Meyer et al., 2024; Cresswell-Clay et al., 2024). Further, diffusion models can be controlled by the user in a variety of ways, using guidance (Song et al., 2020a) to generate conditional distributions. Generative diffusion modeling has had tremendous impact across disciplines ranging from computer vision to science, and enjoys widespread adoption across the industry. Notably, it has been used in weather and climate for ensemble generation and uncertainty quantification (Li et al., 2024), probabilistic forecasting (Price et al., 2024), super-resolution (or downscaling) (Mardani et al., 2025), and data-assimilation (Rozet and Louppe, 2023; Manshausen et al., 2024).
A less explored yet promising direction is its application towards learning a compressed representation of massive climate datasets that generates, on demand, climate states steered by user-specified conditions and scenarios. Early examples included training deterministic emulators to predict the mean-state response of climate models to various forcings (Watson-Parris et al., 2022). More recent efforts have extended this approach to stochastic emulators in the diffusion framework (Bassetti et al., 2024; Quilcaille et al., 2022) as well as traditional dimensionality reduction approaches (Wang et al., 2025). But these approaches have been limited to a small number of channels (typically surface temperature) at a coarse spatial resolution.
We develop a generative diffusion-based framework that we call “Climate in a bottle” (cBottle) to emulate multimodal climate data. cBottle uses a cascaded approach (Ho et al., 2021) to generating km-scale samples: a coarse-resolution generator conditioned on user-specified inputs (time of day, time of year, and monthly-averaged sea surface temperature (SST)) and a 16x super-resolution module. Both steps use generative diffusion models, and either can be used independently, enabling generation of coarse-resolution samples and/or super-resolution to km-scale.
cBottle has a few key properties:
  1. 1.
    Multi-modality and domain transfer: Bridges ERA5 and ICON km-scale datasets, enables generation of km-scale climate states that exploits the information content contained across ERA5 and ICON
  2. 2.
    Versatility across a few useful tasks: Bias-correction, channel in-filling, downscaling
  3. 3.
    Climate faithfulness: Generates realistic climatology, climate variability across timescales, large-scale modes, and extreme weather statistics
  4. 4.
    Data efficiency: Can be trained on as little as 4 weeks of simulation output
  5. 5.
    Extreme compression: Encapsulates km-scale climate model and reanalysis outputs into a few GB of neural network weights, and offers 256x compression ratio per channel
  6. 6.
    Speed and scale: Low-latency generation of km-scale samples and scalability to higher resolutions and more modalities
These properties of cBottle offer potential pathways towards: (i) solving the data and informatics challenges of km-scale climate data; and (ii) building a generative foundation model for an interactive digital twin.
Mean sea-level pressure (generated) (ground truth)
Refer to caption Refer to caption
Surface air temperature (generated) (ground truth)
Refer to caption Refer to caption
Infrared(generated) (ground truth)
Refer to caption Refer to caption
Visible light (generated) (ground truth)
Refer to caption Refer to caption
Surface precipitation (generated) (ground truth)
Refer to caption Refer to caption
Figure 1: High resolution (HPX1024) samples from the ICON data (right) and synthesized by cBottle given sea surface temperature and time conditionings (left). Note these are not identically paired. For each field the same colormap is used, but to highlight the visual similarity we do not include color-bars. More quantiative comparisons will follow.
2 Methods
Figure 2: Climate in a Bottle (cBottle) Model Architecture. A cascaded approach with two complementary diffusion models that produces kilometer-scale climate data on demand. Inputs (left) include time-of-day, date-of-year, and SST, which feed into the Coarse Generation model (center) to produce 100km resolution outputs. These are then processed by the Super Resolution model (right) to transform coarse outputs to 5km resolution. This architecture enables climate emulation, channel in-filling, bias correction, downscaling, and what-if scenario exploration.
2.1 Overview
In this section we describe the components of our diffusion-based generative climate model shown in Figure 2. The architecture and diffusions are formulated on the sphere using the increasing popular HEALPix grid (Section 2.2). We train on two dense global datasets (Section 2.3) using a masked training strategy (Section 2.6). We introduce the preliminary details of diffusion modeling (Section 2.4) and our strategy for scaling to high-resolution data (Section 2.5). Finally, we present the architectures and hyper-parameters used for training our coarse image and video models, and the super-resolution model (Section 2.7).
2.2 HEALPix and Padding
Using AI for global data poses a geometric challenge of appropriately representing data on a spherical surface. To this end, HEALPix (Gorski et al., 2005) grids offer an attractive solution with an equal area grid structure that also has iso-latitude properties and a hierarchical indexing scheme for multi-resolution workflows. The HEALPix grid is constructed of 12 faces with each face containing a square $2^{l}$ x $2^{l}$ grid of pixels where $l$ represents the resolution level of the grid. This structure makes it very friendly for deep-learning algorithms, especially for convolutional architectures and has been used in (Krachmalnicoff and Tomasi, 2019), (Perraudin et al., 2019) and (Karlbauer et al., 2024). The equal area property is also particularly desirable for diffusion training. Due to these desirable properties, we use the HEALPix grid as the spherical discretization method in this work. A common convention for HEALPix grids refers to the length of each face $N_{side}=2^{l}$. We refer to two main global grids: HPX64 ($N_{side}=64$) corresponding to 100km resolution (49152 pixels) and HPX1024 ($N_{side}=1024$) at 6km resolution (12,582,912 pixels).
2.2.1 Global modeling
For coarse resolution applications (below HEALPix level 8), we use architectures similar to (Mardani et al., 2025) that were designed for 2D rectangular domains. This is done by adapting the convolution operations in the architecture to perform convolutions on each HEALPix face independently. However, boundaries between HEALPix faces are not handled by such an approach and can cause artifacts. To address this, we pad each face using the protocol outlined in (Karlbauer et al., 2024) by the number of pixels needed for the convolution operation instead of relying on any implicit padding methods supported by convolution operations.
Consider a field represented on a level $l$ the HEALPix grid, $x\in\mathbb{R}^{c_{in}\times 12\times 2^{l}\times 2^{l}}$. One may apply a discrete convolution kernel $\mathcal{K}_{conv}\in\mathbb{R}^{c_{out}\times c_{in}\times 2k+1\times 2k+1}$ as follows:
$x$ $=\left(x_{1},x_{2},\cdots x_{12}\right),$ $\quad x_{f}\in\mathbb{R}^{c_{in}\times 2^{l}\times 2^{l}}$ (1)
$x^{p}$ $=\mathcal{P}\left(x;k\right)=\left(x^{p}_{1},x^{p}_{2},\cdots x^{p}_{12}\right),$ $\quad x^{p}_{f}\in\mathbb{R}^{c_{in}\times(2^{l}+2k)\times(2^{l}+2k)}$ (2)
$y_{f}$ $=\mathcal{K}_{conv}\ast x^{p}_{f},$ $\quad y_{f}\in\mathbb{R}^{c_{out}\times 2^{l}\times 2^{l}}$ (3)
$y$ $=\left(y_{1},y_{2},\cdots y_{12}\right),$ $\quad y\in\mathbb{R}^{c_{out}\times 12\times 2^{l}\times 2^{l}}$ (4)
where $x_{f}$ and $y_{f}$ are the pixels of $x$ and $y$ respectively on the HEALPix face $f$, $\mathcal{P}(.;k)$ is the padding operation of (Karlbauer et al., 2024) and $\ast$ indicates the convolution operation. Figure SI3 shows a schematic of how the UNet architecture of (Song et al., 2021) is adapted to work on global HEALPix fields.
2.2.2 Local modeling
Modeling of high resolution fields, like kilometer-scale climate model outputs, is prohibitively expensive using global approaches. To make this tractable, we use the multi-diffusion approach outlined in Section 2.5 and adapt that to spherical manifolds using the HEALPix grid and HEALPix padding. We define a new spatial sampling operator ${\mathcal{G}_{i}}:\mathcal{X}\rightarrow\mathcal{X}_{\text{local},i}$ as a composition of the padding operator $\mathcal{P}(.;N/2)$ and the 2D sampling operator $\mathcal{F}_{i}$ defined in Section 2.5. Here $N$ represents the patch size used in the multi-diffusion approach. The global denoiser can then be represented as
$\mathcal{D}_{\theta}(\textbf{x}_{t})$ $=\frac{\sum_{i}\mathcal{G}_{i}^{-1}\left(\Phi_{\theta}\left(\mathcal{G}_{i}(\textbf{x}_{t})\,\middle|\,\mathcal{G}_{i}(y),\sigma\right)\right)}{\sum_{i}\mathcal{G}_{i}^{-1}(J_{N})}$ (5)
$\mathcal{G}_{i}(\textbf{x}_{t})$ $=\mathcal{F}_{i}(\mathcal{P}(\textbf{x}_{t};N/2)),$ (6)
Figure 3 shows a schematic of this approach.
Figure 3: Schematic showing the multi-diffusion approach for spherical geometry with HEALPix grids. A sequence of the overlapping patch sampling protocol is shown going clockwise from the top left. Each colored region indicates a single patch and overlap between the patches minimizes boundary artifacts by combining patch representations into a global representation following 14.
2.3 Datasets
Targets Shape Datasets used for 2D 3D
Single level fields
liquid water path X icon+era
ice water path X icon+era
2m air temperature X icon+era
surface zonal wind X icon+era
surface meridional wind X icon+era
Outgoing long wave radiation X icon
Outgoing shortwave radiation X icon
mean sea level pressure X icon+era
precipitation (total) X icon+era
downwelling SW at surface X icon
sea surface temperature (snapshot) X icon+era
sea ice concentration X icon+era
Pressure level variables @ [1000, 850, 700, 500, 300, 200, 50, 10] mb
$U$ P, X icon+era
$V$ P, X icon+era
$T$ P, X icon+era
$Z$ P, X icon+era
Inputs
Monthly mean SST X
era5 = AMIP Monthly SST
icon = rolling mean
Day of year
Hour of day
Label
Table 1: Input and target variables with dataset sources and dimensionality. X = space, P = levels.
The inputs and outputs we trained with are displayed in Tab. 1. Briefly, the models are trained to generate 3D states of the atmosphere and some key 2D variables relevant to climate impacts (precipitation, surface air temperature, downwelling short-wave) and observations (top of atmosphere radiative fluxes).
Our models are trained on two global datasets: the ECMWF Reanalysis version 5 (ERA5)(Hersbach et al., 2020) which is observationally constrained, and a global storm resolving simulation with the ICON climate model (Hohenegger et al., 2023), which is a free-running simulation unpaired with reality.
ERA5
ERA5 is a commonly used dataset that represent the best guess of the global atmospheric state at a coarse resolution of 25 km. Briefly, it is created by running a data assimilation algorithm that incorporates observations into ECMWF’s numerical model. However, it is not available at a high enough resolution in space or time to resolve meso-scale motions on kilometer scale, and in particular deep convection. The ICON simulation fills in this gap since it is available at a global 5km resolution, but is not paired with reality. For pre-processing, we retrieved hourly ERA5 data for 1980–2018 (inclusive) via the data lake at the National Energy Research Computing Center (NERSC) at $0.25^{\circ}$ resolution on the lat-lon grid. The 3d atmospheric states are available on pressure levels. We then regridded this using bilinear interpolation to the HEALPix grid with $N_{side}=256$. For training the coarse-resolution model, this coarsened by averaging pooling to an $N_{side}=64$. We train on years 1980–2017 (inclusive) and validate on 2018, though many of our climate diagnostics below are also computed on the validation set (as is common in generative modeling).
ICON Cycle 3
The ICON data is a free-running simulation with the ICON atmospheric model coupled to a dynamic ocean and land. The atmospheric component solves the nonhydrostatic fluid mechanics equations on a global ico-sahedral mesh with a resolution of around 5 km. Unlike ERA5 the model explicitly resolves certain convective motions so a convection parameterization is not used. So even beyond the simple increase in resolution, we expect this to increase the fidelity of the precipitation and cloud fields in this dataset relative to ERA5 where these processes are parameterized. Furthermore, because the ocean and land are dynamically coupled to the atmosphere, we expect this run to obey conservation of heat, momentum and moisture between these different components, unlike in the ERA5 where the observations (rather than conservations) are king.
We obtained O(PB) of ICON data (ngc3028) of the nextGEMS cycle 3 simulations (Koldunov et al., 2023) in zarr format from DKRZ’s JupyterHub server. This dataset was stored on the HEALPix grid with $N_{side}=1024$, corresponding roughly to a resolution of $\sqrt{A/(12N_{side}^{2})}=6.3\text{km}$. Prior to our handling of the data it was interpolated using nearest neighbors from the native icosahedral grid. Five years of this data are available at 3-hour (3D) and 30-minute (2D) resolution in time. Conveniently, this dataset featured pre-coarsened data (again using averaging pooling) at our coarse-resolution of $N_{side}=64$. We interpolated this coarse data to fixed pressure levels using linear interpolation in the vertical direction. To fill in values at pressure level locations that are below the surface, we use extrapolation based on hydrostatic balance constraints and an assumption of a constant temperature lapse rate of $6.5$ K/km for temperature and geo-potential. For all other variables, we use constant extrapolation down from the surface. This procedure approximately reproduces ERA5’s undocumented procedure for filling-in below-surface levels. We train on 2020-01-20 03:00:00 through 2024-03-06 12:00:00 (inclusive) and validate on 2024-03-06 15:00:00 through 2025-07-22 00:00:00.
AMIP SST Forcing
Our models take a monthly time-scale SST as an input. For the ERA5 data which is paired with reality, we linearly interpolate the adjusted monthly SST time series following the AMIP protocol (Gates et al., 1999). Specifically we downloaded the file: input4MIPs.CMIP6Plus.CMIP.PCMDI.PCMDI-AMIP-1-1-9.ocean.mon.tosbcs.gn from the Earth System Grid Federation, and use bilinear interpolation to map it to HPX64 resolution. For the ICON data on the other hand, we compute a rolling window average in time with a window size of 30 days based on the HPX64 spatial and daily temporal resolution data in the ICON dataset.
2.4 Diffusion models
Background
Let $p_{\text{data}}(\mathbf{x})$ denote the data distribution with standard deviation $\sigma_{\text{data}}$. We define the diffusion process as a pair of Markov chains over the states $\{\mathbf{x}_{0},\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$. The forward process gradually corrupts the data by adding i.i.d. Gaussian noise with standard deviation $\sigma$, eventually transforming the data into pure noise. Specifically, the final state satisfies $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\text{max}})$, where $\sigma_{\text{max}}\gg\sigma_{\text{data}}$. The reverse process then attempts to denoise samples from this noisy distribution, ideally recovering samples from the original data distribution, i.e., transforming $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\text{max}})$ back into $\mathbf{x}_{0}\sim p_{\text{data}}(\mathbf{x})$.
In this work, we adopt the score-based diffusion framework (Song et al., 2021; Mardani et al., 2025), which learns this denoising process via a score-matching network. The training objective is given by
$\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathbf{x}\sim p_{\text{data}}}\mathbb{E}_{\sigma\sim p_{\sigma}}\mathbb{E}_{\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})}\left\|\mathcal{D}_{\theta}(\mathbf{x}+\mathbf{\epsilon};\sigma)-\mathbf{x}\right\|^{2},$ (7)
where the learned denoiser $\mathcal{D}_{\theta}(\cdot;\sigma)$ approximates the score function according to $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)\approx\frac{\mathcal{D}_{\theta}(\mathbf{x};\sigma)-\mathbf{x}}{\sigma^{2}}$. Here, $p_{\sigma}$ denotes the noise level distribution used during training.
To sample from the model, we employ the Elucidated Diffusion Model (EDM) sampler proposed by (Karras et al., 2022), which solves the corresponding reverse-time ordinary differential equation (ODE):
$fd\mathbf{x}=-\dot{\sigma}(t)\sigma(t)\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma(t))\,$ (8)
where $\dot{\sigma}(t)$ is the time derivative of the noise schedule $\sigma(t)$.
Specifically, we use Algorithm 2 from Karras et al. (2022), with the $S_{churn}$ parameter set to 0.
Diffusions on the sphere
To train diffusion models on the sphere, the noise $\epsilon_{i}\sim~{}N(0,1)$ is drawn independently for every pixel $1\leq i\leq 12\cdot 4^{l}$. Because the HEALPix pixels are all equal area, we conjecture that this is identical to the pixel-wise average of a finite-variance continuous Gaussian process on the sphere.
Tuning noise schedules for non-stationary signals
The choice of noise schedule during training ($p_{\sigma}$) and inference is crucial for the quality of diffusion model. We find that settings commonly used for image generation (Karras et al., 2022) do not work well across the dozens of atmospheric observables, each with different spectral characteristics. Moreover, compared to image datasets, these data are highly non-stationary in space and time. Climatology explains much of the variance of the atmosphere, but the mean of all cat images does not look like a cat. In particular, we find that our models cannot produce the seasonal or diurnal cycles using the settings from Karras et al. (2022) (see Sec 3.2.4).
The basic issue is known as the signal leak bias when a model is not trained with enough noise to swamp the signal (Everaert et al., 2023). As a rough criterion, the noise levels sampled during training should be large enough to completely mask/unmask the signal when represented in any basis—not just the canonical basis in pixel space. Intuitively, if we fail to add enough noise then large-variance signals like season or hour of day can be detected simply by filtering the data enough. More concretely, the smallest $\sigma_{min}$ should be chosen so that
$v^{\prime}(x+\sigma_{min}\epsilon)\approx v^{\prime}x$ (9)
and the largest so that
$v^{\prime}(x+\sigma_{max}\epsilon)\approx\sigma_{max}v^{\prime}\epsilon$ (10)
for any vector $||v||=1$.
Criteria for $\sigma_{\{min,max\}}$ can be derived by starting from the observation that $E|v^{\prime}\epsilon|^{2}=|v|^{2}=1$, so that (9) and (10) are satisfied by choosing
$\sigma^{2}_{min}$ $\lesssim\inf_{||v||=1}E|v^{\prime}x|^{2}$ (11)
$\sigma^{2}_{max}$ $\gtrsim\sup_{||v||=1}E|v^{\prime}x|^{2}.$ (12)
The signal leak bias occurs when $\sigma_{max}^{2}$ is too small. When $\sigma_{min}$ is too large, the diffusion model will not denoise the lower-amplitude modes of the data.
Note that (12) is equal to the variance explained by the maximum principal component of the data. Recall that the overall variance can be decomposed $n=TrCov(x)=\sum_{i}s^{2}_{i}$ where $n$ is the dimension of $x$ (pixels and channels). To understand the behavior for changing $n$ it is useful to discuss the fraction of variance explained by a singular value $\tilde{s}^{2}_{i}=s_{i}^{2}/n$, and expect this fraction to converge as we refine the resolution. This implies then, that $\sigma_{max}^{2}\gtrsim n\tilde{s}_{1}^{2}=O(\sqrt{n})$.
In atmospheric science, the seasonal cycle and diurnal cycles are the dominant modes of variability since they are: 1) planetary in spatial scale; and 2) large in amplitude. In mid-latitude land locations, the daily high and low temperature are separated by $\sim 5-10^{\circ}$, while the seasonal cycle is $\sim 10-20^{\circ}$. Other variables don’t have as strong a seasonal cycle over as large an area so we take the conservative estimate that the seasonal cycle explains 5% of the global variance across our whole variable set. The HPX64 model has about 30 channels in total and 49152 pixels. So this would imply a $\sigma_{max}^{2}\gtrsim 300$. This is far larger than the range of the hyper parameters used for this resolution in typical image generation (Karras et al., 2022). Similar analysis can explain the observation that generating higher-resolution images requires larger noise levels (Chen, 2023).
To estimate $\sigma_{min}$ one could also use principal component analysis in theory, but in practice this is difficult. In practice, power spectral analysis works well for identifying high frequency modes. We also often choose $\sigma_{min}$ based on the precision required for practical applications. Choosing $\sigma_{min}$ too large manifests as an incorrectly flat spectra for the highest wave numbers (see e.g., the Z500 power spectra presented by Price et al. (2024)).
While one can tune the log-normal distribution to fully cover the dynamic range of the data $[\sigma_{min},\sigma_{max}]$, we have found that sample quality is sensitive to shifts in the mean of the log-normal distribution. As a simpler approach, we propose a log-uniform noise schedule
$p_{\sigma}\propto\sigma^{-1},\sigma_{min}\leq\sigma\leq\sigma_{max}.$ (13)
Because this is uniform in $\log\sigma$ we call it a “log-uniform” noise distribution. In earlier tests on the 2D data alone without SST conditioning, we achieved similar results with the power law distribution used by Price et al. (2024). In practice, we set $\sigma_{min}=0.02$ and $\sigma_{max}=200$, settings which span the dynamic range of several of our variables (see Figure SI12). Although in hindsight we should have decreased it further for the smoothest fields like Z500. For inference, we use the same sampler as Karras et al. (2022) but with these settings.
2.5 Multi-diffusion
Emulating ICON at HPX1024 resolution poses substantial computational challenges due to the massive size of this km-scale global dataset. Directly applying a traditional diffusion model to the full-resolution domain is infeasible as it would require over 2,000 GB of GPU memory, assuming a linear relationship between spatial resolution and memory footprint. This far exceeds the capacity of current hardware, even on high-end multi-GPU systems.
To address the computational bottleneck and take advantage of the inherently local nature of super-resolution tasks, multi-diffusion (Bar-Tal et al., 2023) is adapted as a scalable and memory-efficient sampling strategy. It enables high-resolution generation without sacrificing global spatial coherence or requiring model architecture changes.
The central idea behind multi-diffusion is to replace the global denoiser used in standard sampling algorithms with an aggregated local denoiser. At each sampling step, the aggregated local denoiser operates on the global feature map in a sliding-window fashion, denoising it patch by patch. These locally denoised patches are then reassembled to reconstruct the global feature map. This procedure is illustrated in Figure SI5.
Given a pretrained local denoiser $\Phi_{\theta}$ and a set of spatial sampling operators ${\mathcal{F}_{i}}:\mathcal{X}\rightarrow\mathcal{X}_{\text{local},i}$, the global denoiser can be represented as an aggregation of local denoisers:
$\mathcal{D}_{\theta}(\textbf{x}_{t})=\frac{\sum_{i}\mathcal{F}_{i}^{-1}\left(\Phi_{\theta}\left(\mathcal{F}_{i}(\textbf{x}_{t})\,\middle|\,\mathcal{F}_{i}(y),\sigma\right)\right)}{\sum_{i}\mathcal{F}_{i}^{-1}(J_{N})},$ (14)
where $\mathcal{F}_{i}^{-1}:\mathcal{X}_{\text{local},i}\rightarrow\mathcal{X}$ denotes the inverse operation that maps local patches back to the global space, $y\in\mathcal{Y}$ denote the spatial conditions, $\mathcal{X}_{local}=\mathbb{R}^{c_{out}\times N\times N}$ denotes the local target space, $\mathcal{Y}=\mathbb{R}^{c_{in}\times N\times N}$ denotes the local condition space, $N$ is the patch size, and $J\in\mathcal{X}_{\text{local}}$ is an all-one matrix of size $N\times N$ used to account for the number of overlapping patches at each location in the global map.
A fundamental design principle of multi-diffusion is the global sharing of the latent representation $X$ across all local patches. The spatial sampling operators ${\mathcal{F}_{i}}$ define the patch layout, including the extent of overlap between adjacent regions and the total number of patches required to span the full domain. Introducing overlap is essential for reducing boundary artifacts and promoting consistency across patches. However, increasing the overlap also raises the computational cost, as it leads to a larger number of denoiser evaluations during sampling. As demonstrated in the ablation study in Subsection 14, strong performance can be maintained even with minimal overlap, enabling a favorable trade-off between quality and efficiency.
2.6 Multi-modal masking
Since we intend our framework to extend to a variety of data sources, with different coverage in space, time, and the included fields, we have designed the model to handle masked training. For example, we learned that the ERA5 shortwave and longwave radiation fields are corrupted by a known, but undocumented, bug which appears as clear stripe-like artifacts that point North-West (see Fig 17). This is caused be a random seed in the radiation scheme being set to $lat+lon$ in every atmospheric column (Chantry, 2024). Therefore, as noted in Tab. 1, we do not use the fields from ERA5 for training our models, and instead rely on the multi-modal diffusion modeling to fill in this gap.
Our approach is to simply train a diffusion model to generate the union of all channels present in all the datasets, but have modified the code to support masking. This approach may seem inflexible compared to flexible architectures (Bodnar et al., 2024) but it is reasonable to scale since it has long been known how to represent the state of the atmosphere using dense grids of the known state variables.
Specifically, during training, any masked values in the noisy image $x+n$ and target $x$ are filled with 0 before calling the network and computing the loss. We then normalize the loss by the fraction of masked out data points, so that the training equally weights modalities with different numbers of observables. Psuedo-code for this approach is shown in SI E.
2.7 Architectures and hyperparameter settings
cBottle-3d cBottle-video cBottle-SR
Number of parameters 149M 282M 330M
Target channels 3D 3D 2D
Architecture Song UNet Song UNet Song UNet
Learned Position Encoding
Calendar embedding
Trained on ICON+ERA5 ICON+ERA5 ICON
Grid HPX64 (49152) HPX64 (49152) 128x128 crops of HPX1024
Training domain global global crops
Total examples trained 9,856,000 2,160,000 95,520 samples / 191M patches
Batch size 64 (3M images) 256 (3-10M) 32 960
Optimizer Adam Adam SGD
Learning rate 0.0001 0.0001 0.0001
Noise distribution ($p_{\sigma}$) log-uniform 0.02–200 log-uniform 0.02–1000 log-normal
$P_{\text{mean}}=-1.2$ $P_{\text{std}}=1.2$
Computational Profile
Hardware trained on 16xH100 32xH100 64xH100
Examples/GPU-second (train) 6.7 0.32 35.4 patches
Table 2: Comparison of cBottle model variants.. A training example is defined a single HPX64 image for the the coarse models but is a single 128x128 crop for cBottle-superres.
We mostly reuse the standard UNet architecture used in diffusion modeling (the so-called Song-UNet in Karras et al. (2022); Mardani et al. (2025) and others). However, it was necessary to add position and temporal embedding (Figure SI2 to allow the diffusion model to learn non-stationary signals in space and time, and temporal attention layers to support video diffusion. Table 2 summarizes our three model variants.
Calendar Embedding
Producing an accurate seasonal and diurnal cycle required a calendar and hour embedding strategy. We take care to use a embedding that respects the periodicity of these inputs. This is given by
$f_{4k}$ $=\cos\left(2\pi k\frac{\text{doy}}{365.25}\right)$ (15)
$f_{4k+1}$ $=\sin\left(2\pi k\frac{\text{doy}}{365.25}\right)$ (16)
$f_{4k+2}$ $=\cos\left(kt\right)$ (17)
$f_{4k+3}$ $=\sin\left(kt\right).$ (18)
Here, the day of year (doy) is a scalar representing the time since the beginning of the year, and the time of day ($t$) is the local solar time in units radians $t=2\pi(\frac{\text{second of day}}{86400}+\frac{\text{lon}}{360})$ [^1] Due to an implementation error we actually used $second-lon$.. In the architectures below we use $k=1,\ldots,8$, and concatenate these inputs along with other image-like conditioning (e.g., monthly-averaged SST) and noisy target image.
Learned position embedding
To allow the diffusion model to learn non-stationary spatial structures like climatology, we add a learned position embedding. After the initial convolution embedding of the image-like inputs (calendar, noisy image and monthly-averaged SST), a learned positional embedding for each pixel $49152\times c$ is added. The position embedding is initialized with independent standard normal random numbers.
2.7.1 Video Generation Extension
Let $x\in\mathbb{R}^{T\times C\times X}$ denote an atmospheric trajectory of $T$ evenly spaced frames with $X$ pixels each. While the coarse image cBottle produces realistic atmospheric snapshots, $T$ successive samples in time produce a sequence that lacks temporal coherence due to the independence of each sample.
Temporal Attention
We extend cBottle to video generation by augmenting the 2D UNet architecture with temporal self-attention layers, following factorized space-time approaches (Ho et al., 2022; Blattmann et al., 2023). The architecture preserves the UNet backbone of cBottle, but introduces temporal layers at the end of spatial blocks. Within each UNet block, we first apply spatial modules with the time dimension folded into the batch dimension. For blocks with temporal attention, we fold the spatial dimension into the batch dimension, and then apply bidirectional self-attention across frames per pixel, using the channels as the embedding dimension. A learned relative-time embedding $R_{\Delta t}\in\mathbb{R}^{2T-1}$ capturing frame-to-frame dependencies based on temporal distance $\Delta t$ is added to the pre-softmax logits. The temporal attention output is then residually added to the main branch.
Masked-Conditioning
To enable flexible inference capabilities, we adopt a masked-conditioning training scheme inspired by MCVD (Voleti et al., 2022). During training, we randomly sample mask patterns $m\in\{0,1\}^{T}$ (where 1 indicates observed frames) from a categorical distribution $p_{mask}$ that blends several tasks: Bernoulli random frame dropout, masking all but the first or last few frames contiguously, masking middle frames while keeping endpoints visible, and full sequence masking. On top of the calendar embedding and monthly-mean SST conditioning that the image-model uses, we add this masked sequence (ground truth at unmasked positions and zeros at masked positions) and the binary mask indicator as conditioning concatenated channel-wise with the input. The video diffusion training objective is identical to the image diffusion objective, but with this additional per-sample masked-conditioning and the loss averaged across the $T$ frames.
2.7.2 Model Settings
cBottle-3D
The coarse image model generates HPX64 fields. The coarse resolution image generator uses the standard settings for the Song-UNet presented by Karras et al. (2022), but we found the 3D training was made more stable by increasing base model embedding dimension from 128 to 192. The source dataset is identified with a label (ERA5 or ICON) which is randomly dropped for 1/4 samples to encourage the model to learn features across both datasets. This label is provided to the UNet in the same manner as the imagenet class label in Karras et al. (2022). Self-attention is only used on the coarsest resolution stage (8x8). All convolutional layers are implemented on the HPX grid using the padding approach described in Section 2.2.1. 32 channels are devoted to calendar embedding. The model was trained from scratch on a 50-50 mix of ICON and ERA5 data on 32 H100 GPUs. Even (odd) ranks were trained with ERA5 (ICON). The model was trained to 3 million samples (with repeats) with a batch size of 64 ($\sim 50000$ steps) when the log probability loss started to plateau, and then the batch size was increased to 256 until 10 million samples. The Adam optimizer (Kingma and Ba, 2014) was used with a learning rate of $10^{-4}$. During training the model uses the log-uniform noise distribution ($p_{\sigma}$) with $\sigma_{min}=0.02$ and $\sigma_{min}=200$. For inference, we use the power law sampler proposed by Karras et al. (2022) with these same settings.
This unified training approach enables the video model to perform multiple tasks at inference time: generate entire sequences unconditionally, predict future frames given initial states, or infill arbitrary temporal gaps.
cBottle-video
The video model generates 3-day sequences with 6-hour resolution ($T=12$ frames) at the same HPX64 resolution of the image-based cBottle. Because $T$ correlated frames carry more redundant signal, we keep the same log-uniform noise schedule (with $\sigma_{min}=0.02$) and increase $\sigma_{max}$ from 200 to 1000 in both training and inference to preempt signal leak bias (see Subsection 2.4). We insert temporal‐attention layers into every encoder block and into two positions within each decoder resolution stage: the up‑sampling block and the final block. We weight each of the four masking tasks defined above as follows: random frame dropout (0.3), block-wise masking (0.25), endpoint interpolation (0.1), and full sequence masking (0.35).
Similar to the coarse image model, a base model embedding dimension of 128 sufficed to model the 2D dataset’s 12 fields. Given the video model’s expanded 124-channel input on the 3D dataset, we increased the UNet width to 256 channels, finding it to converge markedly faster and achieve higher log-likelihood than at 192 channels. The model was trained from scratch like the image model on 32 H100 GPUs, using the same Adam optimizer (Kingma and Ba, 2014) and learning rate of $10^{-4}$. It was trained to 2.16M video samples with a batch size of 32.
cBottle-SR
The super-resolution model, which performs upscaling from HPX64 to HPX1024 resolution (a factor of 16), is trained within a HEALPix multi-diffusion framework described in 2.2.2. In this setting, the global input data—including latent variables, conditioning fields, and high-resolution targets—are partitioned into spatial patches. Each patch is then independently processed by a local denoising model, which is trained to reconstruct high-resolution content based on temporally and spatially aligned inputs. This approach enables efficient learning of fine-scale features while preserving global coherence through patch-wise training across the full spherical domain.
As illustrated in Figure SI4, this training strategy adapts the standard denoising score matching loss (see Eq. 7) to operate at the patch level. Specifically, the modified objective is defined as
$\min_{\theta}\mathbb{E}_{\mathbf{x}\sim p_{\text{data}}}\mathbb{E}_{\sigma\sim p_{\sigma}}\mathbb{E}_{\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})}\mathbb{E}_{i\sim\mathcal{U}}\Big{|}\Big{|}\Phi_{\theta}\Big{(}\mathcal{F}_{i}(\mathbf{x})+\mathbf{\epsilon};\mathcal{F}_{i}(y),\sigma\Big{)}-\mathcal{F}_{i}(\mathbf{x})\Big{|}\Big{|}^{2},$ (19)
Like cBottle-3d we use the Song-UNet with 128 hidden channels. Since it is a local model, we do not need to use the HEALPix padding shown in SI3.
A patch size of 128$\times$128 is used. This is chosen to be much larger than 16x super-resolution factor but is still affordable. To cope with non-stationary spatial features like topopgraphy, 20 learnable positional embeddings with a spatial size of HPX1024 grid are introduced and initialized using sinusoidal functions. The inputs of the network consist of (1) the HPX64 data re-gridded bilinearly to the HPX1024 128$\times$128 patch in question, (2) the global position embeddings for this patch, and (4) the global HPX64 data interpolated to the 128×128 lat-lon grid. These are concatenated with the latent before further processing by the encoder. The model was trained to 95520 ICON samples (corresponding to 191 million patches, with repeats) with a batch size of 960. The SGD optimizer was used with a learning rate of $10^{-4}$. During training the model uses the log-normal noise distribution of Karras et al. (2022) ($p_{\sigma}$) with $P_{mean}=-1.2$ and $P_{std}=1.2$. For inference, $\sigma_{max}=800$.
3 Results
3.1 ERA5 in a Bottle
3.1.1 Variability
A useful generator of climate states should exhibit realistic variability across independent samples from phenomena operating on diurnal, synoptic, seasonal, and interannual time scales that, when averaged, produce a realistic climatology. To assess this we perform an Atmosphere Model Intercomparison Project (AMIP) control run with prescribed monthly-averaged SSTs (Eyring et al., 2016; Gates et al., 1999). We generate samples every hour from 1940 to 2021 (inclusive). It should be stressed that each atmospheric state is completely independent in each timestep and only shares the same conditioning. This was a large-scale inference, requiring $\sim 1024$ GPU-hours on the H100 GPU.
We now present a series of common diagnostics used to assess the quality of the generated samples and statistics. Additional analyses are presented in Section G.
Figure 4: Seasonal cycle of sea ice concentration for ERA5 emulation and ground truth, computed over the test period.
Climatology and seasonal cycle
Figure 4 shows that cBottle correctly generates the seasonal evolution of Arctic sea ice concentration. As in the ground truth, generated ice approaches 100% concentration during boreal winter (a,f) and spring (b,g). During boreal summer (c,h) and fall (d,i) the sea ice margin retreats and its concentration is reduced locally over the Beaufort Sea, Chukchi Sea and Bering Sea – poleward of Alaska and Siberia, also reflected in the annual climatology (e,j). These results are evidence that the time-of-year conditioning in cBottle works as intended. The seasonal cycle is similarly realistic for other fields (see Figure SI10). The annual mean climate of precipitation of the cBottle generated samples also matches the ERA5 and ICON data (see Figure SI11).
Tropical cyclones
Consistent with realistic synoptic variability, tropical cyclones are generated in appropriate locations, with appropriate occurrence probability magnitudes (Figure 5). This is assessed using a tracking algorithm based on TempestExtremes v2.1 (Ullrich et al., 2021). The following filters are applied to identify candidate features: (1) Mean sea level pressure must increase by at least 200 Pa within a 5.5 great-circle distance (GCD) from the candidate point; (2) The difference between geopotential heights at 300 hPa and 500 hPa must decrease by at least 58.8 m2/s2 within a 6.5 GCD; (3) Wind speed must exceed 10 m/s; (4) The latitude of the candidate point must fall within the range of -45 to 45. As cBottle lacks temporal coherence, cyclone detections for both Target ERA5 and cBottle-ERA5 are computed without temporal continuity tracking to ensure a fair comparison.
Figure 5: Tropical cyclone occurrence probability derived from ERA5 (1980–2017) and cBottle (1940–2021), presented as heatmaps normalized per $km^{2}$.
Diurnal Cycle
Figure 6: Diurnal cycle diagnostics for surface precipitation. (top row) Amplitude of the diurnal cycle. (bottom row) signal to noise ratio.
The diurnal cycle of rainfall should only be detectable above the noise of other modes of variability in certain subregions of the planet. In Figure 6, we compare the amplitude and signal-to-noise ratio of the diurnal cycle in cBottle with ERA5 (see Sec B for details). cBottle appears to reproduce even small signals like the locally detectable subtropical stratocumulus drizzle diurnal cycles (see SNR). Morever, there is a realistic spatial structure of the precipitation diurnal cycle amplitude including spatial details interior to tropical continents – and even the halo that surrounds the Maritime Continent presumably from the emanating diurnally phased convection propagation there. See the SI Fig SI1 for the same analysis with surface air temperature.
3.1.2 Northern Annular Mode
We next assess a planetary scale mode of internal variability. In the northern hemispheric midlatitudes, the Northern Annular Mode is the leading empirical orthogonal function (EOF), which accounts for one-fifth of overall variance in the 1000 hPa geopotential height (REF). Figure 7 shows that the leading EOF computed from cBottle generations exhibits the expected spatial structure of the NAM as well as percent variance explained. Moreover, Figure 8 shows that the daily variability of cBottle’s NAM is realistically large and even contains a low frequency envelope that is in phase with interannual fluctuations in the ERA5 NAM state, suggesting learnt associations between the generated NAM state and the SST input to cBottle.
Figure 7: Arctic Oscillation or the Northern Annular Mode (NAM) computed based on Nov-Apr Z1000 monthly mean anomalies. Text overlay indicates the percentage of the total variance attributed to the plotted EOF mode. (left) cBottle. (right) ERA5. Computed over 1970-2010.
Figure 8: Arctic Oscillation or the Northern Annular Mode (NAM) daily indices between 1980 and 1995 computed as projections of daily Z1000 anomalies onto the ERA5 NAM mode. Top - Red/Blue filled curves: ERA5 3 month running mean, solid black line: cBottle 3 month running mean, grey filled region: cBottle ±$\pm 2\sigma$ interval around the mean. Bottom - Solid line: ERA5 daily AO index, red circles: cBottle daily AO index.
3.1.3 ENSO Variability
Appropriate conditioning on SST is confirmed by focusing on the El Nino Southern Oscillation (ENSO). Figure 9 demonstrates that cBottle produces realistic ENSO precipitation variability, measured as the regression of its generated precipitation against the Nino3.4 index associated with its input SST. Within the tropics, increases of Central Pacific precipitation during the positive phase of ENSO associate with reductions of generated precipitation over tropical Africa, Indonesia and South America, consistent with Walker Cell modulation. These results add to the evidence of the cBottle’s ability to condition generated variability appropriately on SST, in addition to time of year and time of day.
Figure 9: El Niño-Southern Oscillation. Regression coefficient of the Nino3.4 index against monthly average precipitation. (left) ERA5. (right) Climate in a bottle. Computed over 1990-2020 inclusive.
3.1.4 Non-stationarity: Heat waves
Figure 10: Time series of global mean anomalies from ERA5’s monthly climatology computed over 1980-2010. (a) surface air temperature, (b) precipitation, (c) 500-mb geopotential. 95% confidence-intervals for the annual mean anomaly are computed by bootstrapping over the monthly averages. (d) change in probability of exceeding 90-percentile of Z500 (Z500-90p) computed for 1990-2010 for all land points. Also shows Z500-90p trend maps for cBottle-ERA5 (e) and ERA5 (f).
Figure 10 shows the secular trends in global mean quantities for ERA5 and cBottle. cBottle noticeably misses the climate change trend in surface fields (temperature and precipitation), but does better for tropospheric quantities like Z500. Taking this caveat, we opt to use Z500 to analyze the trend in extreme heat—this is reasonable since surface heat waves are often linked to anomalous Z500 at atmospheric blocking patterns. For example, the 2021 Pacific Northwest heat wave feature +4 standard deviation height anomalies (Mass et al., 2024). cBottle is able to reproduce observed trends in the occurrence of heights exceeding the 90th percentile in space (Figure 10e,f), albeit with reduce amplitude in some regions. The land averaged trends are also quite similar (Figure 10d).
3.2 ICON in a bottle
Generation of ICON states at HPX1024 resolution is achieved through a two-stage process in which the macroscale generator of the ICON modality at HPX64 is fed through a patch-based super-resolution module (see Section 2.7.2), which we assess below.
All super-resolved results presented here are produced using a consistent configuration: a maximum noise level ($\sigma_{max}$) of 800, 18 sampling steps, and $S_{churn}$=0.
3.2.1 Super-Resolution Results
Figure 11: (a), (b) Demonstration of super-resolved atmospheric variables outgoing shortwave radiation, and precipitation flux of ICON HPX64 at a selected time during an extreme tropical cyclone event and adjacent midlatitude front. From left to right are the input ICON HPX64, HPX1024 super-resolved by cBottle, and ICON HPX1024. (c), (d) Spherical power spectra plots for 2-meter temperature and precipitation of ICON HPX64 super-resolution inferences across 32 validation samples.
Figure 11 (a), (b) provides a first qualitative assessment of the super-resolution. Here we choose a convenient time sample from the test set, in which a West Pacific tropical cyclone occurred adjacent to a large midlatitude weather system to its northeast. The performance of super-resolution is shown for two field variables – upwelling solar radiation (first row) and precipitation rate (second row). Compared to the ground truth (right column), the results (middle column) are generally encouraging. For the tropical cyclone, appropriate orographic maximization of convection over the interior of Taiwan – not present in the macroscale inputs (left column) – is realistically generated. So too are details of the eyewall structure and surrounding rain bands, albeit with less spatial coherence than in the ground truth. Meanwhile, for the midlatitude weather system, a plausible fine-scale frontal band is generated, extending from the Kiril islands to just east of the Kamchatka peninsula, with consistent horizontal scale and position as the ground truth, across both field variables.
A more quantitative assessment is shown in Figure 11 (c), (d) from the view of spectra. The plots are generated using 32 randomly selected samples spanning from March 6, 2024, to July 23, 2025. The spherical power spectra are computed by averaging the squared magnitudes of the spherical harmonic coefficients over the azimuthal index $m$ and across the temporal dimension. Correct distribution of variance as a function of spatial scale is confirmed in the power spectra for temperature and precipitation.
3.2.2 End-to-end synthesis
We now turn to results from our capstone task of attempting to generate 12M-pixel multivariate ICON states purely from noise, time of year, time of day, and a monthly mean map of SST. The left column of Figure 12 (a), (b), (c), (d) provides a qualitative demonstration of the results (see also Figure 1). Results are shown for an arbitrary subregion spanning the entire North Pacific ocean, across four arbitrary time periods having distinct weather states. It is important to note that the ground truth (corresponding right columns) should not be compared to the model output for any matched spatial positioning of individual weather events, since the task is unconditional generation. That is, the ground truth also has distinct weather states. Rather, what is noteworthy in the comparison is the common modulation of similar microscale weather detail by variably positioned macroscale weather events.
From this view, we find encouraging co-generation of outgoing shortwave radiation and precipitation channels. As in the ground truth, sub-sectors of midlatitude eddies exhibit km-scale organization patterns of non-precipitating shallow cloud (i.e. evident in the shortwave channel but not the precipitation channel). As in the ground truth, zonally elongated macroscale filaments of the intertropical convergence zone are embedded with fine-scale precipitation having its own characteristic length scale of organization. In contrast, midlatitude precipitation is comparatively stratiform – lower intensity and more spatially coherent – with the exception of fronts visible in both brightness and rainfall. While difficult to summarize statistically, such multi-scale, multi-variate spatial associations are readily visible to the trained human eye accustomed to inspecting weather data, and suggest successful synthesis.
Figure 12: (a-d) Demonstration of 2 HPX1024 samples from cascaded cBottle, showing outgoing shortwave radiation, and precipitation flux. Samples from cascaded and ICON are not matched. (e), (f) Spherical power spectra of cascaded cBottle inferences across 64 validation samples, for 2-meter temperature and precipitation.
Spectra analysis again quantitatively confirm this finding (Figures 12 (e), (f)). On the one hand, secondary imperfections are introduced that were not apparent when super-resolution is used in isolation of macro-scale generation (Figures 11 (c), (d)) such as too low occurrence of the very coldest surface temperatures on the planet (less than -60 Centigrade). On the other hand such imperfections are minor such that the results are overall of reasonable enough quality so as to suggest the viability of cascaded super-resolution as a means to generate multivariate ICON states. Figures 12 (e), (f) are generated using 64 randomly selected samples spanning from March 6, 2024, to July 23, 2025.
3.2.3 Video diffusion
Spatio-Temporal Analysis
To evaluate whether the video diffusion model captures realistic atmospheric dynamics through time—beyond just generating independent plausible snapshots—we analyze its ability to reproduce the patterns characteristic of mid-latitude storm track evolution in Figure 13. Following the approach of (Rao et al., 2002), we compute one-point lag correlation maps on mean sea level pressure. The video model generations successfully reproduce both the spatial structure of these patterns and their temporal evolution at comparable propagation speeds to the ICON ground truth. At negative lags, we observe a coherent wave train approaching the reference point from the west, and at positive lags, these correlation continue propagating eastward. The model’s ability to replicate these spatio-temporal patterns indicates the temporal attention mechanism enables the model to capture the underlying physical dynamics and generate coherent trajectories of atmospheric state.
Figure 13: Lagged correlation (left) and regression slope (right) maps for mean sea level pressure computed on unconditional video model generations and ICON ground truth. Maps are computed relative to a seed point at 45S, 85E (green dot), and averaged over 60 non-overlapping 3-day sequences from December to May.
3.2.4 Ablations
Noise schedule
Figure 14 demonstrates the impact of noise schedule choice on diffusion models for atmospheric data. The bottom row shows results using a log-normal schedule ($P_{mean}=-1.2$, $P_{std}=1.2$) typically used for image generation (Karras et al., 2022). This schedule introduces insufficient noise to mask the the dominant diurnal cycle of outgoing shortwave radiation during training. Consequently, the model learns to detect and amplify residual signal patterns from noised inputs rather than using the time-of-day conditioning. However, during inference where starting from pure noise provides no such signal to extract, the model produces physically inconsistent results. Panel (g) shows this most clearly, with a “double sun” artifact with radiation appearing in multiple incorrect locations.
In contrast, the log-uniform schedule (top row) with higher $\sigma_{max}$ provides sufficient noise to completely obscure the diurnal pattern during training, forcing the model to learn the relationship between time-of-day conditioning and the spatial distribution of solar radiation. This yields physically consistent results qualitatively with correctly positioned subsolar points. This visual comparison confirms the earlier theoretical argument (see Subsection2.4) that atmospheric variables, some with planetary scale patterns, require substantially higher noise ceilings than is standard for diffusion to avoid signal leak bias.
Figure 14: Four consecutive frames of reflective visible light (rsut) from a sequence generated by video diffusion models trained with either a log-uniform (top) or log-normal (bottom) noise schedule. Generations are conditioned on the calendar embedding and monthly average SST. Red × marks indicate the subsolar point at each timestep.
Training data size ablation
Figure 15 illustrates the spherical power spectra of inference results from two models trained on datasets of differing sizes: one with 3.75 years of distinct samples (72,360 in total), and another with 4 weeks of distinct samples (1,344 in total), with each week randomly selected from a different season to ensure representative temporal coverage. Both models were trained using 95,000 total samples via repetition to match training iterations.
Across almost all variables, the 4 weeks model closely reproduces the spectral characteristics of the 3.75 years model. This alignment is particularly notable given the drastic reduction in training data diversity. The only substantial deviation appears in the precipitation spectra, where the 4 weeks model exhibits slightly lower zonal energy at high wavenumbers, suggesting a modest underrepresentation of fine-scale variability. These results demonstrate the model’s strong data efficiency: even when trained on a small, seasonally representative subset, it retains robust downscaling performance and preserves key spectral structures across a range of physical variables.
Figure 15: Training data size ablation study: Power spectra from global super-resolution models trained on 3.75 years versus 4 weeks of timesteps.
3.3 Multimodal
In this section we demonstrate how the separate modalities and generative vs. super-resolution modules of cBottle can be composed into novel workflows, providing the beginnings of a foundational versatility.
3.3.1 ICON-Informed Downscaling of ERA5
Although the super-resolution model is trained exclusively on ICON data, it demonstrates zero-shot generalization capability when applied to data from other sources. In this subsection, we assess its skills on ERA5 data regridded to HPX64 resolution. The corrupted radiation channels in ERA5 have been infilled using the method described in Section 3.3.2. The resulting super-resolved outputs are compared against both the 3-km High-Resolution Rapid Refresh (HRRR) dataset over CONUS and the native-resolution ICON model at HPX1024. All datasets used in this evaluation are regridded to a common 3-km Lambert conformal projection over the CONUS region to ensure fair spatial alignment.
Figure 16: (a), (b), (c): Illustration of the super-resolved ERA5 output at 15:00 UTC on December 11, 2018, with radiative channels infilled using the cBottle HPX64 generator. The first three columns correspond to the same timestamp, while the rightmost column (ICON) shows the free-run forecast for 15:00 UTC on December 11, 2024. From top to bottom, the rows represent 10-meter eastward wind velocity, outgoing shortwave radiation, and precipitation. (d), (e) Power spectra of cBottle-resolved ERA5 and various baselines, computed over 10 samples. All datasets are regridded to a Lambert-Conformal projection for comparison. ERA5, cBottle-resolved ERA5, and HRRR share an identical set of evaluation timestamps. ICON HPX64 and ICON HPX1024 use matching times of day and days of the year, but from different years.
Figures 16 (a), (b), (c) presents a qualitative comparison of three representative variables: 10-meter eastward wind velocity, outgoing shortwave radiation, and precipitation rate. The figure shows spatial subsets spanning the Midwest to the Eastern United States, with four columns corresponding to the original ERA5 HPX64 input, the super-resolved ERA5 output generated by our model, the HRRR reference, and the ICON HPX1024 reference. Note that the ICON simulation is a free-running model and is not temporally aligned with the ERA5 or HRRR data; however, it still serves as a useful reference for assessing small-scale structural similarity. While the large-scale patterns in the super-resolved ERA5 resemble those in the original ERA5 and HRRR data, the finer-scale structures more closely resemble those in the native-resolution ICON fields. This suggests that the model effectively emulates high-frequency features in ICON and is capable of generating high-resolution patterns and dynamics that are not explicitly present in the coarse-resolution data.
The zonal-mean spherical power spectra for four representative variables are shown in Figure SI9 (d), (e): 2-meter temperature, outgoing shortwave radiation, and precipitation. Comparisons are made across five datasets: the ERA5 HPX64 input, the super-resolved ERA5 output generated by our model (cBottle), the HRRR reference, ICON at HPX64, and ICON at HPX1024 resolution.
For variables where ERA5 HPX64 and ICON HPX64 exhibit similar spectral characteristics—such as kinetic energy and temperature—the super-resolved outputs from cBottle closely follow the spectral slope and amplitude of the ICON HPX1024 reference, indicating successful reconstruction of high-frequency content. This alignment demonstrates the model’s capacity to synthesize physically plausible fine-scale variability.
For channels where the ERA5 HPX64 input exhibits significant spectral differences from the ICON HPX64 input (precipitation), however, the model’s ability to recover fine-scale variability is limited. While cBottle partially bridges the gap, it underrepresents high-wavenumber energy relative to ICON HPX1024 and HRRR. This highlights the challenge of zero-shot generalization when there is a domain shift in spectral characteristics between the training and evaluation inputs.
To sum up, it is encouraging to observe that the super-resolution model can effectively super-resolve unseen data in a zero-shot setting, while also faithfully emulating the characteristics of its training data. Reducing spectral and distributional discrepancies caused by domain shifts—potentially through the inclusion of more diverse training datasets—may further improve the model’s robustness and enhance its applicability in climate science.
3.3.2 Channel filling
Following Appendix I.2 of Song et al. (2020b), let $\Omega:\mathbb{R}^{C\times\mathbf{d}}\to\mathbb{R}^{C_{\mathcal{K}}\times\mathbf{d}}$ and $\bar{\Omega}:\mathbb{R}^{C\times\mathbf{d}}\to\mathbb{R}^{C_{\mathcal{U}}\times\mathbf{d}}$ denote the slicing operators that extract the known and unknown channels, respectively, where $\mathbf{d}$ denotes an arbitrary set of additional dimensions (e.g., spatial or temporal). The unknown subset at sampling time $t$ is then denoted by $\mathbf{z}_{t}=\bar{\Omega}(\mathbf{x}_{t})$. The conditional score function required to impute the missing channels can be approximated as
$\nabla_{\mathbf{z}}\log p\big{(}\mathbf{z}_{t}\mid\Omega(\mathbf{x}_{0}),\sigma_{t}\big{)}\approx\nabla_{\mathbf{z}}\log p\big{(}[\mathbf{z}_{t};\,\Omega(\mathbf{x}_{t})]\mid\sigma_{t}\big{)},$ (20)
[…]
To evaluate the imputation capability of the model, we focus on infilling radiation channels in the ERA5 and ICON datasets. Specifically, the imputed variables include outgoing shortwave radiation, outgoing longwave radiation, and surface downwelling shortwave radiation. Figure 17 visualizes the infilled outgoing shortwave radiation for both ERA5 and ICON, as produced by the proposed cBottle framework. All ICON images (Target ICON, cBottle-Infill ICON, and cBottle ICON) correspond to the same timestamp, as do all ERA5 samples.
Compared to the respective ground truths, cBottle successfully reconstructs plausible radiation structures for both ERA5 and ICON. The contrast between cBottle-Infill ICON and cBottle ICON is particularly illuminating: although both are conditioned on identical information (timestamp, day of year, and SST), only the result from the infilling workflow (cBottle-Infill ICON) shows accurate reconstruction of the missing radiation channels in alignment with the Target ICON. This highlights that the improvement arises from the imputation technique itself, rather than from conditioning alone.
Furthermore, cBottle with channel filling proves effective in mitigating dataset-specific artifacts. For instance, as seen in the comparison between cBottle-Infill ERA5 and Target ERA5, the model reproduces coherent structures and dynamics while simultaneously removing the streaking artifacts inherent in the original ERA5 fields.
In summary, these results demonstrate that the imputation approach employed by cBottle not only enables the plausible reconstruction of missing physical variables, but also helps mitigate known deficiencies in climate datasets. This underscores the potential of cBottle as a valuable tool for data recovery in atmospheric science, particularly in scenarios where observational records are incomplete, corrupted, or degraded by systematic artifacts.
Figure 17: Imputation of outgoing shortwave radiation using cBottle. cBottle-Infill ICON: ICON data with radiation channels masked and subsequently reconstructed via cBottle. cBottle ICON: ICON data fully generated by cBottle using only SST and timestamp as conditioning. Target ERA5: ERA5 reference with native spatial resolution and known streaking artifacts. cBottle-Infill ERA5: ERA5 data with radiation channels infilled by cBottle.
3.3.3 Bias correction
This approach takes advantage of the latent representation of the diffusion model. To transition a sample from dataset $i$ to $j$ we evaluate the ODE forward to get the latent then reverse it with the $\beta$-denoiser.
It may seem surprising that this yields a realistic sample. Liu et al. (2022) explain that since the flow defined by an ODE is unique, the trajectories between samples cannot cross one another. This property makes the flow map close to an optimal transport map between the distributions.
Figure SI15 shows that cBottle can debias the cloud field of the original ICON data while maintaining temporal coherence across frames. The ICON model prefers a liquid-heavy partitioning between cloud water and ice compared to ERA5(cf Fig. 18). It has nearly the double the liquid water as ERA5. Overall, the location of the clouds in the bias corrected ICON data is similar, though there are some apparent differences.
Overall, the ability to translate between datasets with a single foundation model is an exciting potential application, and we leave more detailed analysis to future studies. Unlike Liu et al. (2022) our model is not trained explicitly to map from one dataset to the other. That approach would require training a separate model for each pair of datasets so does not scale well to many different data sources. By training one model on all potential datasets, we can use the denoiser to define a shared latent space, a desirable property of a foundation model.
Figure 18: Bias correction of the total cloud ice.
4 Conclusions
We have demonstrated that masked denoising diffusion is a feasible approach to generative modeling of the atmosphere. The two-stage cascade comprises: (i) a coarse-resolution multi-modal climate data generator; and (ii) a cascaded patch-based super-resolution module. Together, these have the potential to enhance: (i) the way km-scale climate data is accessed; and (ii) the practice of climate informatics.
In the first stage, provided only 200 kb of inputs comprising two scalars (time of year, time of day) and a 50,000 pixel map of monthly mean SST data, cBottle synthesizes 9 MB of global multivariate samples for 45 channels that emulate the macroscale (HPX64) climatological properties of ERA5 or ICON. Our analysis suggests the generated samples exhibit the same mean state and variance from diurnal, synoptic, seasonal and interannual modes of variability as the ground truth. In the second stage, given such HPX64 fields, cBottle’s cascaded super-resolution module generates HPX1024 states for 12 channels, totaling 600 MB. This stage exploits a patch-based approach to handle scaling to the global (12.5-million pixel) domain. We have shown that the super-resolved states have accurate spectra, can be trained with remarkable sample efficiency – as little as four weeks of ICON HPX1024 output – and that secondary patch boundary artifacts can be mostly controlled using simple overlap techniques. Either stage can be used independently, for instance, to generate coarse-resolution (HPX64) climate samples, or to super-resolve existing HPX64 climate data to HPX1024 resolution with ICON-learnt high-resolution characteristics.
The resulting end-to-end system – ICON in a bottle — offers 3000x compression ratio per sample over the raw data (256x compression ratio per channel for 12 channels). We suggest this technology can thus greatly reduce the bottleneck that storage, I/O, and data movement presents in modern informatics, which has become especially extreme with the advent of global cloud resolving simulation output. This end-to-end system enables users to access realistic km-scale data not by high-latency retrieval of stored data but by low-latency generation of that data, on demand, from a generative ML model.
Moreover, the modular and multi-modal nature of cBottle demonstrates new possibilities characteristic of AI foundation modeling that are as yet unfamiliar to the climate science community under the prevailing autorogressive paradigm for emulation. cBottle exploits climate information across the multiple modalities it is trained on, and transfers learnings across them. For instance, we have shown how high-resolution textures of ICON-learnt cloud-scale organization can be in-painted onto the appropriate sectors of synoptic systems of the macroscale ERA5 modality. We have also shown how channels missing from one-modality can be readily transferred from another, by in-painting the missing shortwave radiation field in the ERA5 modality exploiting what was learnt from ICON. The narrow lesson is that synthetic output from high-resolution simulations can be used more flexibly than it might otherwise appear. The broader lesson is more profound: While our proof of concept has focused on just two data modalities for simplicity, we hope it inspires extensions across the many complementary observing and simulating systems of the Earth System, where such cross-modal learning, in-filling, mode-transfer, and bias-correction becomes increasingly useful for climate science and society. We expect these extensions to help realize a climate foundation model at the km-scale.
In terms of methodology, we have introduced diffusion on the HEALPix grid and highlight the importance of the noise schedule when training denoising diffusion models on scientific data. In particular we show that the defaults most often used in natural image generation literature are not capable of generating extremely large-scale modes like the seasonal and diurnal cycles. Fixing this required two main innovations: i) increasing the range of noise levels sampled during training and inference; and ii) adding appropriate time and position embedding. Specifically, we propose tuning the noise schedule based on spectral analysis of the data, and show that the noise schedule should scale with $\sqrt{n}$ where $n$ is the size of the largest-scales modes in number of scalars (pixels $\times$ channels). This perspective should be helpful to anyone training diffusion models on scientific data.
Additional limitations of our preliminary work are important to note. Non-stationary trends are not optimally represented in the current version of cBottle; while we have demonstrated that 90th percentile 500 hPa geopotential height exceedance trends can be emulated, analogous trends in surface temperature are currently elusive. The mean climate – while reasonable – is imperfect, including 4-10% biases in the zonal mean precipitation rate. Another important limitation is that our image generation framework cannot diagnose the duration of events or commonly used metrics like the return time of an event. Future work will need to add temporal coherence and extend the system to enable the generation of spatio-temporally coherent time series of phenomena.
Appendix A Details for Z500-90p analysis
The methodology mirrors the tx90p analysis feature defined by Seneviratne et al. [2021] (See Fig. 11.9). First, percentiles of Z500 at 0 UTC are computed for each day of the year and spatial location using a 14-day rolling window. Specifically, we use the climate data operators [Schulzweida, 2023] ydrunpctl command with a 14-day window, and compute percentiles over 1990-2010.
Then, the fraction of samples exceeding this daily percentile climatology is computed for each year separately. The regression coefficients in Figure 10e,f are computed by fitting a regression line with least squares over the entire period we had ERA5 data for 1980-2018 (inclusive).
Appendix B Diurnal Cycle: Details
The diurnal amplitude is defined as
$S:=\frac{1}{366}\sum_{d}\max_{h}\overline{T}(d,h)-\min_{h}\overline{T}(d,h).$
$\overline{T}(d,h)$ is the average of $T$ as a function of day of year $d$ and hour of day $h$. For the SNR computation, the noise is given by
$N:=\frac{1}{366}\sum_{d}\frac{1}{24}\sum_{h}s_{T}^{2}(d,h),$
where $s_{T}^{2}$ is the conditional sample variance. In this computation, $s_{T}^{2}$ and $\overline{T}$ are computed over 1990–2018 (inclusive). Then, the $SNR:=S/N$.
Figure SI1: Diurnal cycle diagnostics for surface air temperature. (top row) Amplitude of the diurnal cycle. (bottom row) signal to noise ratio.
Appendix C Architecture: Details
Figure SI2: Position and calendar embedding.
Figure SI3: Schematic showing the UNet architecture for spherical geometry with HEALPix grids. Each convolution is applied on HEALPix faces padded as in [Karlbauer et al., 2024].
Figure SI4: Training workflow of the multi-diffusion framework. Global maps are decomposed into patches for parallel training of local denoisers conditioned on spatially aligned context.
Figure SI5: Illustration of the aggregated local denoiser replacing the global denoiser in the sampling algorithm.
Appendix D Datasets: Details
Appendix E Pseudo-code for masking in the loss function
D_yn = net(where(mask, y + n, 0), sigma)
loss = weight * ((D_yn - where(mask, y, 0)) ** 2)
loss = where(mask, loss, 0) / mask.mean()
Appendix F Super-resolution results: Details
Figure SI6: Spherical power spectra and distribution plots of ICON HPX64 super-resolution inferences across 32 validation samples. (a), (c): power spectra for 2-meter temperature and precipitation. (b), (d): log-scale distributions of 2-meter temperature and precipitation.
Super-resolution probability distribution
Figure SI6 shows the probability distributions of surface air temperature and precipitation aggregating over all temporal and spatial locations. The super-resolution (red line) captures the shape of temperature and precipitation PDFs (Figure SI6a,c) including the correct power law regulating the heavy tail of low-likelihood extreme precipitation occurrence (Figure SI6c).).
Cascaded generations probability distribution
Figure SI7 shows the probability distributions computed using 64 randomly selected samples spanning from March 6, 2024, to July 23, 2025. Overall, the distributions match well, though the cascaded model struggles to produce the coldest temperatures (less than -60 deg C). The super-resolution did not have this problem, which suggests that the bias likely comes from the coarse-resolution generations in this cascaded setup.
Figure SI7: Spherical power spectra and distribution plots of cascaded cBottle inferences across 64 validation samples. (a), (c): power spectra for 2-meter temperature and precipitation. (b), (d): log-scale distributions of 2-meter temperature and precipitation. Samples from cascaded and ICON are not matched.
Multidiffusion Parameter Ablation
The overlap size in multidiffusion sampling is a critical factor influencing both computational efficiency and output quality. Specifically, it determines the number of denoiser evaluations required: fewer evaluations reduce computational cost but can compromise visual consistency due to patch independence. A zero-overlap configuration, while minimizing runtime, often introduces visible artifacts at patch boundaries. Therefore, selecting an appropriate overlap size is essential to balancing reconstruction fidelity and efficiency.
To investigate this trade-off, we evaluate various overlap sizes using the median absolute horizontal Sobel gradient of surface temperature, as shown in Figure SI8. This metric serves as a proxy for boundary artifacts, where higher gradient magnitudes indicate sharper discontinuities. To generate the plot, we aggregate all faces from 16 randomly sampled outputs. A Sobel-X filter is applied to the surface temperature field, and the median of the resulting gradient magnitudes is computed across the face, sample, and vertical dimensions.
As expected, the 0-pixel overlap configuration exhibits pronounced, periodic spikes at 128-pixel intervals, indicating strong discontinuities at patch edges. The 64-pixel overlap shows secondary spikes at 64-pixel intervals. The 4-pixel overlap yields slightly stronger artifacts than the 32-pixel case, suggesting that very narrow overlaps may be insufficient to effectively smooth patch boundaries.
Figure SI8: Severity of patch artifacts across different overlapping pixel sizes, visualized using the mean absolute Sobel-X gradient of surface temperature field as a function of horizontal pixel position.
Figure SI9: Emulation of native-grid ICON with ERA5: Power spectra of cBottle-resolved ERA5 and various baselines, computed over 10 samples. All datasets are regridded to a Lambert-Conformal projection for comparison. ERA5, cBottle-resolved ERA5, and HRRR share an identical set of evaluation timestamps. ICON HPX64 and ICON HPX1024 use matching times of day and days of the year, but from different years.
Appendix G Further climate diagnostics
Figure SI10: Seasonal cycle for (JJA- DJF) for ERA5 and cBottle. Panels are shown for Z500, surface air temperature, precipitation, and sea ice concentration.
G.1 Zonal Mean Precipitation
Figure SI11 examines mean precipitation in cBottle ERA5 and its relationship with latitude computed. Overall cBottle ERA5 captures the existence and correct positioning of local maxima associated with midlatitude storm tracks and intertropical convergence zones straddling the equator. Local imperfections include positive biases in the northern (+4%) and southern (+11%) tropical rain bands, and as much as 11% underestimation of midlatitude rainfall. These differences can in part be explained partially by sampling error.
Figure SI11: Zonal Average of precipitation ICON, ERA5, and their emulators. 95% confidence intervals for the validation set means are estimated by bootstrapping over monthly averages. Only data from the validation periods for ICON and ERA5 are used.
G.2 Power-spectral fidelity of the coarse-resolution generator
Figure SI12: (left) Spherical power spectra of versus the truth. Averages over 100 random samples from the test period are shown. The dotted lines indicate the range of noise levels used for the sampling $\sigma_{\{min,max\}}^{2}\frac{4\pi}{n_{pix}}$. The factor of $\frac{4\pi}{n_{pix}}$ scales the pixel-wise variance to the same units as the power spectra, which is variance per surface area of the unit sphere.
Figure SI13: PDF of instantaneous precipitation. Computed for same samples as in SI12.
Figure SI12 shows that the spectra of the generated coarse samples are consistent with their ICON and ERA5 ground truth. For some fields, the ICON data shows significantly more power than ERA5 at large wave numbers, but we have not confirmed if this comes from the aliasing of smaller scale modes due to the block averaging used to coarsen the native HPX1024 outputs for ICON or an actual difference.
G.3 Submonthly variance
Figure SI14: Sub-monthly standard-deviation (variance from the monthly mean averaged over all months in the test period), surface air temperature, and precipitation. Panels are shown for the ground truth and cBottle for ICON and ERA5.
Figure SI14 maps summarize the pointwise variance of samples after removing the seasonal cycle, i.e. anomalized with respect to their monthly climatology (Figure SI14). This summarizes dispersion produced by a variety of sources including diurnal and synoptic variability. Variance of 500 hPa geopotential height (top) is maximized over the Southern Ocean, North Atlantic Sea and northeast Pacific due to midlatitude weather systems, but avoids the tropics due to the weak temperature gradient constraint there [Sobel et al., 2001]. Midlatitude rainfall variance occurs over the storm tracks including the Gulf Stream and Kuroshio currents. Rainfall variance is especially high within the intertropical convergence zones. In the special case of surface temperature, diurnal variability leads to a distinct land-sea contrast – a result of solar heating causing strong day-night temperature cycles over land, where surface specific heat capacity is relatively low.
The main point of Figure SI14 is that cBottle-ERA5 and cBottle-ICON faithfully generate all of these pointwise variance features summarizing diurnal-to-synoptic variability – and with sufficient precision to detect differences between them, such as the “double ITCZ” – which manifests in cBottle-ICON and cBottle-ICON-Target as a zonal band of excessive dispersion in tropical rainfall in the eastern Pacific just south of the equator.
Appendix H Supplementary Figures
Figure SI15: Bias correction of the total cloud liquid water. (top) original ICON data and (bottom) bias corrected.

### v2

@@
 Abstract
-AI emulators offer a path to compressing, boosting limited ensembles, and improving the latency of interacting with petabyte-scale climate prediction data.
-However, prevailing auto-regressive paradigms offer limited flexibility, and are challenging to train on climate time horizons due to drifts, instabilities and component-coupling challenges.
-Conditionally generative models offer an appealing alternative.
-In this context we demonstrate a generative diffusion-based framework—Climate in a Bottle (cBottle)—for emulating global km-scale climate simulations and reanalysis on the equal-area HEALPix grid.
-cBottle consists of two model stages: a globally-trained coarse-resolution image generator that generates 100km (50k-pixel) fields given monthly average sea surface temperatures and solar conditioning, followed by a locally-trained 16x super-resolution stage that generates 5km (12.5M-pixel) fields; global super-resolution is made affordable using an overlapping patch-based multi-diffusion.
-Overall, cBottle shows promise as an emulator across a battery of climate model diagnostics, including diurnal-to-seasonal scale variability, large-scale modes of variability, tropical cyclone statistics, and trends of climate change and weather extremes.
-Moreover, cBottle is a step towards a foundation model, by bridging multiple data modalities (reanalysis and simulation) with corresponding utility beyond emulation to tasks such as zero-shot bias correction, climate downscaling, and channel in-filling.
-The code is available at https://github.com/NVlabs/cBottle.
+Climate modeling is reaching unprecedented resolution, producing petabytes of climate data [1].
+AI climate model emulators offer a path to computationally cheap analysis of this data, enabling new scientific insight and scenario planning [2].
+Recent advances have shown promise in faithfully emulating climate data [3, 4, 5, 6].
+However, prevailing auto-regressive paradigms are challenging to train on climate time horizons due to drifts, instabilities and component-coupling challenges, they are hard to scale to high resolution, and pose the challenge of having to sift through troves of output to identify rare climate extremes of outsize interest.
+Here, we present Climate in a Bottle (cBottle), a generative diffusion-based framework [7, 8, 9] emulating global 5 km climate simulations and reanalysis on the HEALPix grid.
+cBottle directly samples from the full distribution of atmospheric states without requiring a previous time step and therefore auto-regressive rollout, as well as the first to reach this global resolution (12.5M-pixel).
+cBottle consists of two model stages [10]: a global coarse-resolution generator conditioned on sea surface temperatures and solar position, followed by a patch-based 16x super-resolution stage.
+cBottle passes a battery of climate model tests, including diurnal-to-seasonal scale variability, large-scale modes of variability, tropical cyclone statistics, and trends of climate change and weather extremes.
+Moreover, cBottle is a step towards a foundation model: It bridges multiple data modalities (reanalysis and simulation), and enables zero-shot bias correction, climate downscaling, and data in-filling.
+It also enables a new form of human interactivity by leveraging guided diffusion [8].
+For example, users can request samples of extreme weather located over regions of interest – we demonstrate this by training a tropical cyclone (TC) classifier alongside the generator, guiding towards TC states, and obtaining physically credible samples.
+We anticipate this to kickstart developments of guidance methods for a wide array of user queries and new ways of interacting with climate data.
 1 Introduction
-Knowing more, and planning better for the consequences of climate change are driving tremendous advances in the realms of climate simulation and Earth observation.
-With exascale computing and energy-efficient computational systems, petabytes of kilometer-scale climate data are being generated.
-Digital-twin efforts like the Destination Earth initiative (Bauer et al., 2021) and the Earth Virtualization Engines (EVE) project (Stevens et al., 2024) envision a world where high-fidelity global climate projections with local granularity are available to all with a degree of access and interactivity that provides a scientifically grounded basis for action.
-Further, digital twins aim to efficiently expose the data to new methods of analysis and proactive information production that will enable stakeholders to construct and interact with their own climate scenarios.
+Policymakers, infrastructure planners, the insurance industry and the public rely on accurate and high-resolution information about the present and possible future climate.
+This need is driving tremendous advances in climate simulation and Earth observation technology.
+With exascale computing, energy-efficient computational systems now generate petabytes of kilometer-scale climate data.
+Digital-twin efforts like the Destination Earth initiative [1] and the Earth Virtualization Engines (EVE) project [2] envision such high-fidelity physical climate projections as underpinning next-generation climate informatics infrastructure, which should eventually enable a degree of access and interactivity to allow widespread climate-informed decision making.
 Easy access and low-latency interactivity, however, are nearly impossible with the current paradigm of petabyte-scale datasets generated and hosted by a handful of climate centers.
-Even if the access problem is unthrottled by moving data around, interacting with petabytes of data is virtually impossible.
-Generative artificial intelligence (AI) is poised to disrupt this paradigm by fully leveraging state-of-the-art architectures and algorithms optimized for performance on accelerated computing hardware.
+Even if the access problem is unthrottled by moving data, interacting with petabytes of data is computationally expensive.
+Recent advances in AI to replace the simulation task do not solve this challenge.
+While AI has proved efficient and faithful at emulating models and climate data evaluated across a range of metrics on a wide variety of tasks [3, 4, 5, 6], most of these developments have resorted to the auto-regressive modeling paradigm, much like numerical weather and climate models that step forward in time evolving the state of the system modeled by governing systems of equations.
+Such systems have not been scaled to the km-scale global resolution envisioned by EVE.
+Even if this were possible it would not solve the informatics challenge of nimbly guiding users to samples or statistics of interest, such as rare extreme weather events impacting a particular region.
+That is, autoregressive AI emulators share a common challenge with physical climate models – the need to sift through troves of simulation output to uncover information of interest.
+We propose that generative artificial intelligence (AI) for sampling offers a solution to both the data volume and interactivity problems by fully leveraging state-of-the-art algorithms optimized for performance on accelerated computing hardware.
 For our purposes, we define a generative foundation model as a self-supervised climate simulator capable of producing realistic-looking samples that are climate-faithful measured by a set of diagnostics, and can be used to examine what-if scenarios not explicitly built into the model.
-Of course, climate science has long had such foundation models in the form of physics-based general circulation models.
-However, advances in AI are already fundamentally changing how weather and climate are simulated due to their ease of use, enhanced computational performance, and improved accuracy.
-Specifically, emulators of weather have proven to be skillful and orders-of-magnitude faster than traditional numerical simulation-based approaches to simulating the weather across spatial and temporal scales (Pathak et al., 2022; Lam et al., 2023; Bi et al., 2023; Price et al., 2024; Bodnar et al., 2024).
-Similarly, AI is proving to be efficient and faithful at emulating climate models and climate data, evaluated across a range of metrics on a wide variety of tasks (Watt-Meyer et al., 2024; Chapman et al., 2025; Cresswell-Clay et al., 2024; Watson-Parris et al., 2022).
-A majority of these developments have resorted to the auto-regressive modeling paradigm, much like numerical weather and climate models that step forward in time evolving the state of the system modeled by governing systems of equations.
-Masked training and representation learning have also been proposed (Lessig et al., 2023).
-Diffusion modeling (Ho et al., 2020; Song et al., 2020a; Karras et al., 2022) offers an alternative to the auto-regressive time-stepping approach by learning—and sampling—from a distribution of states.
-This is especially attractive for rapid analyses and informatics that do not require the full temporal evolution of the system.
-By the same token, they are not susceptible to instabilities or drifts (i.e., hallucinations) inherent to auto-regressive models, though admittedly many auto-regressive emulators perform surprisingly well for large numbers of steps (Watt-Meyer et al., 2024; Cresswell-Clay et al., 2024).
-Further, diffusion models can be controlled by the user in a variety of ways, using guidance (Song et al., 2020a) to generate conditional distributions.
-Generative diffusion modeling has had tremendous impact across disciplines ranging from computer vision to science, and enjoys widespread adoption across the industry.
-Notably, it has been used in weather and climate for ensemble generation and uncertainty quantification (Li et al., 2024), probabilistic forecasting (Price et al., 2024), super-resolution (or downscaling) (Mardani et al., 2025), and data-assimilation (Rozet and Louppe, 2023; Manshausen et al., 2024).
+Specifically, diffusion modeling [7, 8, 9] offers an alternative to the auto-regressive time-stepping approach by learning—and sampling—from a distribution of states.
+This is especially attractive for rapid analyses that do not require the full temporal evolution of the system.
+By the same token, they are not susceptible to instabilities or drifts inherent to auto-regressive models, though admittedly many auto-regressive emulators perform surprisingly well for large numbers of steps [3, 5].
+Further, diffusion models can be controlled by the user in a variety of ways, using guidance [8] to generate conditional distributions.
+Generative diffusion modeling has been used in weather and climate for ensemble generation and uncertainty quantification [11], probabilistic forecasting [12], super-resolution (or downscaling) [13], and data-assimilation [14, 15].
 A less explored yet promising direction is its application towards learning a compressed representation of massive climate datasets that generates, on demand, climate states steered by user-specified conditions and scenarios.
-Early examples included training deterministic emulators to predict the mean-state response of climate models to various forcings (Watson-Parris et al., 2022).
-More recent efforts have extended this approach to stochastic emulators in the diffusion framework (Bassetti et al., 2024; Quilcaille et al., 2022) as well as traditional dimensionality reduction approaches (Wang et al., 2025).
-But these approaches have been limited to a small number of channels (typically surface temperature) at a coarse spatial resolution.
-We develop a generative diffusion-based framework that we call “Climate in a bottle” (cBottle) to emulate multimodal climate data.
-cBottle uses a cascaded approach (Ho et al., 2021) to generating km-scale samples: a coarse-resolution generator conditioned on user-specified inputs (time of day, time of year, and monthly-averaged sea surface temperature (SST)) and a 16x super-resolution module.
+Early examples included training deterministic emulators to predict the mean-state response of climate models to various forcings [6].
+More recent efforts have extended this approach to stochastic emulators in the diffusion framework [16, 17] as well as traditional dimensionality reduction approaches [18].
+But these approaches have been limited to a small number of channels (typically surface temperature).
+And, as with current AI simulators, none have been scaled to the ambitious km-scale resolution of next-generation climate predictions such as proposed by EVE.
+In this context, we present “Climate in a bottle” (cBottle), a first-of-its-kind generative diffusion-based framework, faithfully emulating multimodal high-resolution climate data.
+cBottle offers pathways towards: (i) solving the data and informatics challenges of km-scale climate data; and (ii) building a generative foundation model for an interactive Earth digital twin.
+2 cBottle
+cBottle generates km-scale samples from the climate distribution using a cascaded approach [10] as shown in Fig.
+1b): A coarse-resolution generator conditioned on user-specified inputs (time of day, time of year, and monthly-averaged sea surface temperature (SST)) and a 16x super-resolution module capable of generating over ten million pixels globally (see Fig.
+2 for a visualization of the high-resolution output) for dozens of channels (see Table 2).
 Both steps use generative diffusion models, and either can be used independently, enabling generation of coarse-resolution samples and/or super-resolution to km-scale.
-cBottle has a few key properties:
+Special care was required to train the model to produce the largest modes of variability without over-fitting to them.
+See Methods for details of the model formulation.
+Briefly, cBottle’s key properties are:
   1. 1.
-    Multi-modality and domain transfer: Bridges ERA5 and ICON km-scale datasets, enables generation of km-scale climate states that exploits the information content contained across ERA5 and ICON
+    Multi-modality and domain transfer: Trained on reanalysis data and 5-km global atmospheric simulations with a convection-resolving climate model, enables generation of km-scale climate states that exploits the information content contained in both modalities.
+Specifically, we use the ECMWF Reanalysis v5 (ERA5) [19] —a 40-year observationally-constrained dataset available at coarse resolution, and a 5-year simulation with the ICON global 5-km model [20].
   2. 2.
-    Versatility across a few useful tasks: Bias-correction, channel in-filling, downscaling
+    Versatility across several useful tasks: removing bias in climate simulations, diagnosing some fields (e.g.
+precipitation) given others (e.g.
+winds), or down-scaling coarse-resolution climate data to km-scale.
   3. 3.
-    Climate faithfulness: Generates realistic climatology, climate variability across timescales, large-scale modes, and extreme weather statistics
+    Climate faithfulness: Generates realistic climatology, climate variability across timescales, large-scale modes, and extreme weather statistics,
   4. 4.
-    Data efficiency: Can be trained on as little as 4 weeks of simulation output
+    Data efficiency: The super-resolution model can be trained on as little as 4 weeks of simulation output, reducing the need for very long simulations with expensive convection-resolving models.
   5. 5.
-    Extreme compression: Encapsulates km-scale climate model and reanalysis outputs into a few GB of neural network weights, and offers 256x compression ratio per channel
+    Extreme distillation: Encapsulates km-scale climate model and reanalysis outputs into a few GB of neural network weights, and offers 256x compression ratio per channel.
   6. 6.
-    Speed and scale: Low-latency generation of km-scale samples and scalability to higher resolutions and more modalities
-These properties of cBottle offer potential pathways towards: (i) solving the data and informatics challenges of km-scale climate data; and (ii) building a generative foundation model for an interactive digital twin.
-Mean sea-level pressure (generated) (ground truth)
+    Speed and scale: Low-latency generation of km-scale samples and scalability to higher resolutions and more modalities.
+  7. 7.
+    Interactivity: Can be guided flexibly to yield extreme events like Tropical Cyclones in a specific location on demand.
+Figure 1: Climate in a Bottle: A generative foundation model for kilometer-scale climate simulation.
+(a) Training phase: Petabytes of high-resolution climate data from ICON global cloud resolving simulations and ERA5 reanalysis are compressed into a compact neural network representation ($\sim$ few GB).
+(b) Inference phase: The trained model generates realistic 5km-resolution climate states from minimal conditional inputs (time, location, sea surface temperature) through a two-stage process: Coarse global generation (100km) followed by patch-based super-resolution.
+(c) Applications: The foundation model enables diverse climate science applications including interactive what-if scenarios, bias correction between datasets, channel in-filling for missing variables, statistical downscaling, multimodal learning across data sources, and high-fidelity climate emulation—all achievable in minutes rather than hours and accessible without supercomputing resources.
+Surface air temperature (tas) Ground truth Generated
 Refer to caption Refer to caption
-Surface air temperature (generated) (ground truth)
+Infrared (rlut) Ground truth Generated
 Refer to caption Refer to caption
-Infrared(generated) (ground truth)
-Refer to caption Refer to caption
-Visible light (generated) (ground truth)
-Refer to caption Refer to caption
-Surface precipitation (generated) (ground truth)
-Refer to caption Refer to caption
-Figure 1: High resolution (HPX1024) samples from the ICON data (right) and synthesized by cBottle given sea surface temperature and time conditionings (left).
-Note these are not identically paired.
+Figure 2: High resolution (HPX1024) samples from the ICON data (left) and synthesized by cBottle given sea surface temperature and time conditionings (right).
+Note these are not identically paired but use the same input SST condition for 2025-03-16T15:00:00.
 For each field the same colormap is used, but to highlight the visual similarity we do not include color-bars.
-More quantiative comparisons will follow.
+More quantitative comparisons will follow.
+See Figure SI1 for more fields.
+3 Results
+Below, the acronyms HPX64 (HPX1024) refer to the low-resolution 100-km (high-resolution 5-km) components of cBottle (see Section 5.2).
+3.1 ICON in a bottle
+Generation of ICON states at HPX1024 resolution – nearly 13M pixels, a formidable task – is achieved through a two-stage process in which the macroscale generator of the ICON modality at HPX64 is fed through a patch-based super-resolution module (see Methods), which we validate below.
+Figure 2 shows a comparison of (unpaired) samples of ICON and cBottle end-to-end-generated surface temperature and infrared radiation—these are visually similar, and the generated samples show what fine-grained detail cBottle can achieve.
+While we have chosen to highlight the end-to-end synthesis task – where high-dimensional ICON weather states are generated from low-dimensional climate-controlling inputs – it is not the only use-case.
+The compression application, where coarsened ICON output are input to the super-resolution model also is relevant as a means of compressing large global storm-resolving simulations.
+Validation of super-resolution in this context is provided in Section B.4.1 where we show cBottle successfully super-resolves the different length scales of ICON’s convective organization from shallow cloud systems influencing its shortwave radiation channel to deep tropical convection clusters in its precipitation channel (Fig.
+SI2), and provide quantitative validation of PDFs and spectra (Fig.
+SI8).
+3.2 ERA5 in a Bottle
+Realistic conditional samples
+To be useful as a climate sampler, cBottle must demonstrate realistic variability across independent samples from phenomena operating on diurnal, synoptic, seasonal, and interannual time scales that, when averaged, produce a realistic climatology.
+To validate this we perform an Atmosphere Model Intercomparison Project (AMIP) control experiment with prescribed monthly-averaged SSTs [21, 22].
+Samples from AMIP experiments are typically produced by integrating physical climate models (or autoregressive AI simulators) forward in time subject to SST boundary conditions.
+In cBottle, analogous samples can be obtained by simply generating independent samples every hour from 1940 to 2021 (inclusive)—a large-scale inference, requiring 1024$\sim$ 1024∼ 1024 GPU-hours on the H100 GPU, but embarrassingly parallel since each atmospheric state (timestep) is completely independent and only shares the same conditioning.
+The following common diagnostics assess the quality of the generated samples and statistics. Additional analyses are presented in Section B.6.
+Figure 3: Fidelity to conditioning.
+Amplitude of the diurnal cycle of precipitation for (a) cBottle and (b) ERA5.
+DJF-JJA Arctic sea ice concentration computed over the test period for (c) cBottle and (d) ERA5.
+(e) cBottle’s secular temperature trend during the AMIP period compared to ERA5.
+(f) User-provided hurricane conditioning.
+Minima of pressure shown in blue are linked to TCs generated by cBottle using classifier guidance.
+The input to the inference is the location of the cyclones, shown as turquoise boxes.
+Diurnal Cycle
+In Figure 3a,b) we compare the amplitude of the diurnal cycle of precipitation in cBottle with ERA5.
+Although the amplitude is too large in cBottle, there is a realistic spatial structure of the precipitation diurnal cycle amplitude including spatial details interior to tropical continents – and even the halo that surrounds the Maritime Continent from the emanating diurnally phased convection propagation there [23].
+Section B.1 provides additional analysis of the diurnal band including the signal-to-noise ratio which confirms cBottle reproduces even small-amplitude signals like the locally detectable subtropical stratocumulus drizzle diurnal cycles (cf.
+Fig.
+SI5(d)) and the appropriate land-sea contrast in the surface temperature diurnal cycle (Fig.
+SI4).
+Seasonal cycle and Mean Climate
+Figure 3c) shows that cBottle correctly generates the seasonal evolution of Arctic sea ice concentration compared to ERA5 in d) (See Fig.
+SI14 for the concentrations by season).
+During boreal summer (JJA) the sea ice margin retreats and its concentration is reduced locally over the Beaufort Sea, Chukchi Sea and Bering Sea – poleward of Alaska and Siberia.
+These results are evidence that the time-of-year and/or SST-conditioning in cBottle work as intended.
+The seasonal cycle is similarly realistic for other fields (see Figure SI15).
+The annual mean climate of the cBottle generated precipitation samples also matches the ERA5 and ICON data (see Figure SI16).
+Northern Annular Mode and SST conditioning
+We next assess a planetary scale mode of internal atmospheric variability.
+In the northern hemispheric midlatitudes, the Northern Annular Mode (NAM) is the leading empirical orthogonal function (EOF), which accounts for one-fifth of overall variance in the monthly anomalies of 1000 hPa geopotential height computed from November to April [24].
+Figure 4d shows that the daily variability of cBottle’s NAM index is plausible though slightly weak—with a root-mean-square amplitude in 2018 of 0.84 versus 1.02 in ERA5.
+Moreover, its low frequency envelope is in phase with inter-annual fluctuations in the ERA5 NAM state, which is more variable during the winter months, suggesting learnt associations between the generated NAM state and the SST and day of year inputs to cBottle.
+Appropriate conditioning on SST is further confirmed by analyzing the El Nino Southern Oscillation (ENSO; see SI Section B.2 and Fig.
+SI6).
+We note that this success in NAM generation depended on using our proposed regularized ensemble of experts approach (see Section A.3.2) to avoid severe over-fitting to the training data that is otherwise encountered when a single checkpoint trained to 9.8M images is naively used instead.
+Figure 4b demonstrates such overfitting as unrealistic temporal coherence between ERA5 and cBottle in the test set; in contrast, our regularized model has a better fair continuous-ranked-probability score (CRPS) [25] in the validation set and a smaller discrepancy between training and validation scores.
+Figure 4: Internal variability.
+(a,c) Tropical cyclone occurrence probability derived from ERA5 and cBottle (1980–2018), presented as heatmaps normalized per $km^{2}$.
+(b,d) The Northern Annular Mode (NAM) index for cBottle samples with overfitting (b) and large-noise-level regularization (d).
+The NAM indices are computed by projecting the daily Z1000 anomalies onto the NAM loading pattern (Fig SI13.
+For convenience, the CRPS is computed by comparing the 0 UTC from the ground truth index to the 24-member ensemble of cBottle samples taken from ever hour of the same day.
+3.3 Interactivity
+cBottle being a diffusion model enables the unique capability of oversampling parts of the distribution that are particularly relevant, such as extreme events.
+We show this using the example of Tropical Cyclones (TCs), which are generated with a realistic climatology in vast AMIP samples (Fig.
+4a,b) but nonetheless rare in individual generations.
+This relates to a principal motivation of cBottle to enable new interactive experiences in climate informatics, such as sidestepping the need for a user to sift through vast datasets to find e.g.
+extreme events of interest over locations of concern.
+To demonstrate user-informed sampling from the distribution of TCs, we train a simple classifier on historical observations from the International Best Track Archive for Climate Stewardship (IBTrACS) [26].
+We use the classifier for guidance during inference [27] (see section 5.7 for details), enabling us to steer the output state to be compatible with an input map of desired Tropical Cyclone locations.
+Fig.
+3f shows a sample generated using this technique, with three TCs successfully generated in the locations requested in the conditioning input.
+This demonstrates that an AI climate state sampler can generate extreme events on-demand.
+In our experiments, imposing such conditions in certain locations did not necessarily lead to reduced variability elsewhere, and channels are correlated in a realistic way.
+Unphysical conditioning, such as guiding towards a TC on the equator, is possible but is somewhat impeded by the model (see SI section B.3).
+3.4 Multi-modal inference
+We demonstrate how the separate modalities and generative vs.
+super-resolution modules of cBottle can be composed into novel workflows, providing the beginnings of a foundational versatility.
+We show three examples tested on a real-world atmospheric-river that occurred during the validation period [28].
+Channel infilling
+When the channel of interest is not only biased, but damaged, or not available at all in one of the two datasets, we can use cBottle to infill it based on the other dataset.
+In Fig.
+5 e–g) we show infilled radiation channels for the ERA5 sample in a, b), channels which in the original ERA5 data show streaking artifacts and were excluded from the training data.
+Compared to the ERA5 ground truths in Fig.
+5 a, b), cBottle successfully reconstructs plausible ICON fields.
+This feature can be used to diagnose any given field given arbitrary subsets of the other fields so it removes the need to train combinatorially many diagnostic models to predict certain fields (e.g.
+precipitation) given others (e.g.
+free-tropospheric states).
+Translating between modalities
+The ICON model prefers a liquid-heavy partitioning between cloud water and ice compared to ERA5 (cf Fig.
+5a, c).
+It has nearly double the liquid water as ERA5.
+cBottle can translate from such an ERA5 sample (a, b) to an ICON sample (c, d) (see Methods).
+Overall, the location of the macroscale weather phenomena in the translated data is similar, though there are some apparent differences.
+This ability to translate between datasets with a single foundation model is an exciting potential application, and we leave more detailed analysis to future studies.
+Downscaling
+To downscale ERA5 data to higher resolution, we first translate the ERA5 sample to the ICON distribution using the coarse-resolution model, and then apply the super-resolution model.
+The results in Fig.5h shows that the precipitation is concentrated in the same regions as Fig.
+5a,b.
+In summary, these results demonstrate that cBottle can plausibly reconstruct physical variables and unresolved features throughout the globe.
+This underscores the potential of cBottle as a valuable tool for data recovery in atmospheric science, particularly in scenarios where observational records are incomplete, corrupted, or degraded by systematic artifacts or high-resolution dynamical-downscaling simulations have not been performed.
+Figure 5: Multi-modal inference methods.
+The original ERA5 data for vertically integrated liquid water (CLLVI) (a) and $dbR=10log(precip)$ (b).
+(c-d) These fields translated to the ICON distribution.
+(e-g) The missing radiation channels infilled by cBottle-3d: outgoing shortwave (RSUT), surface downwelling shortwave (RSDS), and outgoing longwave (RLUT) (h) The downscaled dbR for this same example.
+This is computed by applying cBottle-SR after translation from ERA5 to ICON.
+These are shown for an atmospheric river example from 2006-07-11 12UTC.
+4 Conclusions
+We have introduced a generative AI model capable of synthesizing full global atmospheric states at 5 km resolution—over 12.5 million pixels per sample—without requiring autoregressive rollout.
+This milestone in resolution and sampling fidelity is achieved through a two-stage diffusion framework: a coarse-resolution generator conditioned on minimal, climate-controlling inputs (monthly SST, time of year/day), followed by a patch-based super-resolution module.
+Together, they enable low-latency, high-fidelity generation of multivariate climate states at the same ambitious resolution of the world’s next-generation operational climate simulation technology.
+In addition to resolution, cBottle demonstrates foundational versatility.
+Trained jointly on reanalysis (ERA5) and global cloud-resolving simulations (ICON), it permits multimodal translation between datasets, channel in-filling for corrupted or missing variables, bias correction and downscaling and zero-shot generalization of its global super-resolution module.
+This is made possible by a shared latent representation and masked training strategy (see Methods), allowing cBottle to flexibly bridge data modalities and tasks.
+Perhaps most importantly, cBottle proposes a new form of interactive climate modeling.
+By integrating classifier guidance into the diffusion process, users can conditionally generate physically plausible extreme events—such as tropical cyclones—at specified locations.
+This demonstrates that an AI climate model can sample rare events on-demand, opening the door to targeted scenario generation and new forms of climate risk analysis.
+Methodologically, we have introduced diffusion modeling on the HEALPix grid and proposed a noise schedule tuned to the spectral characteristics of climate data.
+In particular we show (see Methods) that the defaults most often used in natural image generation literature are insufficient for generating extremely large-scale modes like the seasonal and diurnal cycles.
+Fixing this required two main innovations: i) increasing the range of noise levels sampled during training and inference; and ii) adding appropriate time and position embedding.
+Specifically, we propose tuning the noise schedule based on spectral analysis of the data instead of expensive hyper-parameter tuning.
+We have also identified and mitigated over-fitting at high noise levels using an ensemble-of-experts approach; it is logical to speculate that similar issues may explain the lack of calibration in regional down-scaling models [29].
+By comparison our super-resolution models trained on large global storm-resolving models simulations are much less prone to over-fitting both owing to the quantity of the training data and the enhanced locality of the super-resolution task compared to global state generation.
+These perspectives should be helpful to anyone training diffusion models on scientific data.
+Additional limitations of our work are important to note.
+Non-stationary trends are not optimally represented in the current version of cBottle, with its long-term trend only half as big as observed.
+Similarly, the amplitude of NAM index of the generated samples is 20% too small, suggesting the model under-represents the variance in the largest scales.
+We attribute both of these weaknesses to the heavy-handed regularization required to prevent the coarse model from over-fitting at large noise levels.
+Nonetheless the modularity, scalability, and interactivity of cBottle represent a step towards a truly interactive Earth digital twin.
 5 Methods
-Figure 2: Climate in a Bottle (cBottle) Model Architecture.
-A cascaded approach with two complementary diffusion models that produces kilometer-scale climate data on demand.
-Inputs (left) include time-of-day, date-of-year, and SST, which feed into the Coarse Generation model (center) to produce 100km resolution outputs.
-These are then processed by the Super Resolution model (right) to transform coarse outputs to 5km resolution.
-This architecture enables climate emulation, channel in-filling, bias correction, downscaling, and what-if scenario exploration.
 5.1 Overview
-In this section we describe the components of our diffusion-based generative climate model shown in Figure 2.
-The architecture and diffusions are formulated on the sphere using the increasing popular HEALPix grid (Section 2.2).
-We train on two dense global datasets (Section 2.3) using a masked training strategy (Section 2.6).
-We introduce the preliminary details of diffusion modeling (Section 2.4) and our strategy for scaling to high-resolution data (Section 2.5).
-Finally, we present the architectures and hyper-parameters used for training our coarse image and video models, and the super-resolution model (Section 2.7).
-2.2 HEALPix and Padding
+CBottle is a cascaded diffusion model formulated on the HEALPix discretization of the sphere. It is trained on atmospheric simulations and reanalysis and consists of the following components:
+  • cBottle-3d generates hourly atmospheric conditions at approximately 1 degree resolution given monthly sea surface temperatures (SST), the day of year, data-set label, and the time of day as conditions.
+It is trained on ERA5 and ICON.
+  • cBottle-3d-tc is a variant of cBottle-3d capable of generating tropical cyclones in user-specified locations.
+It is trained with tropical cyclone track information data from the International Best Track Archive for Climate Stewardship (IBTrACS) [26].
+  • cBottle-SR is a super-resolution model trained to generate 5-km atmospheric states given 1 degree coarse resolution inputs. This is trained only on the ICON simulation data.
+The following sections provide a high-level overview of the data sources and our modeling approach.
+Detailed information about selected inputs and targets, architecture details, and hyper-parameter settings are shown in Section A.7.
+5.2 HEALPix
+Background
 Using AI for global data poses a geometric challenge of appropriately representing data on a spherical surface.
-To this end, HEALPix (Gorski et al., 2005) grids offer an attractive solution with an equal area grid structure that also has iso-latitude properties and a hierarchical indexing scheme for multi-resolution workflows.
+To this end, HEALPix [30] grids offer an attractive solution with an equal area grid structure that also has iso-latitude properties and a hierarchical indexing scheme for multi-resolution workflows.
 The HEALPix grid is constructed of 12 faces with each face containing a square $2^{l}$ x $2^{l}$ grid of pixels where $l$ represents the resolution level of the grid.
-This structure makes it very friendly for deep-learning algorithms, especially for convolutional architectures and has been used in (Krachmalnicoff and Tomasi, 2019), (Perraudin et al., 2019) and (Karlbauer et al., 2024).
-The equal area property is also particularly desirable for diffusion training.
+This structure is well-suited for deep-learning algorithms, especially for convolutional architectures and has been used in [31], [32] and [33]; the equal area property is also useful.
 Due to these desirable properties, we use the HEALPix grid as the spherical discretization method in this work.
@@
 We refer to two main global grids: HPX64 ($N_{side}=64$) corresponding to 100km resolution (49152 pixels) and HPX1024 ($N_{side}=1024$) at 6km resolution (12,582,912 pixels).
-2.2.1 Global modeling
-For coarse resolution applications (below HEALPix level 8), we use architectures similar to (Mardani et al., 2025) that were designed for 2D rectangular domains.
+Padding and convolution
+For coarse resolution applications (below HEALPix level 8), we use architectures similar to [13] that were designed for 2D rectangular domains.
 This is done by adapting the convolution operations in the architecture to perform convolutions on each HEALPix face independently.
-However, boundaries between HEALPix faces are not handled by such an approach and can cause artifacts.
-To address this, we pad each face using the protocol outlined in (Karlbauer et al., 2024) by the number of pixels needed for the convolution operation instead of relying on any implicit padding methods supported by convolution operations.
-Consider a field represented on a level $l$ the HEALPix grid, $x\in\mathbb{R}^{c_{in}\times 12\times 2^{l}\times 2^{l}}$.
+However, this approach does not account for the connected topology of the HEALPix grid and prevents the flow of information between adjacent faces.
+To address this, we apply padding by exchanging pixels from the neighboring HEALPix faces using the protocol outlined in [33] (see Section A.1).
+This technique is commonly referred to as a ghost-cell or halo update in numerical methods.
+5.3 Datasets
+The inputs and outputs we trained with are displayed in Tab.
+2.
+Briefly, the models are trained to generate 3D states of the atmosphere and some key 2D variables relevant to climate impacts (precipitation, surface air temperature, downwelling short-wave) and observations (top of atmosphere radiative fluxes).Our models are trained on two global datasets: the ECMWF Reanalysis version 5 (ERA5)[19] which is observationally constrained, and a global storm resolving simulation with the ICON climate model [20], which is a free-running simulation unpaired with reality.
+ERA5
+ERA5 is a commonly used dataset that represents the best guess of the global atmospheric state at a coarse resolution of 25 km.
+Briefly, it is created by running a data assimilation algorithm that incorporates observations into ECMWF’s numerical model.
+However, it is not available at a high enough resolution in space or time to resolve meso-scale motions on kilometer scale, and in particular deep convection.
+We train on hourly data from 1980–2017, leaving 2018 as validation.
+For consistency with the well-known AMIP protocol [22], we use its forcing instead of ERA5 for the sea-surface temperature input.
+The ERA5 shortwave and longwave radiation fields are corrupted by a known, but undocumented, bug which appears as clear stripe-like artifacts that point North-West (see Fig SI3).
+This is caused be a random seed in the radiation scheme being set to $lat+lon$ in every atmospheric column [34].
+Therefore, we do not use the fields from ERA5 for training our models.
+ICON Cycle 3
+The ICON data is a free-running simulation with the ICON atmospheric model coupled to a dynamic ocean and land [35].
+The atmospheric component solves the nonhydrostatic fluid mechanics equations on a global ico-sahedral mesh with a resolution of around 5 km.
+Unlike ERA5 the model explicitly resolves certain convective motions so a convection parameterization is not used.
+So even beyond the simple increase in resolution, we expect this to increase the fidelity of the precipitation and cloud fields in this dataset relative to ERA5 where these processes are parameterized.
+We train on 2020-01-20 03:00:00 through 2024-03-06 12:00:00 (inclusive) and validate on 2024-03-06 15:00:00 through 2025-07-22 00:00:00.
+IBTrACS
+The International Best Track Archive for Climate Stewardship (IBTrACS) [26] is used to train the TC classifier for guiding states towards TCs in a desired location.
+The dataset contains information collected from Regional Specialized Meteorological Centers (RSMCs) about position, intensity, and other parameters of TCs.
+Here, we use only the information about the storms’ position, based on the best track from multiple sources.
+While data is available from 1840 to present day, we use the part that overlaps with our ERA5 training set from 1980–2017.
+IBTrACS thus provides the labels for locations of TCs in the ERA5 data at a given time.
+See Section A.2 for detailed information about these data sources and our pre-processing.
+5.4 Denoising Diffusion
+Background
+As mentioned in the introduction, cBottle is a de-noising diffusion model [36, 37].
+This is a powerful class of generative models that learn to reverse a Gaussian stochastic process which starts from a sample of the data and ends in pure white noise.
+Specifically, we adopt the score-based diffusion framework [36, 13], which learns this de-noising process via a score-matching network.
+The training objective is given by
+$\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathbf{x}\sim p_{\text{data}}}\mathbb{E}_{\sigma\sim p_{\sigma}}\mathbb{E}_{\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})}\left\|\mathcal{D}_{\theta}(\mathbf{x}+\mathbf{\epsilon};\sigma)-\mathbf{x}\right\|^{2},$ (1)
+where the learned denoiser $\mathcal{D}_{\theta}(\cdot;\sigma)$ approximates the score function according to $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)\approx\frac{\mathcal{D}_{\theta}(\mathbf{x};\sigma)-\mathbf{x}}{\sigma^{2}}$.
+Here, $p_{\sigma}$ denotes the noise level distribution used during training.
+To sample from the model, we employ the Elucidated Diffusion Model (EDM) sampler proposed by [9], which solves the corresponding reverse-time ordinary differential equation (ODE):
+$\frac{d\mathbf{x}}{d\sigma}=\frac{\mathbf{x}-\mathcal{D}_{\theta}(\mathbf{x},\sigma)}{\sigma}.$ (2)
+Specifically, we use Algorithm 2 from Karras et al. [9], with the $S_{churn}$ parameter set to 0. This ODE is sampled backwards from $\sigma=\sigma_{max}$ to $\sigma=\sigma_{min}$.
+Diffusions on the sphere
+To train diffusion models on the sphere, the noise $\epsilon_{i}\sim~N(0,1)$ is drawn independently for every pixel $1\leq i\leq 12\cdot 4^{l}$.
+Because the HEALPix pixels are all equal area, we conjecture that this is identical to the pixel-wise average of a finite-variance continuous Gaussian process on the sphere.
+Noise schedule
+The choice of noise levels seen during training ($p_{\sigma}$) and inference ($\sigma_{min}$ and $\sigma_{max}$) is crucial for the quality of a diffusion model.
+Compared to the typical settings used for natural image generation, we find that much larger noise levels must be sampled during training for a model to generate realistic seasonal and diurnal cycles.
+When the noise is too small, the model never learns to create these signal on its own—a phenomenom known as the signal-leak bias [38].
+We therefore select a broadly distributed log-uniform noise distribution with $\sigma_{min}=0.02$ and $\sigma_{max}=200.0$.
+These values are chosen based on a spectral analysis to cover the dynamic range between the lowest and highest variances modes of the data—without a need for expensive hyperparameter tuning.
+Specifically, for data $\mathbf{}\in\mathbb{R}^{n}$, $\sigma_{max}$ should be chosen proportional to $\sqrt{n}s_{1}$ where $s_{1}^{2}$ is the fraction of variance explained by the first EOF of the data.
+See Section A.3.1 for a detailed derivation.
+Ensemble of experts regularization
+Similar spectral considerations reveal that diffusion models are prone to over-fitting at high noise levels when trained on a limited number of samples.
+This occurs because there are few independent observations of such large variance signal such as El Niño or the NAM in the observed record, and a large neural networks is fully capable of memorizing these when given a suitably rich input like SST.
+We avoid this over-fitting by adapting the ensemble of experts approach of Yogesh et al.
+[39].
+While they trained separate models for different noise levels in order to increase overall model capacity, the method can just as easily be adapted to regularize the large noise levels more than the small.
+In our approach, separate denoisers (i.e experts) for small ($\sigma<10$), medium ($10<\sigma\leq)100$, and large ($100<\sigma\leq\sigma_{max}$) noise levels.
+For simplicity and to avoid extra training, we regularize via early stopping, and simply use less-trained checkpoints for the larger noise levels.
+Section A.3.2 contains more details and sensitivity tests.
+5.5 Multi-diffusion
+Emulating ICON at HPX1024 resolution poses substantial computational challenges due to the massive size of this km-scale global dataset.
+Directly applying a traditional diffusion model to the full-resolution domain is infeasible as it would require over 2,000 GB of GPU memory, assuming a linear relationship between spatial resolution and memory footprint.
+This far exceeds the capacity of current hardware, even on high-end multi-GPU systems.
+To address the computational bottleneck and take advantage of the inherently local nature of super-resolution tasks, multi-diffusion [40] is adapted as a scalable and memory-efficient sampling strategy.
+It enables high-resolution generation without sacrificing global spatial coherence or requiring model architecture changes.
+The central idea behind multi-diffusion is to replace the global denoiser used in standard sampling algorithms with an aggregated local denoiser.
+At each sampling step, the aggregated local denoiser operates on the global feature map in a sliding-window fashion, denoising it patch by patch.
+These locally denoised patches are then reassembled to reconstruct the global feature map.
+This approach is simple to adapt to the HEALPix grid by exchanging neighboring information between the 12 faces via the padding procedure described above.
+See Section A.4 for details.
+5.6 Multi-modal training and inference
+Masked Training
+Since we intend our framework to extend to a variety of data sources, with different coverage in space, time, and the included fields, we use masked training.
+Our approach is to simply train a diffusion model to generate the union of all channels present in all the datasets and mask any missing values for a given dataset—such as ERA5’s corrupted radiation fluxes—with 0 after normalization.
+More specifically, any masked values in the noisy image $x+n$ and target $x$ are filled with 0 before calling the network and computing the loss.
+We then normalize the loss by the fraction of masked out data points, so that the training equally weights modalities with different numbers of observables.
+Pseudo-code for this approach is shown in SI A.5.
+Channel infilling
+Given the masked training described above, we can use the trained model to infill channels such as the ERA5 radiation fields given the other ERA5 channels.
+Following Appendix I.2 of Song et al.
+[41], let $\Omega:\mathbb{R}^{C\times\mathbf{d}}\to\mathbb{R}^{C_{\mathcal{K}}\times\mathbf{d}}$ and $\bar{\Omega}:\mathbb{R}^{C\times\mathbf{d}}\to\mathbb{R}^{C_{\mathcal{U}}\times\mathbf{d}}$ denote the slicing operators that extract the known and unknown channels, respectively, where $\mathbf{d}$ denotes an arbitrary set of additional dimensions (e.g., spatial or temporal).
+The unknown subset at sampling time $t$ is then denoted by $\mathbf{z}_{t}=\bar{\Omega}(\mathbf{x}_{t})$.
+The conditional score function required to impute the missing channels can be approximated as
+$\nabla_{\mathbf{z}}\log p\big{(}\mathbf{z}_{t}\mid\Omega(\mathbf{x}_{0}),\sigma_{t}\big{)}\approx\nabla_{\mathbf{z}}\log p\big{(}[\mathbf{z}_{t};\,\Omega(\mathbf{x}_{t})]\mid\sigma_{t}\big{)},$ (3)
+[…]
+Translating between modalities
+To map an ICON sample to the distribution of the ERA5 data (closer to observations), we take advantage of the latent representation of the diffusion model.
+Our diffusion model being trained with ICON and ERA5 data can be regarded as two different models (with shared weights, the only difference being whether we train or inference with an ICON or with an ERA5 flag passed to the model).
+With these two models, we can follow a Dual Diffusion Implicit Bridge approach [42], which consists in evaluating the diffusion ODE forward (mapping an ICON sample to noise), switch the flag to ERA5 and evaluate the ODE backwards (mapping the noise to ERA5).
+It may seem surprising that this yields a realistic sample, but as Liu et al.
+[43] explains the uniqueness property of differential equations ensures that trajectories between the data and noise distributions cannot cross one another.
+Therefore using the deterministic mapping from data to noise for one domain and from noise to data for the other domain, allows us to discover mappings between the two domain with low transport cost.
+Unlike e.g.
+rectified flow [43] our model is not trained explicitly to map from one dataset to the other.
+That approach would require training a separate model for each pair of datasets so does not scale well to many different data sources.
+By training one model on all potential datasets, we can use the denoiser to define a shared latent space, a desirable property of a foundation model.
+5.7 Classifier Guidance
+We demonstrate the capacity of cBottle to be used in a more interactive way than both auto-regressive climate emulators such as ACE2 [3] and numerical GCMs with the example of Tropical Cyclones (TCs).
+To do this, we implement classifier guidance [27] with an image segmentation model trained on historical cyclone data from IBTracs [26].
+We initially experimented with explicitly conditioning on the training by randomly drawing a coarse-resolution HPX8 pixel and labeling it based on whether it contained TC or not.
+However, this approach failed to generate TCs in the desired location—likely because of the sparseness of this input.
+By comparison a classifier trained with cross-entropy loss to segment the image into TC and not-TC classes learns from every TC present in any given sample, enabling a more efficient use of the data.
+This classifier is trained with the noise image $x(\sigma)$ and noise level $\sigma$ as an input so that it can be used in the sampler ODE (2) where only $x(\sigma)$ is available, rather than clean state $x(0)$.
+During the inference phase, this classifier is used to guide the sampler based on a user-specified mask indicating the desired TC locations.
+See Section A.6 for more details.
+6 Data Availability
+The ICON data is archived at the Max Planck Institute for Meteorology [44].
+ERA5[19] is available for download via the Copernicus Climate Data Store and is mirrored at many sites around the globe.
+The monthly sea surface temperature inputs for the ERA5 model were downloaded from the inputs4MIPS section of Earth System Grid Federation.
+Our trained model parameters are available on NVIDIA NGC at this URL:
+https://catalog.ngc.nvidia.com/orgs/nvidia/teams/earth-2/models/cbottle.
+The code is available at https://github.com/NVlabs/cBottle.
+Acknowledgements
+This work would not have occurred had Bjorn Stevens and Daniel Klocke of the Max-Planck Institute for Meteorology not provided the ICON Cycle 3 simulation in a format perfect for training machine learning models, and we also owe David Pruitt of NVIDIA a debt for helping transmit this sizeable dataset onto our systems.
+We also thank Oliver Watt-Meyer and Christopher Bretherton of the Allen Institute of Artificial Intelligence for their critical observation that the released version cBottle was over-fitting to the NAM, and Karsten Kries (NVIDIA) for helpful discussions about the spectral perspective on diffusion models.
+Appendix A Supplemental Methods
+A.1 HEALPix convolutions
+Figure 6: Schematic showing the multi-diffusion approach for spherical geometry with HEALPix grids.
+A sequence of the overlapping patch sampling protocol is shown going clockwise from the top left.
+Each colored region indicates a single patch and overlap between the patches minimizes boundary artifacts by combining patch representations into a global representation following 11.
+Figure 7: Schematic showing the global UNet architecture for spherical geometry with HEALPix grids. Each convolution is applied on HEALPix faces padded as in [33].
+Consider a field represented on a level $l$ of the HEALPix grid, $x\in\mathbb{R}^{c_{in}\times 12\times 2^{l}\times 2^{l}}$.
 One may apply a discrete convolution kernel $\mathcal{K}_{conv}\in\mathbb{R}^{c_{out}\times c_{in}\times 2k+1\times 2k+1}$ as follows:
@@
 $y$ $=\left(y_{1},y_{2},\cdots y_{12}\right),$ $\quad y\in\mathbb{R}^{c_{out}\times 12\times 2^{l}\times 2^{l}}$ (7)
-where $x_{f}$ and $y_{f}$ are the pixels of $x$ and $y$ respectively on the HEALPix face $f$, $\mathcal{P}(.;k)$ is the padding operation of (Karlbauer et al., 2024) and $\ast$ indicates the convolution operation.
-Figure SI3 shows a schematic of how the UNet architecture of (Song et al., 2021) is adapted to work on global HEALPix fields.
-2.2.2 Local modeling
-Modeling of high resolution fields, like kilometer-scale climate model outputs, is prohibitively expensive using global approaches.
-To make this tractable, we use the multi-diffusion approach outlined in Section 2.5 and adapt that to spherical manifolds using the HEALPix grid and HEALPix padding.
-We define a new spatial sampling operator ${\mathcal{G}_{i}}:\mathcal{X}\rightarrow\mathcal{X}_{\text{local},i}$ as a composition of the padding operator $\mathcal{P}(.;N/2)$ and the 2D sampling operator $\mathcal{F}_{i}$ defined in Section 2.5.
+where $x_{f}$ and $y_{f}$ are the pixels of $x$ and $y$ respectively on the HEALPix face $f$, $\mathcal{P}(.;k)$ is the padding operation of [33] and $\ast$ indicates the convolution operation.
+Figure 7 shows a schematic of how the UNet architecture of [36] is adapted to work on global HEALPix fields.
+A.2 Data sources and pre-processing
+ERA5
+For pre-processing, we retrieved hourly ERA5 data for 1980–2018 (inclusive) via the data lake at the National Energy Research Computing Center (NERSC) at $0.25^{\circ}$ resolution on the lat-lon grid.
+The 3d atmospheric states are available on pressure levels.
+We then regridded this using bilinear interpolation to the HEALPix grid with $N_{side}=256$.
+For training the coarse-resolution model, this coarsened by averaging pooling to an $N_{side}=64$.
+We train on years 1980–2017 (inclusive) and validate on 2018, though many of our climate diagnostics below are also computed on the validation set (as is common in generative modeling).
+AMIP SST Forcing
+Our models take a monthly time-scale SST as an input. For the ERA5 data which is paired with reality, we linearly interpolate the adjusted monthly SST time series following the AMIP protocol [22].
+ICON Cycle 3
+We obtained O(PB) of ICON data (ngc3028) of the nextGEMS cycle 3 simulations [35] in zarr format from DKRZ’s JupyterHub server.
+This dataset was stored on the HEALPix grid with $N_{side}=1024$, corresponding roughly to a resolution of $\sqrt{A/(12N_{side}^{2})}=6.3\text{km}$.
+Prior to our handling of the data it was interpolated using nearest neighbors from the native icosahedral grid.
+Five years of this data are available at 3-hour (3D) and 30-minute (2D) resolution in time.
+Conveniently, this dataset featured pre-coarsened data (again using averaging pooling) at our coarse-resolution of $N_{side}=64$.
+We interpolated this coarse data to fixed pressure levels using linear interpolation in the vertical direction.
+To fill in values at pressure level locations that are below the surface, we use extrapolation based on hydrostatic balance constraints and an assumption of a constant temperature lapse rate of 6.56.56.5 K/km for temperature and geo-potential..
+For all other variables, we use constant extrapolation down from the surface.
+This procedure approximately reproduces ERA5’s undocumented procedure for filling-in below-surface levels.
+We compute the SST input for ICON with a rolling window average in time with a window size of 30 days based on the HPX64 spatial and daily temporal resolution data in the ICON dataset.
+AMIP SST Forcing
+We downloaded the file: input4MIPs.CMIP6Plus.CMIP.PCMDI.PCMDI-AMIP-1-1-9.ocean.mon.tosbcs.gn from the Earth System Grid Federation, and use bilinear interpolation in space to map it to HPX64 resolution.
+A.3 Diffusion Training
+A.3.1 Tuning noise schedules based on the data distribution
+The choice of noise levels seen during training ($p_{\sigma}$) and inference ($\sigma_{min}$ and $\sigma_{max}$) is crucial for the quality of a diffusion models.
+Unfortunately, our models failed to produce seasonal or diurnal cycles when using the standard settings from [9], suggesting that these hyper-parameters settings tuned for image generation are not suitable for generating atmospheric states (see Figure SI22).
+The basic issue is known as the signal leak bias when a model is not trained with enough noise to entirely swamp the largest modes present in a dataset [38].
+Without enough noise, it is easy to recover high-amplitude and large-scale signals like the seasonal or diurnal cycles—for example by spatial averaging—so the model never learns to create these signals on its own.
+In this section, we propose a criterion for adjusting this settings based only on a spectral analysis of the data, without a need for expensive hyper-parameter tuning.
+Specifically, the upper and lower bound of the noise ranges spanned during training $p_{\sigma}$ and the initial and final times for (2), $\sigma_{min,max}$, should be determined by
+$\sigma^{2}_{min}$ $\lesssim\inf_{||v||=1}E|v^{\prime}x|^{2}$ (8)
+$\sigma^{2}_{max}$ $\gtrsim\sup_{||v||=1}E|v^{\prime}x|^{2}.$ (9)
+The right hand sides are the minimal and maximal variances along any direction of the data spaces—and are equivalent to the maximum and minimum singular values of the covariance matrix $E|xx^{\prime}|$.
+The signal leak bias occurs when $\sigma_{max}^{2}$ is too small, such that the highest variances mode of the data is not covered up.
+When $\sigma_{min}$ is too large, the diffusion model will not denoise the lower-amplitude modes of the data—typically the highest spatial frequencies.
+Choosing $\sigma_{min}$ too large manifests as an incorrectly flat spectra for the highest wave numbers (see e.g., the Z500 power spectra presented by Price et al.
+[12]).
+[…]
+$|v^{\prime}x|\gg\sigma_{min}|v^{\prime}\epsilon|$
+Likewise for high noise levels we demand that the signal is lost in the noise which means that
+$|v^{\prime}x|\ll\sigma_{max}|v^{\prime}\epsilon|.$
+(9) are recovered from this equations by taking the expectation of these inequalities and noting that $E|v^{\prime}\epsilon|^{2}=|v|^{2}E|\epsilon|^{2}=1$.
+Note that (9) is equal to the variance explained by the maximum principal component of the data.
+Recall that the overall variance can be decomposed $n=TrCov(x)=\sum_{i}s^{2}_{i}$ where $n$ is the dimension of $x$ (pixels and channels).
+To understand the behavior for changing $n$ it is useful to discuss the fraction of variance explained by a singular value $\tilde{s}^{2}_{i}=s_{i}^{2}/n$, and expect this fraction to converge as we refine the resolution.
+This implies then, that $\sigma_{max}^{2}\gtrsim n\tilde{s}_{1}^{2}=O(n)$.
+In atmospheric science, the seasonal cycle and diurnal cycles are the dominant modes of variability since they are: 1) planetary in spatial scale; and 2) large in amplitude.
+In mid-latitude land locations, the daily high and low temperature are separated by $\sim 5-10^{\circ}$, while the seasonal cycle is $\sim 10-20^{\circ}$.
+Other variables don’t have as strong a seasonal cycle over as large an area so we take the conservative estimate that the seasonal cycle explains 5% of the global variance across our whole variable set.
+The HPX64 model has about 30 channels in total and 49152 pixels.
+So this would imply a $\sigma_{max}^{2}\gtrsim 300$.
+This is far larger than the range of the hyper parameters used for this resolution in typical image generation [9].
+Similar analysis can explain the observation that generating higher-resolution images requires larger noise levels [45].
+While one can tune the log-normal distribution to fully cover the dynamic range of the data $[\sigma_{min},\sigma_{max}]$, we have found that sample quality is sensitive to shifts in the mean of the log-normal distribution.
+As a simpler approach, we propose the following noise schedule
+$p_{\sigma}\propto\sigma^{-1},\sigma_{min}\leq\sigma\leq\sigma_{max}.$ (10)
+Because this is uniform in $\log\sigma$ we call it a “log-uniform” noise distribution.
+In earlier tests on the 2D data alone without SST conditioning, we achieved similar distortions of the power law distribution as reported by Price et al.
+[12], i.e.
+incorrectly flat spectra for the highest wave numbers.
+In practice, we set $\sigma_{min}=0.02$ and $\sigma_{max}=200$, settings which span the dynamic range of several of our variables (see Figure SI18; in hindsight we should have decreased $\sigma_{min}$ even further for the smoothest fields like Z500).
+For inference, we use the same sampler as Karras et al.
+[9] but with these same settings.
+A.3.2 Avoiding over-fitting at large noise levels
+Similar spectral considerations reveal that diffusion models are prone to over-fitting at high noise levels when trained on a limited number of samples.
+This occurs because there are very few independent observations of such large variance signals in the training data due to these conspiring factors:
+  • Training at large noise-levels reduces the number of degrees of freedom per sample.
+This can be illustrated by assuming that $\sigma$ is sufficiently large such that only 10 singular values of the data are larger than $\sqrt{n}\sigma$ and that corresponding principal components are independent random variables.
+The same reasoning would apply if we chose the 10 most energetic Fourier modes.
+In either case, only 10 independent degrees of freedom per sample can be recovered at this noise level.
+  • Time series data is auto-correlated in time.
+Moreover, in many physical systems like the atmosphere, the auto-correlation of the largest modes of the data tend to be the slowest.
+For example, the highest-variance modes of the global ocean and extra-tropical atmosphere, respectively, are ENSO, which has an inter-annual timescale, and the annular modes with a multi-week timescale.
+So in 40-year duration of ERA5 there are only a few dozen ESNO events and few hundred oscillations of the annular modes.
+  • Conditional inputs that are easy to memorize. For example, the network can easily learn to map SST input to the timestamp of the sample and then use this timestamp to recover the desired outputs.
+Since the large noise levels are exponentially less likely to be sampled during training, they comprise a small portion of the overall loss so this over-fitting can be easy to miss.
+Therefore, it is critical to monitor the relationship between the training and validation loss and the noise level $\sigma$.
+Indeed, we see that this is a problem for the coarse-resolution cBottle-3d model especially for $\sigma\geq 100$ (Fig.
+SI25).
+The small noise levels, by comparison, show no signs of over-fitting to the validation set.
+We control this over-fitting with an ensemble of experts [39] approach where separate models are used for small (0,10](0,10]( 0 , 10 ], medium (10,100](10,100]( 10 , 100 ], and large (100,)(100$\infty$)( 100 , ∞ ) noise levels $\sigma$ during inference.
+This allows us to regularize the large noise level models more than the small.
+For simplicity and to avoid extra training, we regularize via early stopping.
+Specifically, we select the small, medium, and large models to be the checkpoints after 9.856M, 2.0248M, and 512K samples of training, respectively.
+Experiments with activation and condition dropout yielded better validation losses, but the gains were marginal and required extensive tuning.
+Other studies have also noted the tendency of diffusion models to over-fit to their training data.
+Similar to the present setting, Wang et al.
+[46] note that the generalization gap between training and out of distribution data for diffusion language models is larger for increasing noise levels.
+Favero et al.
+[47] also propose a criteria for early-stopping based regularization that depends on the dataset size, but do not explicitly connect this with the need to regularize high levels more than small or spectral analysis.
+A.4 Multi-diffusion: details
+Figure 8: (a) Training workflow of the multi-diffusion framework.
+Global maps are decomposed into patches for parallel training of local denoisers conditioned on spatially aligned context.
+(b) Illustration of the aggregated local denoiser replacing the global denoiser in the sampling algorithm.
+This procedure is illustrated in Figure 8(b).
+More concretely, given a pretrained local denoiser $\Phi_{\theta}$ and a set of spatial sampling operators ${\mathcal{F}_{i}}:\mathcal{X}\rightarrow\mathcal{X}_{\text{local},i}$, the global denoiser can be represented as an aggregation of local denoisers:
+$\mathcal{D}_{\theta}(\textbf{x}_{t})=\frac{\sum_{i}\mathcal{F}_{i}^{-1}\left(\Phi_{\theta}\left(\mathcal{F}_{i}(\textbf{x}_{t})\,\middle|\,\mathcal{F}_{i}(y),\sigma\right)\right)}{\sum_{i}\mathcal{F}_{i}^{-1}(J_{N})},$ (11)
+where $\mathcal{F}_{i}^{-1}:\mathcal{X}_{\text{local},i}\rightarrow\mathcal{X}$ denotes the inverse operation that maps local patches back to the global space, $y\in\mathcal{Y}$ denote the spatial conditions, $\mathcal{X}_{local}=\mathbb{R}^{c_{out}\times N\times N}$ denotes the local target space, $\mathcal{Y}=\mathbb{R}^{c_{in}\times N\times N}$ denotes the local condition space, $N$ is the patch size, and $J\in\mathcal{X}_{\text{local}}$ is an all-one matrix of size $N\times N$ used to account for the number of overlapping patches at each location in the global map.
+Multi-diffusion can easily be adapted to the HEALPix grid.
+Define a new spatial sampling operator ${\mathcal{G}_{i}}:\mathcal{X}\rightarrow\mathcal{X}_{\text{local},i}$ as a composition of the padding operator $\mathcal{P}(.;N/2)$ and the 2D sampling operator $\mathcal{F}_{i}$ defined in Section 5.5.
 Here $N$ represents the patch size used in the multi-diffusion approach.
@@
 Figure 6 shows a schematic of this approach.
-Figure 3: Schematic showing the multi-diffusion approach for spherical geometry with HEALPix grids.
-A sequence of the overlapping patch sampling protocol is shown going clockwise from the top left.
-Each colored region indicates a single patch and overlap between the patches minimizes boundary artifacts by combining patch representations into a global representation following 14.
-2.3 Datasets
+A fundamental design principle of multi-diffusion is the global sharing of the latent representation $X$ across all local patches.
+The spatial sampling operators ${\mathcal{F}_{i}}$ define the patch layout, including the extent of overlap between adjacent regions and the total number of patches required to span the full domain.
+Introducing overlap is essential for reducing boundary artifacts and promoting consistency across patches.
+However, increasing the overlap also raises the computational cost, as it leads to a larger number of denoiser evaluations during sampling.
+As demonstrated in the ablation study in Subsection B.7, strong performance can be maintained even with minimal overlap, enabling a favorable trade-off between quality and efficiency.
+A.5 Pseudo-code for masking in the loss function
+D_yn = net(where(mask, y + n, 0), sigma)
+loss = weight * ((D_yn - where(mask, y, 0)) ** 2)
+loss = where(mask, loss, 0) / mask.mean()
+A.6 Classifier guidance details
+The classifier outputs a 12×8×812$\times$ 8$\times$ 812 × 8 × 8 (HPX8) map of the probability that a given pixel contains a TC.
+For the architecture, we exploit the HPX8 resolution feature maps learned by the coarse-resolution denoiser (see Sec.
+A.7), and obtain the probability map by taking the lowest-resolution state of the UNet and feeding it through two convolutional layers (with 50% dropout in between).
+During training, we minimize binary cross-entropy between the TC labels from IBTracs and the outputs of the classifier.
+In the inference phase, guidance is implemented by adding a term to the ODE in eq. 2:
+$\frac{d\mathbf{x}}{d\sigma}=\frac{\mathbf{x}-\mathcal{D}_{\theta}(\mathbf{x},\sigma)}{\sigma}+\gamma\nabla_{\mathbf{x}}\mathcal{L}_{BCE}(\mathcal{C}(\mathbf{x}),y).$ (14)
+Here, $\gamma$ represents a guidance scale hyperparameter, and $\mathcal{L}_{BCE}(\mathcal{C}(\mathbf{x}),\mathbf{y})$ the binary cross-entropy loss of the classifier output $\mathcal{C}(\mathbf{x})$ and the guidance input $\mathbf{y}$.
+In practice, we found that a constant guidance scale did not produce the desired effect, and so used an adaptive guidance scale $\gamma=\hat{\gamma}\rho$, where $\hat{\gamma}$ is constant and
+$\rho=\frac{||\frac{\mathbf{x}-\mathcal{D}_{\theta}(\mathbf{x},\sigma)}{\sigma}||^{2}}{||\nabla_{\mathbf{x}}\mathcal{L}_{BCE}(\mathcal{C}(\mathbf{x}),y)||^{2}}$ (15)
+with the Frobenius norm $||\cdot||^{2}$.
+This ensures that at each step, the fraction of the norm of the update in the score direction and of that in the classifier direction is constant.
+Without this normalization, we found the classifier gradient to be small compared to the score in particular at early denoising times.
+A.7 Architectures and hyperparameter settings
+Figure 9: Position and calendar embedding.
+cBottle-3d cBottle-video cBottle-SR
+Number of parameters 149M 282M 330M
+Target channels 3D 3D 2D
+Architecture Song UNet Song UNet Song UNet
+Learned Position Encoding
+Calendar embedding
+Trained on ICON+ERA5 ICON+ERA5 ICON
+Grid HPX64 (49152) HPX64 (49152) 128x128 crops of HPX1024
+Training domain global global crops
+Total examples trained 9,856,000 2,160,000 95,520 samples / 191M patches
+Batch size 64 (3M images) 256 (3-10M) 32 960
+Optimizer Adam Adam SGD
+Learning rate 0.0001 0.0001 0.0001
+Noise distribution ($p_{\sigma}$) log-uniform 0.02–200 log-uniform 0.02–1000 log-normal
+$P_{\text{mean}}=-1.2$ $P_{\text{std}}=1.2$
+Computational Profile
+Hardware trained on 16xH100 32xH100 64xH100
+Examples/GPU-second (train) 6.7 0.32 35.4 patches
+Table 1: Comparison of cBottle model variants.. A training example is defined a single HPX64 image for the coarse models but is a single 128x128 crop for cBottle-superres.
+We mostly reuse the standard UNet architecture used in diffusion modeling (the so-called Song-UNet in Karras et al.
+[9], Mardani et al.
+[13] and others).
+However, it was necessary to add position and temporal embedding (Figure 9 to allow the diffusion model to learn non-stationary signals in space and time, and temporal attention layers to support video diffusion.
+Table 1 summarizes our three model variants.
+Calendar Embedding
+Producing an accurate seasonal and diurnal cycle required a calendar and hour embedding strategy. We take care to use a embedding that respects the periodicity of these inputs. This is given by
+$f_{4k}$ $=\cos\left(2\pi k\frac{\text{doy}}{365.25}\right)$ (16)
+$f_{4k+1}$ $=\sin\left(2\pi k\frac{\text{doy}}{365.25}\right)$ (17)
+$f_{4k+2}$ $=\cos\left(kt\right)$ (18)
+$f_{4k+3}$ $=\sin\left(kt\right).$ (19)
+Here, the day of year (doy) is a scalar representing the time since the beginning of the year, and the time of day ($t$) is the local solar time in units radians $t=2\pi(\frac{\text{second of day}}{86400}+\frac{\text{lon}}{360})$ [^2] Due to an implementation error we actually used $second-lon$..
+In the architectures below we use $k=1,\ldots,8$, and concatenate these inputs along with other image-like conditioning (e.g., monthly-averaged SST) and noisy target image.
+Learned position embedding
+To allow the diffusion model to learn non-stationary spatial structures like climatology, we add a learned position embedding.
+After the initial convolution embedding of the image-like inputs (calendar, noisy image and monthly-averaged SST), a learned positional embedding for each pixel $49152\times c$ is added.
+The position embedding is initialized with independent standard normal random numbers.
+A.7.1 Model Settings
+cBottle-3D
+The coarse image model generates HPX64 fields.
+The coarse resolution image generator uses the standard settings for the Song-UNet presented by Karras et al.
+[9], but we found the 3D training was made more stable by increasing base model embedding dimension from 128 to 192.
+The source dataset is identified with a label (ERA5 or ICON) which is randomly dropped for 1/4 samples to encourage the model to learn features across both datasets.
+This label is provided to the UNet in the same manner as the imagenet class label in Karras et al.
+[9].
+Self-attention is only used on the coarsest resolution stage (8x8).
+All convolutional layers are implemented on the HPX grid using the padding approach described in Section 5.2.
+32 channels are devoted to calendar embedding.
+The model was trained from scratch on a 50-50 mix of ICON and ERA5 data on 32 H100 GPUs.
+Even (odd) ranks were trained with ERA5 (ICON).
+The model was trained to 3 million samples (with repeats) with a batch size of 64 (50000$\sim$ 50000∼ 50000 steps) when the log probability loss started to plateau, and then the batch size was increased to 256 until 10 million samples.
+The Adam optimizer [48] was used with a learning rate of $10^{-4}$.
+During training the model uses the log-uniform noise distribution ($p_{\sigma}$) with $\sigma_{min}=0.02$ and $\sigma_{min}=200$.
+For inference, we use the power law sampler proposed by Karras et al.
+[9] with these same settings.
+cBottle-SR
+The super-resolution model, which performs upscaling from HPX64 to HPX1024 resolution (a factor of 16), is trained within a HEALPix multi-diffusion framework described in 5.5.
+In this setting, the global input data—including latent variables, conditioning fields, and high-resolution targets—are partitioned into spatial patches.
+Each patch is then independently processed by a local denoising model, which is trained to reconstruct high-resolution content based on temporally and spatially aligned inputs.
+This approach enables efficient learning of fine-scale features while preserving global coherence through patch-wise training across the full spherical domain.
+As illustrated in Figure 8(a), this training strategy adapts the standard denoising score matching loss (see Eq. 1) to operate at the patch level. Specifically, the modified objective is defined as
+$\min_{\theta}\mathbb{E}_{\mathbf{x}\sim p_{\text{data}}}\mathbb{E}_{\sigma\sim p_{\sigma}}\mathbb{E}_{\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})}\mathbb{E}_{i\sim\mathcal{U}}\Big{|}\Big{|}\Phi_{\theta}\Big{(}\mathcal{F}_{i}(\mathbf{x})+\mathbf{\epsilon};\mathcal{F}_{i}(y),\sigma\Big{)}-\mathcal{F}_{i}(\mathbf{x})\Big{|}\Big{|}^{2},$ (20)
+Like cBottle-3d we use the Song-UNet with 128 hidden channels. Since it is a local model, we do not need to use the HEALPix padding shown in 7.
+A patch size of 128$\times$128 is used.
+This is chosen to be much larger than 16x super-resolution factor but is still affordable.
+To cope with non-stationary spatial features like topography, 20 learnable positional embeddings with a spatial size of HPX1024 grid are introduced and initialized using sinusoidal functions.
+The inputs of the network consist of (1) the HPX64 data re-gridded bilinearly to the HPX1024 128$\times$128 patch in question, (2) the global position embeddings for this patch, and (4) the global HPX64 data interpolated to the 128×128 lat-lon grid.
+These are concatenated with the latent before further processing by the encoder.
+The model was trained to 95520 ICON samples (corresponding to 191 million patches, with repeats) with a batch size of 960.
+The SGD optimizer was used with a learning rate of $10^{-4}$.
+During training the model uses the log-normal noise distribution of Karras et al.
+[9] ($p_{\sigma}$) with $P_{mean}=-1.2$ and $P_{std}=1.2$.
+For inference, the super-resolution model uses a maximum noise level ($\sigma_{max}$) of 800, 18 sampling steps, and $S_{churn}=0$.
+A.8 Video Generation extension
+Let $x\in\mathbb{R}^{T\times C\times X}$ denote an atmospheric trajectory of $T$ evenly spaced frames with $X$ pixels each.
+While the coarse image cBottle produces realistic atmospheric snapshots, $T$ successive samples in time produce a sequence that lacks temporal coherence due to the independence of each sample.
+A.8.1 Architecture Modifications
+Temporal Attention
+We extend cBottle to video generation by augmenting the 2D UNet architecture with temporal self-attention layers, following factorized space-time approaches [49, 50].
+The architecture preserves the UNet backbone of cBottle, but introduces temporal layers at the end of spatial blocks.
+Within each UNet block, we first apply spatial modules with the time dimension folded into the batch dimension.
+For blocks with temporal attention, we fold the spatial dimension into the batch dimension, and then apply bidirectional self-attention across frames per pixel, using the channels as the embedding dimension.
+A learned relative-time embedding $R_{\Delta t}\in\mathbb{R}^{2T-1}$ capturing frame-to-frame dependencies based on temporal distance $\Delta t$ is added to the pre-softmax logits.
+The temporal attention output is then residually added to the main branch.
+Masked-Conditioning
+To enable flexible inference capabilities, we adopt a masked-conditioning training scheme inspired by MCVD [51].
+During training, we randomly sample mask patterns $m\in\{0,1\}^{T}$ (where 1 indicates observed frames) from a categorical distribution $p_{mask}$ that blends several tasks: Bernoulli random frame dropout, masking all but the first or last few frames contiguously, masking middle frames while keeping endpoints visible, and full sequence masking.
+On top of the calendar embedding and monthly-mean SST conditioning that the image-model uses, we add this masked sequence (ground truth at unmasked positions and zeros at masked positions) and the binary mask indicator as conditioning concatenated channel-wise with the input.
+The video diffusion training objective is identical to the image diffusion objective, but with this additional per-sample masked-conditioning and the loss averaged across the $T$ frames.
+This unified training approach enables the video model to perform multiple tasks at inference time: generate entire sequences unconditionally, predict future frames given initial states, or infill arbitrary temporal gaps.
+A.8.2 Model settings and training
+The video model generates 3-day sequences with 6-hour resolution ($T=12$ frames) at the same HPX64 resolution of the image-based cBottle.
+Because $T$ correlated frames carry more redundant signal, we keep the same log-uniform noise schedule (with $\sigma_{min}=0.02$) and increase $\sigma_{max}$ from 200 to 1000 in both training and inference to preempt signal leak bias (see Subsection A.3.1).
+We insert temporal‐attention layers into every encoder block and into two positions within each decoder resolution stage: the up‑sampling block and the final block.
+We weight each of the four masking tasks defined above as follows: random frame dropout (0.3), block-wise masking (0.25), endpoint interpolation (0.1), and full sequence masking (0.35).
+Similar to the coarse image model, a base model embedding dimension of 128 sufficed to model the 2D dataset’s 12 fields.
+Given the video model’s expanded 124-channel input on the 3D dataset, we increased the UNet width to 256 channels, finding it to converge markedly faster and achieve higher log-likelihood than at 192 channels.
+The model was trained from scratch like the image model on 32 H100 GPUs, using the same Adam optimizer [48] and learning rate of $10^{-4}$.
+It was trained to 2.16M video samples with a batch size of 32.
+A.8.3 Results
+To evaluate whether the video diffusion model captures realistic atmospheric dynamics through time—beyond just generating independent plausible snapshots—we analyze its ability to reproduce the patterns characteristic of mid-latitude storm track evolution in Figure 10.
+Following the approach of [52], we compute one-point lag correlation maps on mean sea level pressure.
+The video model generations successfully reproduce both the spatial structure of these patterns and their temporal evolution at comparable propagation speeds to the ICON ground truth.
+At negative lags, we observe a coherent wave train approaching the reference point from the west, and at positive lags, these correlation continue propagating eastward.
+The model’s ability to replicate these spatio-temporal patterns indicates the temporal attention mechanism enables the model to capture the underlying physical dynamics and generate coherent trajectories of atmospheric state.
+Figure 10: Lagged correlation (left) and regression slope (right) maps for mean sea level pressure computed on unconditional video model generations and ICON ground truth.
+Maps are computed relative to a seed point at 45S, 85E (green dot), and averaged over 60 non-overlapping 3-day sequences from December 2024 to May 2025.
 Targets Shape Datasets used for 2D 3D
@@
 Table 2: Input and target variables with dataset sources and dimensionality. X = space, P = levels.
-The inputs and outputs we trained with are displayed in Tab.
-1.
-Briefly, the models are trained to generate 3D states of the atmosphere and some key 2D variables relevant to climate impacts (precipitation, surface air temperature, downwelling short-wave) and observations (top of atmosphere radiative fluxes).
-Our models are trained on two global datasets: the ECMWF Reanalysis version 5 (ERA5)(Hersbach et al., 2020) which is observationally constrained, and a global storm resolving simulation with the ICON climate model (Hohenegger et al., 2023), which is a free-running simulation unpaired with reality.
-ERA5
-ERA5 is a commonly used dataset that represent the best guess of the global atmospheric state at a coarse resolution of 25 km.
-Briefly, it is created by running a data assimilation algorithm that incorporates observations into ECMWF’s numerical model.
-However, it is not available at a high enough resolution in space or time to resolve meso-scale motions on kilometer scale, and in particular deep convection.
-The ICON simulation fills in this gap since it is available at a global 5km resolution, but is not paired with reality.
-For pre-processing, we retrieved hourly ERA5 data for 1980–2018 (inclusive) via the data lake at the National Energy Research Computing Center (NERSC) at $0.25^{\circ}$ resolution on the lat-lon grid.
-The 3d atmospheric states are available on pressure levels.
-We then regridded this using bilinear interpolation to the HEALPix grid with $N_{side}=256$.
-For training the coarse-resolution model, this coarsened by averaging pooling to an $N_{side}=64$.
-We train on years 1980–2017 (inclusive) and validate on 2018, though many of our climate diagnostics below are also computed on the validation set (as is common in generative modeling).
-ICON Cycle 3
-The ICON data is a free-running simulation with the ICON atmospheric model coupled to a dynamic ocean and land.
-The atmospheric component solves the nonhydrostatic fluid mechanics equations on a global ico-sahedral mesh with a resolution of around 5 km.
-Unlike ERA5 the model explicitly resolves certain convective motions so a convection parameterization is not used.
-So even beyond the simple increase in resolution, we expect this to increase the fidelity of the precipitation and cloud fields in this dataset relative to ERA5 where these processes are parameterized.
-Furthermore, because the ocean and land are dynamically coupled to the atmosphere, we expect this run to obey conservation of heat, momentum and moisture between these different components, unlike in the ERA5 where the observations (rather than conservations) are king.
-We obtained O(PB) of ICON data (ngc3028) of the nextGEMS cycle 3 simulations (Koldunov et al., 2023) in zarr format from DKRZ’s JupyterHub server.
-This dataset was stored on the HEALPix grid with $N_{side}=1024$, corresponding roughly to a resolution of $\sqrt{A/(12N_{side}^{2})}=6.3\text{km}$.
-Prior to our handling of the data it was interpolated using nearest neighbors from the native icosahedral grid.
-Five years of this data are available at 3-hour (3D) and 30-minute (2D) resolution in time.
-Conveniently, this dataset featured pre-coarsened data (again using averaging pooling) at our coarse-resolution of $N_{side}=64$.
-We interpolated this coarse data to fixed pressure levels using linear interpolation in the vertical direction.
-To fill in values at pressure level locations that are below the surface, we use extrapolation based on hydrostatic balance constraints and an assumption of a constant temperature lapse rate of $6.5$ K/km for temperature and geo-potential.
-For all other variables, we use constant extrapolation down from the surface.
-This procedure approximately reproduces ERA5’s undocumented procedure for filling-in below-surface levels.
-We train on 2020-01-20 03:00:00 through 2024-03-06 12:00:00 (inclusive) and validate on 2024-03-06 15:00:00 through 2025-07-22 00:00:00.
-AMIP SST Forcing
-Our models take a monthly time-scale SST as an input.
-For the ERA5 data which is paired with reality, we linearly interpolate the adjusted monthly SST time series following the AMIP protocol (Gates et al., 1999).
-Specifically we downloaded the file: input4MIPs.CMIP6Plus.CMIP.PCMDI.PCMDI-AMIP-1-1-9.ocean.mon.tosbcs.gn from the Earth System Grid Federation, and use bilinear interpolation to map it to HPX64 resolution.
-For the ICON data on the other hand, we compute a rolling window average in time with a window size of 30 days based on the HPX64 spatial and daily temporal resolution data in the ICON dataset.
-2.4 Diffusion models
-Background
-Let $p_{\text{data}}(\mathbf{x})$ denote the data distribution with standard deviation $\sigma_{\text{data}}$.
-We define the diffusion process as a pair of Markov chains over the states $\{\mathbf{x}_{0},\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$.
-The forward process gradually corrupts the data by adding i.i.d.
-Gaussian noise with standard deviation $\sigma$, eventually transforming the data into pure noise.
-Specifically, the final state satisfies $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\text{max}})$, where $\sigma_{\text{max}}\gg\sigma_{\text{data}}$.
-The reverse process then attempts to denoise samples from this noisy distribution, ideally recovering samples from the original data distribution, i.e., transforming $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\text{max}})$ back into $\mathbf{x}_{0}\sim p_{\text{data}}(\mathbf{x})$.
-In this work, we adopt the score-based diffusion framework (Song et al., 2021; Mardani et al., 2025), which learns this denoising process via a score-matching network.
-The training objective is given by
-$\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathbf{x}\sim p_{\text{data}}}\mathbb{E}_{\sigma\sim p_{\sigma}}\mathbb{E}_{\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})}\left\|\mathcal{D}_{\theta}(\mathbf{x}+\mathbf{\epsilon};\sigma)-\mathbf{x}\right\|^{2},$ (7)
-where the learned denoiser $\mathcal{D}_{\theta}(\cdot;\sigma)$ approximates the score function according to $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)\approx\frac{\mathcal{D}_{\theta}(\mathbf{x};\sigma)-\mathbf{x}}{\sigma^{2}}$.
-Here, $p_{\sigma}$ denotes the noise level distribution used during training.
-To sample from the model, we employ the Elucidated Diffusion Model (EDM) sampler proposed by (Karras et al., 2022), which solves the corresponding reverse-time ordinary differential equation (ODE):
-$fd\mathbf{x}=-\dot{\sigma}(t)\sigma(t)\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma(t))\,$ (8)
-where $\dot{\sigma}(t)$ is the time derivative of the noise schedule $\sigma(t)$.
-Specifically, we use Algorithm 2 from Karras et al. (2022), with the $S_{churn}$ parameter set to 0.
-Diffusions on the sphere
-To train diffusion models on the sphere, the noise $\epsilon_{i}\sim~{}N(0,1)$ is drawn independently for every pixel $1\leq i\leq 12\cdot 4^{l}$.
-Because the HEALPix pixels are all equal area, we conjecture that this is identical to the pixel-wise average of a finite-variance continuous Gaussian process on the sphere.
-Tuning noise schedules for non-stationary signals
-The choice of noise schedule during training ($p_{\sigma}$) and inference is crucial for the quality of diffusion model.
-We find that settings commonly used for image generation (Karras et al., 2022) do not work well across the dozens of atmospheric observables, each with different spectral characteristics.
-Moreover, compared to image datasets, these data are highly non-stationary in space and time.
-Climatology explains much of the variance of the atmosphere, but the mean of all cat images does not look like a cat.
-In particular, we find that our models cannot produce the seasonal or diurnal cycles using the settings from Karras et al.
-(2022) (see Sec 3.2.4).
-The basic issue is known as the signal leak bias when a model is not trained with enough noise to swamp the signal (Everaert et al., 2023).
-As a rough criterion, the noise levels sampled during training should be large enough to completely mask/unmask the signal when represented in any basis—not just the canonical basis in pixel space.
-Intuitively, if we fail to add enough noise then large-variance signals like season or hour of day can be detected simply by filtering the data enough.
-More concretely, the smallest $\sigma_{min}$ should be chosen so that
-$v^{\prime}(x+\sigma_{min}\epsilon)\approx v^{\prime}x$ (9)
-and the largest so that
-$v^{\prime}(x+\sigma_{max}\epsilon)\approx\sigma_{max}v^{\prime}\epsilon$ (10)
-for any vector $||v||=1$.
-Criteria for $\sigma_{\{min,max\}}$ can be derived by starting from the observation that $E|v^{\prime}\epsilon|^{2}=|v|^{2}=1$, so that (9) and (10) are satisfied by choosing
-$\sigma^{2}_{min}$ $\lesssim\inf_{||v||=1}E|v^{\prime}x|^{2}$ (11)
-$\sigma^{2}_{max}$ $\gtrsim\sup_{||v||=1}E|v^{\prime}x|^{2}.$ (12)
-The signal leak bias occurs when $\sigma_{max}^{2}$ is too small. When $\sigma_{min}$ is too large, the diffusion model will not denoise the lower-amplitude modes of the data.
-Note that (12) is equal to the variance explained by the maximum principal component of the data.
-Recall that the overall variance can be decomposed $n=TrCov(x)=\sum_{i}s^{2}_{i}$ where $n$ is the dimension of $x$ (pixels and channels).
-To understand the behavior for changing $n$ it is useful to discuss the fraction of variance explained by a singular value $\tilde{s}^{2}_{i}=s_{i}^{2}/n$, and expect this fraction to converge as we refine the resolution.
-This implies then, that $\sigma_{max}^{2}\gtrsim n\tilde{s}_{1}^{2}=O(\sqrt{n})$.
-In atmospheric science, the seasonal cycle and diurnal cycles are the dominant modes of variability since they are: 1) planetary in spatial scale; and 2) large in amplitude.
-In mid-latitude land locations, the daily high and low temperature are separated by $\sim 5-10^{\circ}$, while the seasonal cycle is $\sim 10-20^{\circ}$.
-Other variables don’t have as strong a seasonal cycle over as large an area so we take the conservative estimate that the seasonal cycle explains 5% of the global variance across our whole variable set.
-The HPX64 model has about 30 channels in total and 49152 pixels.
-So this would imply a $\sigma_{max}^{2}\gtrsim 300$.
-This is far larger than the range of the hyper parameters used for this resolution in typical image generation (Karras et al., 2022).
-Similar analysis can explain the observation that generating higher-resolution images requires larger noise levels (Chen, 2023).
-To estimate $\sigma_{min}$ one could also use principal component analysis in theory, but in practice this is difficult.
-In practice, power spectral analysis works well for identifying high frequency modes.
-We also often choose $\sigma_{min}$ based on the precision required for practical applications.
-Choosing $\sigma_{min}$ too large manifests as an incorrectly flat spectra for the highest wave numbers (see e.g., the Z500 power spectra presented by Price et al.
-(2024)).
-While one can tune the log-normal distribution to fully cover the dynamic range of the data $[\sigma_{min},\sigma_{max}]$, we have found that sample quality is sensitive to shifts in the mean of the log-normal distribution.
-As a simpler approach, we propose a log-uniform noise schedule
-$p_{\sigma}\propto\sigma^{-1},\sigma_{min}\leq\sigma\leq\sigma_{max}.$ (13)
-Because this is uniform in $\log\sigma$ we call it a “log-uniform” noise distribution.
-In earlier tests on the 2D data alone without SST conditioning, we achieved similar results with the power law distribution used by Price et al.
-(2024).
-In practice, we set $\sigma_{min}=0.02$ and $\sigma_{max}=200$, settings which span the dynamic range of several of our variables (see Figure SI12).
-Although in hindsight we should have decreased it further for the smoothest fields like Z500.
-For inference, we use the same sampler as Karras et al.
-(2022) but with these settings.
-2.5 Multi-diffusion
-Emulating ICON at HPX1024 resolution poses substantial computational challenges due to the massive size of this km-scale global dataset.
-Directly applying a traditional diffusion model to the full-resolution domain is infeasible as it would require over 2,000 GB of GPU memory, assuming a linear relationship between spatial resolution and memory footprint.
-This far exceeds the capacity of current hardware, even on high-end multi-GPU systems.
-To address the computational bottleneck and take advantage of the inherently local nature of super-resolution tasks, multi-diffusion (Bar-Tal et al., 2023) is adapted as a scalable and memory-efficient sampling strategy.
-It enables high-resolution generation without sacrificing global spatial coherence or requiring model architecture changes.
-The central idea behind multi-diffusion is to replace the global denoiser used in standard sampling algorithms with an aggregated local denoiser.
-At each sampling step, the aggregated local denoiser operates on the global feature map in a sliding-window fashion, denoising it patch by patch.
-These locally denoised patches are then reassembled to reconstruct the global feature map.
-This procedure is illustrated in Figure SI5.
-Given a pretrained local denoiser $\Phi_{\theta}$ and a set of spatial sampling operators ${\mathcal{F}_{i}}:\mathcal{X}\rightarrow\mathcal{X}_{\text{local},i}$, the global denoiser can be represented as an aggregation of local denoisers:
-$\mathcal{D}_{\theta}(\textbf{x}_{t})=\frac{\sum_{i}\mathcal{F}_{i}^{-1}\left(\Phi_{\theta}\left(\mathcal{F}_{i}(\textbf{x}_{t})\,\middle|\,\mathcal{F}_{i}(y),\sigma\right)\right)}{\sum_{i}\mathcal{F}_{i}^{-1}(J_{N})},$ (14)
-where $\mathcal{F}_{i}^{-1}:\mathcal{X}_{\text{local},i}\rightarrow\mathcal{X}$ denotes the inverse operation that maps local patches back to the global space, $y\in\mathcal{Y}$ denote the spatial conditions, $\mathcal{X}_{local}=\mathbb{R}^{c_{out}\times N\times N}$ denotes the local target space, $\mathcal{Y}=\mathbb{R}^{c_{in}\times N\times N}$ denotes the local condition space, $N$ is the patch size, and $J\in\mathcal{X}_{\text{local}}$ is an all-one matrix of size $N\times N$ used to account for the number of overlapping patches at each location in the global map.
-A fundamental design principle of multi-diffusion is the global sharing of the latent representation $X$ across all local patches.
-The spatial sampling operators ${\mathcal{F}_{i}}$ define the patch layout, including the extent of overlap between adjacent regions and the total number of patches required to span the full domain.
-Introducing overlap is essential for reducing boundary artifacts and promoting consistency across patches.
-However, increasing the overlap also raises the computational cost, as it leads to a larger number of denoiser evaluations during sampling.
-As demonstrated in the ablation study in Subsection 14, strong performance can be maintained even with minimal overlap, enabling a favorable trade-off between quality and efficiency.
-2.6 Multi-modal masking
-Since we intend our framework to extend to a variety of data sources, with different coverage in space, time, and the included fields, we have designed the model to handle masked training.
-For example, we learned that the ERA5 shortwave and longwave radiation fields are corrupted by a known, but undocumented, bug which appears as clear stripe-like artifacts that point North-West (see Fig 17).
-This is caused be a random seed in the radiation scheme being set to $lat+lon$ in every atmospheric column (Chantry, 2024).
-Therefore, as noted in Tab.
-1, we do not use the fields from ERA5 for training our models, and instead rely on the multi-modal diffusion modeling to fill in this gap.
-Our approach is to simply train a diffusion model to generate the union of all channels present in all the datasets, but have modified the code to support masking.
-This approach may seem inflexible compared to flexible architectures (Bodnar et al., 2024) but it is reasonable to scale since it has long been known how to represent the state of the atmosphere using dense grids of the known state variables.
-Specifically, during training, any masked values in the noisy image $x+n$ and target $x$ are filled with 0 before calling the network and computing the loss.
-We then normalize the loss by the fraction of masked out data points, so that the training equally weights modalities with different numbers of observables.
-Psuedo-code for this approach is shown in SI E.
-2.7 Architectures and hyperparameter settings
-cBottle-3d cBottle-video cBottle-SR
-Number of parameters 149M 282M 330M
-Target channels 3D 3D 2D
-Architecture Song UNet Song UNet Song UNet
-Learned Position Encoding
-Calendar embedding
-Trained on ICON+ERA5 ICON+ERA5 ICON
-Grid HPX64 (49152) HPX64 (49152) 128x128 crops of HPX1024
-Training domain global global crops
-Total examples trained 9,856,000 2,160,000 95,520 samples / 191M patches
-Batch size 64 (3M images) 256 (3-10M) 32 960
-Optimizer Adam Adam SGD
-Learning rate 0.0001 0.0001 0.0001
-Noise distribution ($p_{\sigma}$) log-uniform 0.02–200 log-uniform 0.02–1000 log-normal
-$P_{\text{mean}}=-1.2$ $P_{\text{std}}=1.2$
-Computational Profile
-Hardware trained on 16xH100 32xH100 64xH100
-Examples/GPU-second (train) 6.7 0.32 35.4 patches
-Table 2: Comparison of cBottle model variants.. A training example is defined a single HPX64 image for the the coarse models but is a single 128x128 crop for cBottle-superres.
-We mostly reuse the standard UNet architecture used in diffusion modeling (the so-called Song-UNet in Karras et al.
-(2022); Mardani et al.
-(2025) and others).
-However, it was necessary to add position and temporal embedding (Figure SI2 to allow the diffusion model to learn non-stationary signals in space and time, and temporal attention layers to support video diffusion.
-Table 2 summarizes our three model variants.
-Calendar Embedding
-Producing an accurate seasonal and diurnal cycle required a calendar and hour embedding strategy. We take care to use a embedding that respects the periodicity of these inputs. This is given by
-$f_{4k}$ $=\cos\left(2\pi k\frac{\text{doy}}{365.25}\right)$ (15)
-$f_{4k+1}$ $=\sin\left(2\pi k\frac{\text{doy}}{365.25}\right)$ (16)
-$f_{4k+2}$ $=\cos\left(kt\right)$ (17)
-$f_{4k+3}$ $=\sin\left(kt\right).$ (18)
-Here, the day of year (doy) is a scalar representing the time since the beginning of the year, and the time of day ($t$) is the local solar time in units radians $t=2\pi(\frac{\text{second of day}}{86400}+\frac{\text{lon}}{360})$ [^1] Due to an implementation error we actually used $second-lon$..
-In the architectures below we use $k=1,\ldots,8$, and concatenate these inputs along with other image-like conditioning (e.g., monthly-averaged SST) and noisy target image.
-Learned position embedding
-To allow the diffusion model to learn non-stationary spatial structures like climatology, we add a learned position embedding.
-After the initial convolution embedding of the image-like inputs (calendar, noisy image and monthly-averaged SST), a learned positional embedding for each pixel $49152\times c$ is added.
-The position embedding is initialized with independent standard normal random numbers.
-2.7.1 Video Generation Extension
-Let $x\in\mathbb{R}^{T\times C\times X}$ denote an atmospheric trajectory of $T$ evenly spaced frames with $X$ pixels each.
-While the coarse image cBottle produces realistic atmospheric snapshots, $T$ successive samples in time produce a sequence that lacks temporal coherence due to the independence of each sample.
-Temporal Attention
-We extend cBottle to video generation by augmenting the 2D UNet architecture with temporal self-attention layers, following factorized space-time approaches (Ho et al., 2022; Blattmann et al., 2023).
-The architecture preserves the UNet backbone of cBottle, but introduces temporal layers at the end of spatial blocks.
-Within each UNet block, we first apply spatial modules with the time dimension folded into the batch dimension.
-For blocks with temporal attention, we fold the spatial dimension into the batch dimension, and then apply bidirectional self-attention across frames per pixel, using the channels as the embedding dimension.
-A learned relative-time embedding $R_{\Delta t}\in\mathbb{R}^{2T-1}$ capturing frame-to-frame dependencies based on temporal distance $\Delta t$ is added to the pre-softmax logits.
-The temporal attention output is then residually added to the main branch.
-Masked-Conditioning
-To enable flexible inference capabilities, we adopt a masked-conditioning training scheme inspired by MCVD (Voleti et al., 2022).
-During training, we randomly sample mask patterns $m\in\{0,1\}^{T}$ (where 1 indicates observed frames) from a categorical distribution $p_{mask}$ that blends several tasks: Bernoulli random frame dropout, masking all but the first or last few frames contiguously, masking middle frames while keeping endpoints visible, and full sequence masking.
-On top of the calendar embedding and monthly-mean SST conditioning that the image-model uses, we add this masked sequence (ground truth at unmasked positions and zeros at masked positions) and the binary mask indicator as conditioning concatenated channel-wise with the input.
-The video diffusion training objective is identical to the image diffusion objective, but with this additional per-sample masked-conditioning and the loss averaged across the $T$ frames.
-2.7.2 Model Settings
-cBottle-3D
-The coarse image model generates HPX64 fields.
-The coarse resolution image generator uses the standard settings for the Song-UNet presented by Karras et al.
-(2022), but we found the 3D training was made more stable by increasing base model embedding dimension from 128 to 192.
-The source dataset is identified with a label (ERA5 or ICON) which is randomly dropped for 1/4 samples to encourage the model to learn features across both datasets.
-This label is provided to the UNet in the same manner as the imagenet class label in Karras et al.
-(2022).
-Self-attention is only used on the coarsest resolution stage (8x8).
-All convolutional layers are implemented on the HPX grid using the padding approach described in Section 2.2.1.
-32 channels are devoted to calendar embedding.
-The model was trained from scratch on a 50-50 mix of ICON and ERA5 data on 32 H100 GPUs.
-Even (odd) ranks were trained with ERA5 (ICON).
-The model was trained to 3 million samples (with repeats) with a batch size of 64 ($\sim 50000$ steps) when the log probability loss started to plateau, and then the batch size was increased to 256 until 10 million samples.
-The Adam optimizer (Kingma and Ba, 2014) was used with a learning rate of $10^{-4}$.
-During training the model uses the log-uniform noise distribution ($p_{\sigma}$) with $\sigma_{min}=0.02$ and $\sigma_{min}=200$.
-For inference, we use the power law sampler proposed by Karras et al.
-(2022) with these same settings.
-This unified training approach enables the video model to perform multiple tasks at inference time: generate entire sequences unconditionally, predict future frames given initial states, or infill arbitrary temporal gaps.
-cBottle-video
-The video model generates 3-day sequences with 6-hour resolution ($T=12$ frames) at the same HPX64 resolution of the image-based cBottle.
-Because $T$ correlated frames carry more redundant signal, we keep the same log-uniform noise schedule (with $\sigma_{min}=0.02$) and increase $\sigma_{max}$ from 200 to 1000 in both training and inference to preempt signal leak bias (see Subsection 2.4).
-We insert temporal‐attention layers into every encoder block and into two positions within each decoder resolution stage: the up‑sampling block and the final block.
-We weight each of the four masking tasks defined above as follows: random frame dropout (0.3), block-wise masking (0.25), endpoint interpolation (0.1), and full sequence masking (0.35).
-Similar to the coarse image model, a base model embedding dimension of 128 sufficed to model the 2D dataset’s 12 fields.
-Given the video model’s expanded 124-channel input on the 3D dataset, we increased the UNet width to 256 channels, finding it to converge markedly faster and achieve higher log-likelihood than at 192 channels.
-The model was trained from scratch like the image model on 32 H100 GPUs, using the same Adam optimizer (Kingma and Ba, 2014) and learning rate of $10^{-4}$.
-It was trained to 2.16M video samples with a batch size of 32.
-cBottle-SR
-The super-resolution model, which performs upscaling from HPX64 to HPX1024 resolution (a factor of 16), is trained within a HEALPix multi-diffusion framework described in 2.2.2.
-In this setting, the global input data—including latent variables, conditioning fields, and high-resolution targets—are partitioned into spatial patches.
-Each patch is then independently processed by a local denoising model, which is trained to reconstruct high-resolution content based on temporally and spatially aligned inputs.
-This approach enables efficient learning of fine-scale features while preserving global coherence through patch-wise training across the full spherical domain.
-As illustrated in Figure SI4, this training strategy adapts the standard denoising score matching loss (see Eq. 7) to operate at the patch level. Specifically, the modified objective is defined as
-$\min_{\theta}\mathbb{E}_{\mathbf{x}\sim p_{\text{data}}}\mathbb{E}_{\sigma\sim p_{\sigma}}\mathbb{E}_{\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})}\mathbb{E}_{i\sim\mathcal{U}}\Big{|}\Big{|}\Phi_{\theta}\Big{(}\mathcal{F}_{i}(\mathbf{x})+\mathbf{\epsilon};\mathcal{F}_{i}(y),\sigma\Big{)}-\mathcal{F}_{i}(\mathbf{x})\Big{|}\Big{|}^{2},$ (19)
-Like cBottle-3d we use the Song-UNet with 128 hidden channels. Since it is a local model, we do not need to use the HEALPix padding shown in SI3.
-A patch size of 128$\times$128 is used.
-This is chosen to be much larger than 16x super-resolution factor but is still affordable.
-To cope with non-stationary spatial features like topopgraphy, 20 learnable positional embeddings with a spatial size of HPX1024 grid are introduced and initialized using sinusoidal functions.
-The inputs of the network consist of (1) the HPX64 data re-gridded bilinearly to the HPX1024 128$\times$128 patch in question, (2) the global position embeddings for this patch, and (4) the global HPX64 data interpolated to the 128×128 lat-lon grid.
-These are concatenated with the latent before further processing by the encoder.
-The model was trained to 95520 ICON samples (corresponding to 191 million patches, with repeats) with a batch size of 960.
-The SGD optimizer was used with a learning rate of $10^{-4}$.
-During training the model uses the log-normal noise distribution of Karras et al.
-(2022) ($p_{\sigma}$) with $P_{mean}=-1.2$ and $P_{std}=1.2$.
-For inference, $\sigma_{max}=800$.
-3 Results
-3.1 ERA5 in a Bottle
-3.1.1 Variability
-A useful generator of climate states should exhibit realistic variability across independent samples from phenomena operating on diurnal, synoptic, seasonal, and interannual time scales that, when averaged, produce a realistic climatology.
-To assess this we perform an Atmosphere Model Intercomparison Project (AMIP) control run with prescribed monthly-averaged SSTs (Eyring et al., 2016; Gates et al., 1999).
-We generate samples every hour from 1940 to 2021 (inclusive).
-It should be stressed that each atmospheric state is completely independent in each timestep and only shares the same conditioning.
-This was a large-scale inference, requiring $\sim 1024$ GPU-hours on the H100 GPU.
-We now present a series of common diagnostics used to assess the quality of the generated samples and statistics. Additional analyses are presented in Section G.
-Figure 4: Seasonal cycle of sea ice concentration for ERA5 emulation and ground truth, computed over the test period.
-Climatology and seasonal cycle
-Figure 4 shows that cBottle correctly generates the seasonal evolution of Arctic sea ice concentration.
-As in the ground truth, generated ice approaches 100% concentration during boreal winter (a,f) and spring (b,g).
-During boreal summer (c,h) and fall (d,i) the sea ice margin retreats and its concentration is reduced locally over the Beaufort Sea, Chukchi Sea and Bering Sea – poleward of Alaska and Siberia, also reflected in the annual climatology (e,j).
-These results are evidence that the time-of-year conditioning in cBottle works as intended.
-The seasonal cycle is similarly realistic for other fields (see Figure SI10).
-The annual mean climate of precipitation of the cBottle generated samples also matches the ERA5 and ICON data (see Figure SI11).
-Tropical cyclones
-Consistent with realistic synoptic variability, tropical cyclones are generated in appropriate locations, with appropriate occurrence probability magnitudes (Figure 5).
-This is assessed using a tracking algorithm based on TempestExtremes v2.1 (Ullrich et al., 2021).
-The following filters are applied to identify candidate features: (1) Mean sea level pressure must increase by at least 200 Pa within a 5.5 great-circle distance (GCD) from the candidate point; (2) The difference between geopotential heights at 300 hPa and 500 hPa must decrease by at least 58.8 m2/s2 within a 6.5 GCD; (3) Wind speed must exceed 10 m/s; (4) The latitude of the candidate point must fall within the range of -45 to 45.
-As cBottle lacks temporal coherence, cyclone detections for both Target ERA5 and cBottle-ERA5 are computed without temporal continuity tracking to ensure a fair comparison.
-Figure 5: Tropical cyclone occurrence probability derived from ERA5 (1980–2017) and cBottle (1940–2021), presented as heatmaps normalized per $km^{2}$.
-Diurnal Cycle
-Figure 6: Diurnal cycle diagnostics for surface precipitation. (top row) Amplitude of the diurnal cycle. (bottom row) signal to noise ratio.
+Appendix B Supplemental information
+Mean Sea-Level Pressure Ground Truth Generated
+Refer to caption Refer to caption
+Surface Wind Speed Ground Truth Generated
+Refer to caption Refer to caption
+Visible Light Ground Truth Generated
+Refer to caption Refer to caption
+Surface Precipitation Ground Truth Generated
+Refer to caption Refer to caption
+Figure SI1: More fields for Figure 2
+Figure SI2: Demonstration of super-resolved atmospheric variables outgoing shortwave radiation, and precipitation flux of ICON HPX64 at a selected time during an extreme tropical cyclone event and adjacent midlatitude front.
+From left to right are the input ICON HPX64, HPX1024 super-resolved by cBottle, and ICON HPX1024.
+Figure SI3: Imputation of outgoing shortwave radiation using cBottle.
+cBottle-Infill ICON: ICON data with radiation channels masked and subsequently reconstructed via cBottle.
+cBottle ICON: ICON data fully generated by cBottle using only SST and timestamp as conditioning.
+Target ERA5: ERA5 reference with native spatial resolution and known streaking artifacts.
+cBottle-Infill ERA5: ERA5 data with radiation channels infilled by cBottle.
+B.1 Diurnal Cycle: Details
 The diurnal cycle of rainfall should only be detectable above the noise of other modes of variability in certain subregions of the planet.
-In Figure 6, we compare the amplitude and signal-to-noise ratio of the diurnal cycle in cBottle with ERA5 (see Sec B for details).
-cBottle appears to reproduce even small signals like the locally detectable subtropical stratocumulus drizzle diurnal cycles (see SNR).
-Morever, there is a realistic spatial structure of the precipitation diurnal cycle amplitude including spatial details interior to tropical continents – and even the halo that surrounds the Maritime Continent presumably from the emanating diurnally phased convection propagation there.
-See the SI Fig SI1 for the same analysis with surface air temperature.
-3.1.2 Northern Annular Mode
-We next assess a planetary scale mode of internal variability.
-In the northern hemispheric midlatitudes, the Northern Annular Mode is the leading empirical orthogonal function (EOF), which accounts for one-fifth of overall variance in the 1000 hPa geopotential height (REF).
-Figure 7 shows that the leading EOF computed from cBottle generations exhibits the expected spatial structure of the NAM as well as percent variance explained.
-Moreover, Figure 8 shows that the daily variability of cBottle’s NAM is realistically large and even contains a low frequency envelope that is in phase with interannual fluctuations in the ERA5 NAM state, suggesting learnt associations between the generated NAM state and the SST input to cBottle.
-Figure 7: Arctic Oscillation or the Northern Annular Mode (NAM) computed based on Nov-Apr Z1000 monthly mean anomalies.
-Text overlay indicates the percentage of the total variance attributed to the plotted EOF mode.
-(left) cBottle.
-(right) ERA5.
-Computed over 1970-2010.
-Figure 8: Arctic Oscillation or the Northern Annular Mode (NAM) daily indices between 1980 and 1995 computed as projections of daily Z1000 anomalies onto the ERA5 NAM mode.
-Top - Red/Blue filled curves: ERA5 3 month running mean, solid black line: cBottle 3 month running mean, grey filled region: cBottle ±$\pm 2\sigma$ interval around the mean.
-Bottom - Solid line: ERA5 daily AO index, red circles: cBottle daily AO index.
-3.1.3 ENSO Variability
-Appropriate conditioning on SST is confirmed by focusing on the El Nino Southern Oscillation (ENSO).
-Figure 9 demonstrates that cBottle produces realistic ENSO precipitation variability, measured as the regression of its generated precipitation against the Nino3.4 index associated with its input SST.
+The diurnal amplitude is defined as
+$S:=\frac{1}{366}\sum_{d}\max_{h}\overline{T}(d,h)-\min_{h}\overline{T}(d,h).$
+$\overline{T}(d,h)$ is the average of $T$ as a function of day of year $d$ and hour of day $h$. For the SNR computation, the noise is given by
+$N:=\frac{1}{366}\sum_{d}\frac{1}{24}\sum_{h}s_{T}^{2}(d,h),$
+where $s_{T}^{2}$ is the conditional sample variance. In this computation, $s_{T}^{2}$ and $\overline{T}$ are computed over 1990–2018 (inclusive). Then, the $SNR:=S/N$.
+a. ERA5 b. cBottle
+Refer to caption Refer to caption
+c. ERA5 d. cBottle
+Refer to caption Refer to caption
+Figure SI4: Diurnal cycle diagnostics for surface air temperature. (top row) Amplitude of the diurnal cycle. (bottom row) signal to noise ratio.
+a. ERA5 b. cBottle
+Refer to caption Refer to caption
+c. ERA5 d. cBottle
+Refer to caption Refer to caption
+Figure SI5: Same as Figure SI4 but for total surface precipitation.
+B.2 ENSO Response
+Figure SI6: El Niño-Southern Oscillation.
+Regression coefficient of the Nino3.4 index against monthly average precipitation.
+(left) ERA5.
+(right) Climate in a bottle.
+Computed over 1990-2020 inclusive.
+Figure SI6 demonstrates that cBottle produces realistic ENSO precipitation variability, measured as the regression of its generated precipitation against the Nino3.4 index associated with its input SST.
 Within the tropics, increases of Central Pacific precipitation during the positive phase of ENSO associate with reductions of generated precipitation over tropical Africa, Indonesia and South America, consistent with Walker Cell modulation.
-These results add to the evidence of the cBottle’s ability to condition generated variability appropriately on SST, in addition to time of year and time of day.
-Figure 9: El Niño-Southern Oscillation. Regression coefficient of the Nino3.4 index against monthly average precipitation. (left) ERA5. (right) Climate in a bottle. Computed over 1990-2020 inclusive.
-3.1.4 Non-stationarity: Heat waves
-Figure 10: Time series of global mean anomalies from ERA5’s monthly climatology computed over 1980-2010.
-(a) surface air temperature, (b) precipitation, (c) 500-mb geopotential.
-95% confidence-intervals for the annual mean anomaly are computed by bootstrapping over the monthly averages.
-(d) change in probability of exceeding 90-percentile of Z500 (Z500-90p) computed for 1990-2010 for all land points.
-Also shows Z500-90p trend maps for cBottle-ERA5 (e) and ERA5 (f).
-Figure 10 shows the secular trends in global mean quantities for ERA5 and cBottle.
-cBottle noticeably misses the climate change trend in surface fields (temperature and precipitation), but does better for tropospheric quantities like Z500.
-Taking this caveat, we opt to use Z500 to analyze the trend in extreme heat—this is reasonable since surface heat waves are often linked to anomalous Z500 at atmospheric blocking patterns.
-For example, the 2021 Pacific Northwest heat wave feature +4 standard deviation height anomalies (Mass et al., 2024).
-cBottle is able to reproduce observed trends in the occurrence of heights exceeding the 90th percentile in space (Figure 10e,f), albeit with reduce amplitude in some regions.
-The land averaged trends are also quite similar (Figure 10d).
-3.2 ICON in a bottle
-Generation of ICON states at HPX1024 resolution is achieved through a two-stage process in which the macroscale generator of the ICON modality at HPX64 is fed through a patch-based super-resolution module (see Section 2.7.2), which we assess below.
-All super-resolved results presented here are produced using a consistent configuration: a maximum noise level ($\sigma_{max}$) of 800, 18 sampling steps, and $S_{churn}$=0.
-3.2.1 Super-Resolution Results
-Figure 11: (a), (b) Demonstration of super-resolved atmospheric variables outgoing shortwave radiation, and precipitation flux of ICON HPX64 at a selected time during an extreme tropical cyclone event and adjacent midlatitude front.
-From left to right are the input ICON HPX64, HPX1024 super-resolved by cBottle, and ICON HPX1024.
-(c), (d) Spherical power spectra plots for 2-meter temperature and precipitation of ICON HPX64 super-resolution inferences across 32 validation samples.
-Figure 11 (a), (b) provides a first qualitative assessment of the super-resolution.
+These results add to the evidence of cBottle’s ability to condition generated variability appropriately on SST, in addition to time of year and time of day.
+B.3 Interactivity: Details
+Tropical Cyclones can in principle be requested in any location, and arbitrarily many can be requested in a single sample.
+Some such guidance inputs may be unphysical, such as TCs on the equator or at high latitudes.
+We find that while the guided model can produce such unphysical samples, TCs there tend to be more muted than physical ones, as demonstrated in Fig.
+SI7a).
+Here, we guide the same way as before, but with two locations on the equator and one off the coast of Greenland.
+All three locations show visible pressure minima, albeit somewhat weaker in the equator cases than in Fig.
+3f).
+The guidance could also lead to a loss of variability in the output states.
+In the limit of very strong guidance and many locations, we cannot expect large variability in samples.
+However, in the moderate guidance setting described here, variability qualitatively seems to be preserved: In Fig SI7b) the ensemble standard deviations of a guided and unguided ensemble are similar (if the variability of the guided samples was very low, we would see a mostly blue map).
+a. Unphysical samples b. Variability
+Refer to caption Refer to caption
+Figure SI7: Realism of generated TC states.
+a) The same Figure as Fig.
+3f), but with TC guidance applied in different location.
+b) Difference between guided (TC guidance applied in one location near Florida) and unguided standard deviation in Z500 of a 10 member ensemble of states.
+B.4 ICON in a bottle
+B.4.1 Super-Resolution
+Figure SI2 provides a first qualitative assessment of the super-resolution on its own (starting from low-resolution ICON data).
 Here we choose a convenient time sample from the test set, in which a West Pacific tropical cyclone occurred adjacent to a large midlatitude weather system to its northeast.
@@
 Meanwhile, for the midlatitude weather system, a plausible fine-scale frontal band is generated, extending from the Kiril islands to just east of the Kamchatka peninsula, with consistent horizontal scale and position as the ground truth, across both field variables.
-A more quantitative assessment is shown in Figure 11 (c), (d) from the view of spectra.
+A more quantitative assessment is shown in Figure SI8 from the view of distributions and spectra.
 The plots are generated using 32 randomly selected samples spanning from March 6, 2024, to July 23, 2025.
@@
 Correct distribution of variance as a function of spatial scale is confirmed in the power spectra for temperature and precipitation.
-3.2.2 End-to-end synthesis
+Interestingly, similar quality results can be obtained when only training on small 4-week subset of the 5 year ICON dataset and patching artifacts can be mitigated by the choice of the patching overlap (see Appendix B.7).
+Figure SI8(a,c) shows the probability distributions of surface air temperature and precipitation aggregating over all temporal and spatial locations.
+The super-resolution (red line) captures the shape of temperature and precipitation PDFs including the correct power law regulating the heavy tail of low-likelihood extreme precipitation occurrence (Figure SI8c).).
+Figure SI8: Spherical power spectra and distribution plots of ICON HPX64 super-resolution inferences across 32 validation samples.
+(a), (c): power spectra for 2-meter temperature and precipitation.
+(b), (d): log-scale distributions of 2-meter temperature and precipitation.
+B.4.2 End-to-end synthesis: Coarse generation + Super-resolution
 We now turn to results from our capstone task of attempting to generate 12M-pixel multivariate ICON states purely from noise, time of year, time of day, and a monthly mean map of SST.
-The left column of Figure 12 (a), (b), (c), (d) provides a qualitative demonstration of the results (see also Figure 1).
+The left column of Figure SI9 provides a qualitative demonstration of the results (see also Figure 2).
 Results are shown for an arbitrary subregion spanning the entire North Pacific ocean, across four arbitrary time periods having distinct weather states.
@@
 While difficult to summarize statistically, such multi-scale, multi-variate spatial associations are readily visible to the trained human eye accustomed to inspecting weather data, and suggest successful synthesis.
-Figure 12: (a-d) Demonstration of 2 HPX1024 samples from cascaded cBottle, showing outgoing shortwave radiation, and precipitation flux.
-Samples from cascaded and ICON are not matched.
-(e), (f) Spherical power spectra of cascaded cBottle inferences across 64 validation samples, for 2-meter temperature and precipitation.
-Spectra analysis again quantitatively confirm this finding (Figures 12 (e), (f)).
-On the one hand, secondary imperfections are introduced that were not apparent when super-resolution is used in isolation of macro-scale generation (Figures 11 (c), (d)) such as too low occurrence of the very coldest surface temperatures on the planet (less than -60 Centigrade).
+Figure SI9: Demonstration of 2 HPX1024 samples from cascaded cBottle, showing outgoing shortwave radiation, and precipitation flux. Samples from cascaded and ICON are not matched.
+Figure SI10 presents the probability distributions and spherical power spectra computed from 64 randomly selected samples generated by the cascaded cBottle model, spanning the period from January 3, 2020, to July 23, 2025.
+Spectra analysis again quantitatively confirm the finding.
+On the one hand, secondary imperfections are introduced that were not apparent when super-resolution is used in isolation of macro-scale generation (Figures SI2) such as an increased presence of high-frequency components in the precipitation spectrum.
 On the other hand such imperfections are minor such that the results are overall of reasonable enough quality so as to suggest the viability of cascaded super-resolution as a means to generate multivariate ICON states.
-Figures 12 (e), (f) are generated using 64 randomly selected samples spanning from March 6, 2024, to July 23, 2025.
-3.2.3 Video diffusion
-Spatio-Temporal Analysis
-To evaluate whether the video diffusion model captures realistic atmospheric dynamics through time—beyond just generating independent plausible snapshots—we analyze its ability to reproduce the patterns characteristic of mid-latitude storm track evolution in Figure 13.
-Following the approach of (Rao et al., 2002), we compute one-point lag correlation maps on mean sea level pressure.
-The video model generations successfully reproduce both the spatial structure of these patterns and their temporal evolution at comparable propagation speeds to the ICON ground truth.
-At negative lags, we observe a coherent wave train approaching the reference point from the west, and at positive lags, these correlation continue propagating eastward.
-The model’s ability to replicate these spatio-temporal patterns indicates the temporal attention mechanism enables the model to capture the underlying physical dynamics and generate coherent trajectories of atmospheric state.
-Figure 13: Lagged correlation (left) and regression slope (right) maps for mean sea level pressure computed on unconditional video model generations and ICON ground truth.
-Maps are computed relative to a seed point at 45S, 85E (green dot), and averaged over 60 non-overlapping 3-day sequences from December to May.
-3.2.4 Ablations
-Noise schedule
-Figure 14 demonstrates the impact of noise schedule choice on diffusion models for atmospheric data.
-The bottom row shows results using a log-normal schedule ($P_{mean}=-1.2$, $P_{std}=1.2$) typically used for image generation (Karras et al., 2022).
-This schedule introduces insufficient noise to mask the the dominant diurnal cycle of outgoing shortwave radiation during training.
-Consequently, the model learns to detect and amplify residual signal patterns from noised inputs rather than using the time-of-day conditioning.
-However, during inference where starting from pure noise provides no such signal to extract, the model produces physically inconsistent results.
-Panel (g) shows this most clearly, with a “double sun” artifact with radiation appearing in multiple incorrect locations.
-In contrast, the log-uniform schedule (top row) with higher $\sigma_{max}$ provides sufficient noise to completely obscure the diurnal pattern during training, forcing the model to learn the relationship between time-of-day conditioning and the spatial distribution of solar radiation.
-This yields physically consistent results qualitatively with correctly positioned subsolar points.
-This visual comparison confirms the earlier theoretical argument (see Subsection2.4) that atmospheric variables, some with planetary scale patterns, require substantially higher noise ceilings than is standard for diffusion to avoid signal leak bias.
-Figure 14: Four consecutive frames of reflective visible light (rsut) from a sequence generated by video diffusion models trained with either a log-uniform (top) or log-normal (bottom) noise schedule.
-Generations are conditioned on the calendar embedding and monthly average SST.
-Red × marks indicate the subsolar point at each timestep.
-Training data size ablation
-Figure 15 illustrates the spherical power spectra of inference results from two models trained on datasets of differing sizes: one with 3.75 years of distinct samples (72,360 in total), and another with 4 weeks of distinct samples (1,344 in total), with each week randomly selected from a different season to ensure representative temporal coverage.
-Both models were trained using 95,000 total samples via repetition to match training iterations.
-Across almost all variables, the 4 weeks model closely reproduces the spectral characteristics of the 3.75 years model.
-This alignment is particularly notable given the drastic reduction in training data diversity.
-The only substantial deviation appears in the precipitation spectra, where the 4 weeks model exhibits slightly lower zonal energy at high wavenumbers, suggesting a modest underrepresentation of fine-scale variability.
-These results demonstrate the model’s strong data efficiency: even when trained on a small, seasonally representative subset, it retains robust downscaling performance and preserves key spectral structures across a range of physical variables.
-Figure 15: Training data size ablation study: Power spectra from global super-resolution models trained on 3.75 years versus 4 weeks of timesteps.
-3.3 Multimodal
-In this section we demonstrate how the separate modalities and generative vs.
-super-resolution modules of cBottle can be composed into novel workflows, providing the beginnings of a foundational versatility.
-3.3.1 ICON-Informed Downscaling of ERA5
+In general, the model-generated distributions exhibit good agreement with the target, though the cascaded approach shows a tendency to exaggerate extreme temperature values and slightly overpredict precipitation.
+The super-resolution (Figures SI2) did not have these problems, which suggests that the bias likely comes from the coarse-resolution generations in this cascaded setup.
+Figure SI10: Distribution and spherical power spectra plots of cascaded cBottle inferences across 64 validation samples.
+Column 1: log-scale distributions of 2-meter temperature and precipitation.
+Column 2: power spectra for 2-meter temperature and precipitation.
+B.5 Multimodal: ICON-Informed Downscaling of ERA5
 Although the super-resolution model is trained exclusively on ICON data, it demonstrates zero-shot generalization capability when applied to data from other sources.
 In this subsection, we assess its skills on ERA5 data regridded to HPX64 resolution.
-The corrupted radiation channels in ERA5 have been infilled using the method described in Section 3.3.2.
+The corrupted radiation channels in ERA5 have been infilled using the method described in Section 3.4.
+Please note that this method differs slightly from the translate-then-super-resolve approach in Figure 5h, which translates ERA5 to ICON and then super-resolves instead.
 The resulting super-resolved outputs are compared against both the 3-km High-Resolution Rapid Refresh (HRRR) dataset over CONUS and the native-resolution ICON model at HPX1024.
 All datasets used in this evaluation are regridded to a common 3-km Lambert conformal projection over the CONUS region to ensure fair spatial alignment.
-Figure 16: (a), (b), (c): Illustration of the super-resolved ERA5 output at 15:00 UTC on December 11, 2018, with radiative channels infilled using the cBottle HPX64 generator.
+Figure SI11: Illustration of the downscaled ERA5 output at 15:00 UTC on December 11, 2018, with radiative channels infilled using the cBottle HPX64 generator.
 The first three columns correspond to the same timestamp, while the rightmost column (ICON) shows the free-run forecast for 15:00 UTC on December 11, 2024.
 From top to bottom, the rows represent 10-meter eastward wind velocity, outgoing shortwave radiation, and precipitation.
-(d), (e) Power spectra of cBottle-resolved ERA5 and various baselines, computed over 10 samples.
-All datasets are regridded to a Lambert-Conformal projection for comparison.
-ERA5, cBottle-resolved ERA5, and HRRR share an identical set of evaluation timestamps.
-ICON HPX64 and ICON HPX1024 use matching times of day and days of the year, but from different years.
-Figures 16 (a), (b), (c) presents a qualitative comparison of three representative variables: 10-meter eastward wind velocity, outgoing shortwave radiation, and precipitation rate.
+Figures SI11 presents a qualitative comparison of three representative variables: 10-meter eastward wind velocity, outgoing shortwave radiation, and precipitation rate.
 The figure shows spatial subsets spanning the Midwest to the Eastern United States, with four columns corresponding to the original ERA5 HPX64 input, the super-resolved ERA5 output generated by our model, the HRRR reference, and the ICON HPX1024 reference.
@@
 This suggests that the model effectively emulates high-frequency features in ICON and is capable of generating high-resolution patterns and dynamics that are not explicitly present in the coarse-resolution data.
-The zonal-mean spherical power spectra for four representative variables are shown in Figure SI9 (d), (e): 2-meter temperature, outgoing shortwave radiation, and precipitation.
+Figure SI12: Emulation of native-grid ICON with ERA5: Power spectra of cBottle-resolved ERA5 and various baselines, computed over 10 samples.
+All datasets are regridded to a Lambert-Conformal projection for comparison.
+ERA5, cBottle-resolved ERA5, and HRRR share an identical set of evaluation timestamps.
+ICON HPX64 and ICON HPX1024 use matching times of day and days of the year, but from different years.
+The zonal-mean spherical power spectra for four representative variables are shown in Figure SI12: 2-meter temperature, outgoing shortwave radiation, and precipitation.
 Comparisons are made across five datasets: the ERA5 HPX64 input, the super-resolved ERA5 output generated by our model (cBottle), the HRRR reference, ICON at HPX64, and ICON at HPX1024 resolution.
@@
 Reducing spectral and distributional discrepancies caused by domain shifts—potentially through the inclusion of more diverse training datasets—may further improve the model’s robustness and enhance its applicability in climate science.
-3.3.2 Channel filling
-Following Appendix I.2 of Song et al.
-(2020b), let $\Omega:\mathbb{R}^{C\times\mathbf{d}}\to\mathbb{R}^{C_{\mathcal{K}}\times\mathbf{d}}$ and $\bar{\Omega}:\mathbb{R}^{C\times\mathbf{d}}\to\mathbb{R}^{C_{\mathcal{U}}\times\mathbf{d}}$ denote the slicing operators that extract the known and unknown channels, respectively, where $\mathbf{d}$ denotes an arbitrary set of additional dimensions (e.g., spatial or temporal).
-The unknown subset at sampling time $t$ is then denoted by $\mathbf{z}_{t}=\bar{\Omega}(\mathbf{x}_{t})$.
-The conditional score function required to impute the missing channels can be approximated as
-$\nabla_{\mathbf{z}}\log p\big{(}\mathbf{z}_{t}\mid\Omega(\mathbf{x}_{0}),\sigma_{t}\big{)}\approx\nabla_{\mathbf{z}}\log p\big{(}[\mathbf{z}_{t};\,\Omega(\mathbf{x}_{t})]\mid\sigma_{t}\big{)},$ (20)
-[…]
-To evaluate the imputation capability of the model, we focus on infilling radiation channels in the ERA5 and ICON datasets.
-Specifically, the imputed variables include outgoing shortwave radiation, outgoing longwave radiation, and surface downwelling shortwave radiation.
-Figure 17 visualizes the infilled outgoing shortwave radiation for both ERA5 and ICON, as produced by the proposed cBottle framework.
-All ICON images (Target ICON, cBottle-Infill ICON, and cBottle ICON) correspond to the same timestamp, as do all ERA5 samples.
-Compared to the respective ground truths, cBottle successfully reconstructs plausible radiation structures for both ERA5 and ICON.
-The contrast between cBottle-Infill ICON and cBottle ICON is particularly illuminating: although both are conditioned on identical information (timestamp, day of year, and SST), only the result from the infilling workflow (cBottle-Infill ICON) shows accurate reconstruction of the missing radiation channels in alignment with the Target ICON.
-This highlights that the improvement arises from the imputation technique itself, rather than from conditioning alone.
-Furthermore, cBottle with channel filling proves effective in mitigating dataset-specific artifacts.
-For instance, as seen in the comparison between cBottle-Infill ERA5 and Target ERA5, the model reproduces coherent structures and dynamics while simultaneously removing the streaking artifacts inherent in the original ERA5 fields.
-In summary, these results demonstrate that the imputation approach employed by cBottle not only enables the plausible reconstruction of missing physical variables, but also helps mitigate known deficiencies in climate datasets.
-This underscores the potential of cBottle as a valuable tool for data recovery in atmospheric science, particularly in scenarios where observational records are incomplete, corrupted, or degraded by systematic artifacts.
-Figure 17: Imputation of outgoing shortwave radiation using cBottle.
-cBottle-Infill ICON: ICON data with radiation channels masked and subsequently reconstructed via cBottle.
-cBottle ICON: ICON data fully generated by cBottle using only SST and timestamp as conditioning.
-Target ERA5: ERA5 reference with native spatial resolution and known streaking artifacts.
-cBottle-Infill ERA5: ERA5 data with radiation channels infilled by cBottle.
-3.3.3 Bias correction
-This approach takes advantage of the latent representation of the diffusion model.
-To transition a sample from dataset $i$ to $j$ we evaluate the ODE forward to get the latent then reverse it with the $\beta$-denoiser.
-It may seem surprising that this yields a realistic sample.
-Liu et al.
-(2022) explain that since the flow defined by an ODE is unique, the trajectories between samples cannot cross one another.
-This property makes the flow map close to an optimal transport map between the distributions.
-Figure SI15 shows that cBottle can debias the cloud field of the original ICON data while maintaining temporal coherence across frames.
-The ICON model prefers a liquid-heavy partitioning between cloud water and ice compared to ERA5(cf Fig.
-18).
-It has nearly the double the liquid water as ERA5.
-Overall, the location of the clouds in the bias corrected ICON data is similar, though there are some apparent differences.
-Overall, the ability to translate between datasets with a single foundation model is an exciting potential application, and we leave more detailed analysis to future studies.
-Unlike Liu et al.
-(2022) our model is not trained explicitly to map from one dataset to the other.
-That approach would require training a separate model for each pair of datasets so does not scale well to many different data sources.
-By training one model on all potential datasets, we can use the denoiser to define a shared latent space, a desirable property of a foundation model.
-Figure 18: Bias correction of the total cloud ice.
-4 Conclusions
-We have demonstrated that masked denoising diffusion is a feasible approach to generative modeling of the atmosphere.
-The two-stage cascade comprises: (i) a coarse-resolution multi-modal climate data generator; and (ii) a cascaded patch-based super-resolution module.
-Together, these have the potential to enhance: (i) the way km-scale climate data is accessed; and (ii) the practice of climate informatics.
-In the first stage, provided only 200 kb of inputs comprising two scalars (time of year, time of day) and a 50,000 pixel map of monthly mean SST data, cBottle synthesizes 9 MB of global multivariate samples for 45 channels that emulate the macroscale (HPX64) climatological properties of ERA5 or ICON.
-Our analysis suggests the generated samples exhibit the same mean state and variance from diurnal, synoptic, seasonal and interannual modes of variability as the ground truth.
-In the second stage, given such HPX64 fields, cBottle’s cascaded super-resolution module generates HPX1024 states for 12 channels, totaling 600 MB.
-This stage exploits a patch-based approach to handle scaling to the global (12.5-million pixel) domain.
-We have shown that the super-resolved states have accurate spectra, can be trained with remarkable sample efficiency – as little as four weeks of ICON HPX1024 output – and that secondary patch boundary artifacts can be mostly controlled using simple overlap techniques.
-Either stage can be used independently, for instance, to generate coarse-resolution (HPX64) climate samples, or to super-resolve existing HPX64 climate data to HPX1024 resolution with ICON-learnt high-resolution characteristics.
-The resulting end-to-end system – ICON in a bottle — offers 3000x compression ratio per sample over the raw data (256x compression ratio per channel for 12 channels).
-We suggest this technology can thus greatly reduce the bottleneck that storage, I/O, and data movement presents in modern informatics, which has become especially extreme with the advent of global cloud resolving simulation output.
-This end-to-end system enables users to access realistic km-scale data not by high-latency retrieval of stored data but by low-latency generation of that data, on demand, from a generative ML model.
-Moreover, the modular and multi-modal nature of cBottle demonstrates new possibilities characteristic of AI foundation modeling that are as yet unfamiliar to the climate science community under the prevailing autorogressive paradigm for emulation.
-cBottle exploits climate information across the multiple modalities it is trained on, and transfers learnings across them.
-For instance, we have shown how high-resolution textures of ICON-learnt cloud-scale organization can be in-painted onto the appropriate sectors of synoptic systems of the macroscale ERA5 modality.
-We have also shown how channels missing from one-modality can be readily transferred from another, by in-painting the missing shortwave radiation field in the ERA5 modality exploiting what was learnt from ICON.
-The narrow lesson is that synthetic output from high-resolution simulations can be used more flexibly than it might otherwise appear.
-The broader lesson is more profound: While our proof of concept has focused on just two data modalities for simplicity, we hope it inspires extensions across the many complementary observing and simulating systems of the Earth System, where such cross-modal learning, in-filling, mode-transfer, and bias-correction becomes increasingly useful for climate science and society.
-We expect these extensions to help realize a climate foundation model at the km-scale.
-In terms of methodology, we have introduced diffusion on the HEALPix grid and highlight the importance of the noise schedule when training denoising diffusion models on scientific data.
-In particular we show that the defaults most often used in natural image generation literature are not capable of generating extremely large-scale modes like the seasonal and diurnal cycles.
-Fixing this required two main innovations: i) increasing the range of noise levels sampled during training and inference; and ii) adding appropriate time and position embedding.
-Specifically, we propose tuning the noise schedule based on spectral analysis of the data, and show that the noise schedule should scale with $\sqrt{n}$ where $n$ is the size of the largest-scales modes in number of scalars (pixels $\times$ channels).
-This perspective should be helpful to anyone training diffusion models on scientific data.
-Additional limitations of our preliminary work are important to note.
-Non-stationary trends are not optimally represented in the current version of cBottle; while we have demonstrated that 90th percentile 500 hPa geopotential height exceedance trends can be emulated, analogous trends in surface temperature are currently elusive.
-The mean climate – while reasonable – is imperfect, including 4-10% biases in the zonal mean precipitation rate.
-Another important limitation is that our image generation framework cannot diagnose the duration of events or commonly used metrics like the return time of an event.
-Future work will need to add temporal coherence and extend the system to enable the generation of spatio-temporally coherent time series of phenomena.
-Appendix A Details for Z500-90p analysis
-The methodology mirrors the tx90p analysis feature defined by Seneviratne et al.
-[2021] (See Fig.
-11.9).
-First, percentiles of Z500 at 0 UTC are computed for each day of the year and spatial location using a 14-day rolling window.
-Specifically, we use the climate data operators [Schulzweida, 2023] ydrunpctl command with a 14-day window, and compute percentiles over 1990-2010.
-Then, the fraction of samples exceeding this daily percentile climatology is computed for each year separately.
-The regression coefficients in Figure 10e,f are computed by fitting a regression line with least squares over the entire period we had ERA5 data for 1980-2018 (inclusive).
-Appendix B Diurnal Cycle: Details
-The diurnal amplitude is defined as
-$S:=\frac{1}{366}\sum_{d}\max_{h}\overline{T}(d,h)-\min_{h}\overline{T}(d,h).$
-$\overline{T}(d,h)$ is the average of $T$ as a function of day of year $d$ and hour of day $h$. For the SNR computation, the noise is given by
-$N:=\frac{1}{366}\sum_{d}\frac{1}{24}\sum_{h}s_{T}^{2}(d,h),$
-where $s_{T}^{2}$ is the conditional sample variance. In this computation, $s_{T}^{2}$ and $\overline{T}$ are computed over 1990–2018 (inclusive). Then, the $SNR:=S/N$.
-Figure SI1: Diurnal cycle diagnostics for surface air temperature. (top row) Amplitude of the diurnal cycle. (bottom row) signal to noise ratio.
-Appendix C Architecture: Details
-Figure SI2: Position and calendar embedding.
-Figure SI3: Schematic showing the UNet architecture for spherical geometry with HEALPix grids. Each convolution is applied on HEALPix faces padded as in [Karlbauer et al., 2024].
-Figure SI4: Training workflow of the multi-diffusion framework. Global maps are decomposed into patches for parallel training of local denoisers conditioned on spatially aligned context.
-Figure SI5: Illustration of the aggregated local denoiser replacing the global denoiser in the sampling algorithm.
-Appendix D Datasets: Details
-Appendix E Pseudo-code for masking in the loss function
-D_yn = net(where(mask, y + n, 0), sigma)
-loss = weight * ((D_yn - where(mask, y, 0)) ** 2)
-loss = where(mask, loss, 0) / mask.mean()
-Appendix F Super-resolution results: Details
-Figure SI6: Spherical power spectra and distribution plots of ICON HPX64 super-resolution inferences across 32 validation samples.
-(a), (c): power spectra for 2-meter temperature and precipitation.
-(b), (d): log-scale distributions of 2-meter temperature and precipitation.
-Super-resolution probability distribution
-Figure SI6 shows the probability distributions of surface air temperature and precipitation aggregating over all temporal and spatial locations.
-The super-resolution (red line) captures the shape of temperature and precipitation PDFs (Figure SI6a,c) including the correct power law regulating the heavy tail of low-likelihood extreme precipitation occurrence (Figure SI6c).).
-Cascaded generations probability distribution
-Figure SI7 shows the probability distributions computed using 64 randomly selected samples spanning from March 6, 2024, to July 23, 2025.
-Overall, the distributions match well, though the cascaded model struggles to produce the coldest temperatures (less than -60 deg C).
-The super-resolution did not have this problem, which suggests that the bias likely comes from the coarse-resolution generations in this cascaded setup.
-Figure SI7: Spherical power spectra and distribution plots of cascaded cBottle inferences across 64 validation samples.
-(a), (c): power spectra for 2-meter temperature and precipitation.
-(b), (d): log-scale distributions of 2-meter temperature and precipitation.
-Samples from cascaded and ICON are not matched.
-Multidiffusion Parameter Ablation
+B.6 Further climate diagnostics
+Figure SI13: Loading pattern for the Northern Annular Mode computed from ERA5.
+This is the first empirical orthogonal function of the ERA5 Z1000 monthly anomalies computed over 1970-2010, north of 20 degrees N, and for the winter months of November-April.
+Figure SI14: Cycle of sea ice concentration for ERA5 emulation and ground truth, computed over the test period
+Figure SI15: Seasonal cycle for (JJA- DJF) for ERA5 and cBottle. Panels are shown for Z500, surface air temperature, precipitation, and sea ice concentration.
+B.6.1 Zonal Mean Precipitation
+Figure SI16 examines mean precipitation in cBottle ERA5 and its relationship with latitude computed.
+Overall cBottle ERA5 captures the existence and correct positioning of local maxima associated with midlatitude storm tracks and intertropical convergence zones straddling the equator.
+Local imperfections include positive biases in the northern (+4%) and southern (+11%) tropical rain bands, and as much as 11% underestimation of midlatitude rainfall.
+These differences can in part be explained partially by sampling error.
+Figure SI16: Zonal Average of precipitation ICON, ERA5, and their emulators.
+95% confidence intervals for the validation set means are estimated by bootstrapping over monthly averages.
+Only data from the validation periods for ICON and ERA5 are used.
+B.6.2 Climate change
+a
+b
+c
+Figure SI17: Climate change.
+Time series of global mean anomalies from ERA5’s monthly climatology computed over 1980-2010.
+(a) surface air temperature, (b) precipitation, (c) 500-mb geopotential.
+95% confidence-intervals for the annual mean anomaly are computed by bootstrapping over the monthly averages.
+B.6.3 Power-spectral fidelity of the coarse-resolution generator
+Figure SI18: (left) Spherical power spectra of versus the truth.
+Averages over 100 random samples from the test period are shown.
+The dotted lines indicate the range of noise levels used for the sampling $\sigma_{\{min,max\}}^{2}\frac{4\pi}{n_{pix}}$.
+The factor of $\frac{4\pi}{n_{pix}}$ scales the pixel-wise variance to the same units as the power spectra, which is variance per surface area of the unit sphere.
+Figure SI19: PDF of instantaneous precipitation. Computed for same samples as in SI18.
+Figure SI18 shows that the spectra of the generated coarse samples are consistent with their ICON and ERA5 ground truth.
+For some fields, the ICON data shows significantly more power than ERA5 at large wave numbers, but we have not confirmed if this comes from the aliasing of smaller scale modes due to the block averaging used to coarsen the native HPX1024 outputs for ICON or an actual difference.
+B.6.4 Variance maps
+Figure SI20: Total standard deviation maps with respect to the annual mean computed over the train and test periods for surface air temperature, and precipitation.
+Panels are shown for the ground truth and cBottle for ICON and ERA5.
+The globally-averaged RMS is shown in the label for each subpanel.
+Figure SI20 maps summarize the pointwise variance of samples after removing the time mean (Figure SI20).
+This summarizes dispersion produced by a variety of sources including diurnal, synoptic variability, seasonal, and inter-seasonal.
+Variance of 500 hPa geopotential height (top) is maximized over the Southern Ocean, North Atlantic Sea and northeast Pacific due to midlatitude weather systems, but avoids the tropics due to the weak temperature gradient constraint there [53].
+Midlatitude rainfall variance occurs over the storm tracks including the Gulf Stream and Kuroshio currents.
+Rainfall variance is especially high within the intertropical convergence zones.
+In the special case of surface temperature, diurnal variability leads to a distinct land-sea contrast – a result of solar heating causing strong day-night temperature cycles over land, where surface specific heat capacity is relatively low.
+The main point of Figure SI20 is that cBottle-ERA5 and cBottle-ICON faithfully generate all of these point-wise variance features summarizing all sources variability – and with sufficient precision to detect differences between them, such as the “double ITCZ” – which manifests in cBottle-ICON and cBottle-ICON-Target as a zonal band of excessive dispersion in tropical rainfall in the eastern Pacific just south of the equator.
+B.7 Ablations
+Training data size
+Figure SI21 illustrates the spherical power spectra of inference results from two models trained on datasets of differing sizes: one with 3.75 years of distinct samples (72,360 in total), and another with 4 weeks of distinct samples (1,344 in total), with each week randomly selected from a different season to ensure representative temporal coverage.
+Both models were trained using 95,000 total samples via repetition to match training iterations.
+Across almost all variables, the 4 weeks model closely reproduces the spectral characteristics of the 3.75 years model.
+This alignment is particularly notable given the drastic reduction in training data diversity.
+The only substantial deviation appears in the precipitation spectra, where the 4 weeks model exhibits slightly lower zonal energy at high wavenumbers, suggesting a modest underrepresentation of fine-scale variability.
+These results demonstrate the model’s strong data efficiency: even when trained on a small, seasonally representative subset, it retains robust downscaling performance and preserves key spectral structures across a range of physical variables.
+Figure SI21: Training data size ablation study: Power spectra from global super-resolution models trained on 3.75 years versus 4 weeks of timesteps.
+Noise schedule
+Figure SI22: Four consecutive frames of reflective visible light (rsut) from a sequence generated by video diffusion models trained with either a log-uniform (top) or log-normal (bottom) noise schedule.
+Generations are conditioned on the calendar embedding and monthly average SST.
+Red × marks indicate the subsolar point at each timestep.
+Figure SI22 demonstrates the impact of noise schedule choice on diffusion models for atmospheric data.
+The bottom row shows results using a log-normal schedule ($P_{mean}=-1.2$, $P_{std}=1.2$) typically used for image generation [9].
+This schedule introduces insufficient noise to mask the dominant diurnal cycle of outgoing shortwave radiation during training.
+Consequently, the model learns to detect and amplify residual signal patterns from noised inputs rather than using the time-of-day conditioning.
+However, during inference where starting from pure noise provides no such signal to extract, the model produces physically inconsistent results.
+Panel (g) shows this most clearly, with a “double sun” artifact with radiation appearing in multiple incorrect locations.
+In contrast, the log-uniform schedule (top row) with higher $\sigma_{max}$ provides sufficient noise to completely obscure the diurnal pattern during training, forcing the model to learn the relationship between time-of-day conditioning and the spatial distribution of solar radiation.
+This yields physically consistent results qualitatively with correctly positioned subsolar points.
+This visual comparison confirms the earlier theoretical argument (see SubsectionA.3.1) that atmospheric variables, some with planetary scale patterns, require substantially higher noise ceilings than is standard for diffusion to avoid signal leak bias.
+Multidiffusion overlap
 The overlap size in multidiffusion sampling is a critical factor influencing both computational efficiency and output quality.
@@
 Therefore, selecting an appropriate overlap size is essential to balancing reconstruction fidelity and efficiency.
-To investigate this trade-off, we evaluate various overlap sizes using the median absolute horizontal Sobel gradient of surface temperature, as shown in Figure SI8.
+To investigate this trade-off, we evaluate various overlap sizes using the median absolute horizontal Sobel gradient of surface temperature, as shown in Figure SI23.
 This metric serves as a proxy for boundary artifacts, where higher gradient magnitudes indicate sharper discontinuities.
@@
 The 4-pixel overlap yields slightly stronger artifacts than the 32-pixel case, suggesting that very narrow overlaps may be insufficient to effectively smooth patch boundaries.
-Figure SI8: Severity of patch artifacts across different overlapping pixel sizes, visualized using the mean absolute Sobel-X gradient of surface temperature field as a function of horizontal pixel position.
-Figure SI9: Emulation of native-grid ICON with ERA5: Power spectra of cBottle-resolved ERA5 and various baselines, computed over 10 samples.
-All datasets are regridded to a Lambert-Conformal projection for comparison.
-ERA5, cBottle-resolved ERA5, and HRRR share an identical set of evaluation timestamps.
-ICON HPX64 and ICON HPX1024 use matching times of day and days of the year, but from different years.
-Appendix G Further climate diagnostics
-Figure SI10: Seasonal cycle for (JJA- DJF) for ERA5 and cBottle. Panels are shown for Z500, surface air temperature, precipitation, and sea ice concentration.
-G.1 Zonal Mean Precipitation
-Figure SI11 examines mean precipitation in cBottle ERA5 and its relationship with latitude computed.
-Overall cBottle ERA5 captures the existence and correct positioning of local maxima associated with midlatitude storm tracks and intertropical convergence zones straddling the equator.
-Local imperfections include positive biases in the northern (+4%) and southern (+11%) tropical rain bands, and as much as 11% underestimation of midlatitude rainfall.
-These differences can in part be explained partially by sampling error.
-Figure SI11: Zonal Average of precipitation ICON, ERA5, and their emulators.
-95% confidence intervals for the validation set means are estimated by bootstrapping over monthly averages.
-Only data from the validation periods for ICON and ERA5 are used.
-G.2 Power-spectral fidelity of the coarse-resolution generator
-Figure SI12: (left) Spherical power spectra of versus the truth.
-Averages over 100 random samples from the test period are shown.
-The dotted lines indicate the range of noise levels used for the sampling $\sigma_{\{min,max\}}^{2}\frac{4\pi}{n_{pix}}$.
-The factor of $\frac{4\pi}{n_{pix}}$ scales the pixel-wise variance to the same units as the power spectra, which is variance per surface area of the unit sphere.
-Figure SI13: PDF of instantaneous precipitation. Computed for same samples as in SI12.
-Figure SI12 shows that the spectra of the generated coarse samples are consistent with their ICON and ERA5 ground truth.
-For some fields, the ICON data shows significantly more power than ERA5 at large wave numbers, but we have not confirmed if this comes from the aliasing of smaller scale modes due to the block averaging used to coarsen the native HPX1024 outputs for ICON or an actual difference.
-G.3 Submonthly variance
-Figure SI14: Sub-monthly standard-deviation (variance from the monthly mean averaged over all months in the test period), surface air temperature, and precipitation.
-Panels are shown for the ground truth and cBottle for ICON and ERA5.
-Figure SI14 maps summarize the pointwise variance of samples after removing the seasonal cycle, i.e.
-anomalized with respect to their monthly climatology (Figure SI14).
-This summarizes dispersion produced by a variety of sources including diurnal and synoptic variability.
-Variance of 500 hPa geopotential height (top) is maximized over the Southern Ocean, North Atlantic Sea and northeast Pacific due to midlatitude weather systems, but avoids the tropics due to the weak temperature gradient constraint there [Sobel et al., 2001].
-Midlatitude rainfall variance occurs over the storm tracks including the Gulf Stream and Kuroshio currents.
-Rainfall variance is especially high within the intertropical convergence zones.
-In the special case of surface temperature, diurnal variability leads to a distinct land-sea contrast – a result of solar heating causing strong day-night temperature cycles over land, where surface specific heat capacity is relatively low.
-The main point of Figure SI14 is that cBottle-ERA5 and cBottle-ICON faithfully generate all of these pointwise variance features summarizing diurnal-to-synoptic variability – and with sufficient precision to detect differences between them, such as the “double ITCZ” – which manifests in cBottle-ICON and cBottle-ICON-Target as a zonal band of excessive dispersion in tropical rainfall in the eastern Pacific just south of the equator.
-Appendix H Supplementary Figures
-Figure SI15: Bias correction of the total cloud liquid water. (top) original ICON data and (bottom) bias corrected.
+Figure SI23: Severity of patch artifacts across different overlapping pixel sizes, visualized using the mean absolute Sobel-X gradient of surface temperature field as a function of horizontal pixel position.
+Ensemble of experts and over-fitting
+As discussed in Section A.3.2 the coarse-resolution cBottle-3d is prone to over-fitting the largest modes of the data and propose a regularized ensemble of experts approach as a fix.
+This was initially pointed out to us by astute readers of an early version of this manuscript, who noticed the over-fitting in the Northern Annular Mode (see 4b).
+Intriguingly, this also improves the sample translation abilities demonstrated in Fig 5.
+The translated sample by the regularized approach Fig SI24c much more closely resembles the original sample ( Fig.
+SI24a).
+On the other hand, the over-fit model SI25 hallucinates an entirely new storm system in the south pacific and elsewhere.
+a. ERA5 b. ICON (overfit) c. ICON (regularized)
+Refer to caption Refer to caption Refer to caption
+c. d. e.
+Refer to caption Refer to caption Refer to caption
+Figure SI24: Impact of ensemble of experts regularization on the data-set translation example shown in Fig.
+5.
+(a-c) dBR.
+(a) The original ERA5 dBR field.
+This sample translated to ICON by (b) the over-fit denoiser trained to 9.8M images and (c) the regularized ensemble of experts approach copied from Fig 5d.
+(d-e) is the same but for column integrated liquid water
+Figure SI25: Diffusion loss versus $\sigma$ for different training iterations for the training and a subset of the validation set.
+From upper left, $\sigma\in(0,0.1]$, (0.1, 1], (10, 100], (100, 1000].
+The x-axis is the training progress in 1000s of samples.